### v1

Abstract
We investigate the Steklov eigenvalue problem in an exterior Euclidean domain. First, we present several formulations of this problem and establish the equivalences between them. Next, we examine various properties of the exterior Steklov eigenvalues and eigenfunctions. One of our main findings is an Escobar-type lower bound for the first exterior Steklov eigenvalue on convex domains in dimensions three and higher. This bound is expressed in terms of the principal curvatures of the boundary and is sharp, with equality attained for a ball. Moreover, it implies the existence of a sequence of convex domains with fixed volume and the first exterior Steklov eigenvalues tending to infinity. This contrasts with the interior case, as well as with the two-dimensional exterior case, for which we show that an analogue of the Weinstock isoperimetric inequality holds.
§1.  Introduction and main results
§$1.$  Statement of the problem
Let $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$, be a bounded open set with Lipschitz boundary $\partial\Omega$. The classical Steklov eigenvalue problem in $\Omega$ is given by
$\begin{cases}\Delta u=0\qquad&\text{ in }\Omega,\\ \partial_{\overline{\nu}}u=\sigma u\qquad&\text{ on }\partial\Omega,\end{cases}$ (1.1)
where $\partial_{\overline{\nu}}u=\left\langle\nabla u,\overline{\nu}\right\rangle$ is the normal derivative of $u$ in the direction of the unit normal $\overline{\nu}$ pointing towards the exterior of $\Omega$, and $\sigma$ is a spectral parameter. Since the trace operator $H^{1}(\Omega)\to L^{2}(\partial\Omega)$ is compact, the Steklov spectrum is discrete. It consists of a sequence of eigenvalues
$0=\sigma_{1}(\Omega)\leq\sigma_{2}(\Omega)\leq\sigma_{3}(\Omega)\leq\ldots\nearrow+\infty,$
accumulating at infinity. The Steklov eigenvalues can also be understood as the eigenvalues of the Dirichlet-to-Neumann operator $\mathcal{D}:H^{\frac{1}{2}}(\partial\Omega)\to H^{-\frac{1}{2}}(\partial\Omega)$, defined by $\mathcal{D}:f\mapsto\partial_{\overline{\nu}}(\mathcal{H}f)$, where $\mathcal{H}f\in H^{1}(\Omega)$ is the harmonic extension of a function $f\in H^{\frac{1}{2}}(\partial\Omega)$ from the boundary to the interior. The eigenfunctions of the Dirichlet-to-Neumann operator are the boundary traces of the Steklov eigenfunctions, and they can be chosen to form an orthonormal basis of $L^{2}(\partial\Omega)$. For a comprehensive survey and detailed analysis of the Steklov spectrum on bounded domains (as well as on compact Riemannian manifolds with boundary) we refer to [GirPo17, CoGGS24] as well as to [LeMaPo23, Chapter 7].
The Steklov problem and the Dirichlet-to-Neumann map have many physical applications, notably, to the modelling of diffusion processes with a particular behaviour of the particles when they reach the boundary surface, see [Gr20, GrCh25]. Problem (1.1) arises if the diffusion occurs in the interior of $\Omega$. However, in many situations it is of interest to consider the diffusion in the exterior domain
$\Omega^{\mathrm{ext}}:=\mathbb{R}^{n}\setminus\overline{\Omega}.$
It is therefore natural to look for an analogue of the Steklov problem (1.1) on an unbounded domain $\Omega^{\mathrm{ext}}$. Problems of this type also arise in scattering theory, when modelling the interaction of waves with bounded obstacles, see, e.g., [CaCMM16]. Another motivation for the exterior Steklov problem comes from the study of the exterior Robin eigenvalue problem [KrLo18, KrLo20, KrLo24, Bun25], which is in a sense dual to the Steklov problem, see §1.3.4.
Let us fix some basic notations. A point $x\in\mathbb{R}^{n}$ will be represented in Cartesian coordinates as $x=(x_{1},\dots,x_{n})$ and in polar coordinates as $x=(r,\theta)$, where $r=|x|$ and $\theta\in\mathbb{S}^{n-1}$. We will denote by
$B_{R}=\left\{x\in\mathbb{R}^{n}:|x|<R\right\}$
the ball of radius $R$ centred at the origin, and set
$\Omega^{\mathrm{ext}}_{R}:=\Omega^{\mathrm{ext}}\cap B_{R},$
for each
$R>R_{0}(\Omega):=\inf\,\left\{R>0:\overline{\Omega}\subset B_{R}\right\}.$
We also denote
$H^{1}_{\mathrm{loc}}\left(\Omega^{\mathrm{ext}}\right):=\left\{u:\Omega^{\mathrm{ext}}\to\mathbb{R}:u\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)\text{ for all }R>R_{0}(\Omega)\right\}.$
Simply replacing $\Omega$ by $\Omega^{\mathrm{ext}}$ in (1.1) does not lead to a well-defined exterior Dirichlet-to-Neumann map. Indeed, a harmonic extension of a function to an unbounded domain is not unique unless certain conditions are imposed at infinity. This can be immediately seen by considering the function $f\equiv 0$ on the boundary of the unit disk $B_{1}\subset\mathbb{R}^{2}$: it can be extended harmonically to the exterior of the disk as $u\equiv 0$ or as $u=\log r$. One way to fix the choice of the harmonic extension is to assume that
$u(x)=O\left(|x|^{2-n}\right)\qquad\text{ as }|x|\to\infty,\quad x\in\Omega^{\mathrm{ext}}\subset\mathbb{R}^{n}.$ (1.2)
Theorem 1.1 (see [KacNaPa25, Theorem 3.1] and [McL00, Theorem 8.10]).
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. For any $f\in H^{1/2}(\partial\Omega)$, there exists a unique harmonic extension $u=\mathcal{H}^{\mathrm{ext}}f\in H^{1}_{\mathrm{loc}}\left(\Omega^{\mathrm{ext}}\right)$ satisfying (1.2).
Definition 1.2.
The exterior Steklov problem on $\Omega^{\mathrm{ext}}\subset\mathbb{R}^{n}$ is defined as follows: find $\sigma\in\mathbb{R}$ for which there exists a nonzero $u:\Omega^{\mathrm{ext}}\to\mathbb{R}$ such that $u\in H^{1}_{\mathrm{loc}}\left(\Omega^{\mathrm{ext}}\right)$, and
[…]
In ($\mathrm{ES}$) and further on, $\partial_{\nu}u$ is the normal derivative of $u$ in the direction of the unit normal $\nu=-\overline{\nu}$ pointing towards the exterior of $\Omega^{\mathrm{ext}}$ (and therefore towards the interior of $\Omega$), see Figure 1. ∎
Figure 1: The geometry of an exterior problem.
We denote by $\mathcal{D}^{\mathrm{ext}}:H^{\frac{1}{2}}(\partial\Omega)\to H^{-\frac{1}{2}}(\partial\Omega)$, $\mathcal{D}^{\mathrm{ext}}f=\partial_{\nu}\left(\mathcal{H}^{\mathrm{ext}}f\right)$, the corresponding exterior Dirichlet-to-Neumann operator. The following basic result holds.
Theorem 1.3.
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. The spectrum of the exterior Steklov problem ($\mathrm{ES}$) in $\Omega^{\mathrm{ext}}$ is discrete, and consists of a sequence of eigenvalues
$0\leq\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)\leq\sigma_{2}\left(\Omega^{\mathrm{ext}}\right)\leq\sigma_{3}\left(\Omega^{\mathrm{ext}}\right)\leq\ldots\nearrow+\infty,$
accumulating at infinity. The boundary traces $f_{k}=u_{k}|_{\partial\Omega}$ of the corresponding Steklov eigenfunctions can be chosen to form an orthonormal basis of $L^{2}(\partial\Omega)$.
The proof of this theorem can be essentially deduced from the results of [AuHa14b, AretE15], see §6.1.
Remark 1.4.
Without loss of generality, it will always be assumed in what follows that $\Omega^{\mathrm{ext}}$ is a connected unbounded domain. Otherwise, $\Omega^{\mathrm{ext}}$ would be a disjoint union of finitely many bounded domains and a single connected unbounded domain. Hence, it suffices to consider the exterior Steklov problem on connected $\Omega^{\mathrm{ext}}$ only. ∎
Remark 1.5.
Note that for any planar open set $\Omega$ with connected $\Omega^{\mathrm{ext}}$, the function $u_{1}=\operatorname{const}$ solves ($\mathrm{ES}$) with $\sigma=\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)=0$. In contrast, $\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)$ is strictly positive in dimensions $n\geq 3$, as will follow from Theorems 3.2 and 3.5. ∎
§1.2.  Different approaches to the exterior Steklov problem and their equivalence
One of the difficulties arising in the study of the exterior Steklov problem is that the solutions of ($\mathrm{ES}$) may no longer be square-integrable, like the constant function in dimension two, or the function $1/r$ in the exterior of a ball in $\mathbb{R}^{3}$, see §2.1 and §2.2 for further details. In particular, the space $H^{1}\left(\Omega^{\mathrm{ext}}\right)$ is not suitable for the variational characterisation of exterior eigenvalues.
One of the goals of the present paper is to unify different approaches to the exterior Steklov problem by showing that they are all equivalent in an appropriate sense to Definition 1.2. We list these approaches in an informal way below and provide references to precise formulations:
  • Approach I: Finite energy spaces in dimensions $n\geq 3$. Solve the exterior Steklov problem using variational characterisation ($\mathrm{FE}$) in the space of functions of finite energy, i.e., functions with their gradients in $L^{2}\left(\Omega^{\mathrm{ext}}\right)$, and satisfying a certain decay condition at infinity, see §3.1.2. This approach was developed in [AuHa14b, AretE15].
  • Approach II: Conformal mapping in dimension two. Use a conformal mapping to transform the exterior Steklov problem to a weighted interior Steklov problem ($\mathrm{CT}$), see §3.4.
  • Approach III: Truncated domains. Given $R>R_{0}(\Omega)$, consider a mixed Steklov–Dirichlet eigenvalue problem ($\mathrm{SD}_{R}$) on $\Omega^{\mathrm{ext}}_{R}$ with Steklov condition on $\partial\Omega$ and Dirichlet condition on $\partial B_{R}$, and take the limit of eigenvalues and eigenfunctions as $R\to\infty$, see §3.2. This approach was explored in [AretE15] in dimension $n\geq 3$.
  • Approach IV: Helmholtz equation. Consider the exterior Steklov problem ($\mathrm{H}_{\Lambda}$) for the Helmholtz equation $(-\Delta+\Lambda^{2})u=0$ with a given $\Lambda>0$. The decaying solutions of ($\mathrm{H}_{\Lambda}$) belong to $H^{1}\left(\Omega^{\mathrm{ext}}\right)$, and one can use the variational characterisation of eigenvalues in the usual way. One can then take the limit of eigenvalues and eigenfunctions as $\Lambda\searrow 0$, see §3.3. This approach was developed in [GrCh25, ChrDa23].
  • Approach V: Layer potentials. Finally, one can define the exterior Steklov problem using the layer potential approach, see ($\mathrm{LP}_{n}$) for dimensions $n\geq 3$ and ($\mathrm{LP}_{2}$) for $n=2$, in a similar way to how it is usually done for the interior Steklov problem. We refer to §4 for details. An analogous approach was used in [KacNaPa25].
The equivalence of these formulations is understood as follows. We will show that for the Approaches I, II, and V, the corresponding eigenvalues and eigenfunctions coincide with those of the exterior Steklov problem ($\mathrm{ES}$), see Definition 1.2. For approaches III and IV, we will prove that the $k$th eigenvalues of the problems ($\mathrm{SD}_{R}$) as $R\to\infty$ and ($\mathrm{H}_{\Lambda}$) as $\Lambda\searrow 0$ converge to $\sigma_{k}\left(\Omega^{\mathrm{ext}}\right)$ for all $k\geq 1$, and the corresponding eigenfunctions converge in an appropriate sense to the eigenfunctions $u_{k}$ of the exterior Steklov problem. For schematic relation between different approaches, see Figures 2 and 3. For the precise statements, we refer to §5.
Remark 1.6.
There exist other possible approaches to the exterior Steklov problem. In particular, one can modify the truncated domains approach and impose Neumann conditions on $\partial B_{R}$ instead of the Dirichlet conditions. In two dimensions this formulation is equivalent to the others, see §5.2. At the same time, as was shown in [AretE15], in dimensions $n\geq 3$ the Neumann truncation does not provide an equivalent formulation to the ones described above. In fact, this can be seen from the fact that in dimensions $n\geq 3$ the first exterior Steklov eigenvalue is positive, while the Neumann truncation yields a zero eigenvalue. In these dimensions, the Neumann truncation corresponds instead to the exterior Steklov problem with a decay condition on the gradient (vanishing flow), as described in §4.3.2.
A problem related to ($\mathrm{H}_{\Lambda}$) has been studied in [HelNi26]. Instead of introducing a Helmholtz parameter $\Lambda$, one can consider a magnetic Laplacian and examine the spectral behaviour as the magnetic potential vanishes. It would be interesting to verify whether in the limit one gets the same exterior Steklov problem as above. ∎
FiniteenergyTruncatedDirichletTruncatedNeumannHelmholtzequationLayerPotentialsOur ExternalSteklovVanishingflowTheorem 5.4Theorem 5.6Corollary 4.6Proof of Theorem 1.3Theorem 5.8
Figure 2: Relations between approaches for $n\geq 3$.
ConformalmappingTruncatedDirichletTruncatedNeumannHelmholtzequationLayerPotentialsOur ExternalSteklovVanishingflowTheorem 5.9Theorem 5.13Theorem 5.15Proof of Theorem 1.3Theorem 5.10Remark 4.9
Figure 3: Relations between approaches for $n=2$.
Remark 1.7.
Aside from the pure Steklov conditions on $\partial\Omega$, one can also consider exterior problems with mixed Steklov–Dirichlet–Neumann boundary conditions. This setting naturally arises in some applications, see [HenTW70, Gr25]. The formulations of the exterior problem discussed above, as well as of Theorem 1.3 (with $L^{2}(\partial\Omega)$ replaced by $L^{2}(\partial_{S}\Omega)$, where $\partial_{S}\Omega$ is the part of the boundary on which the Steklov condition is imposed), extend directly to this case. ∎
§1.3.  Spectral properties of the exterior Steklov problem
The approaches outlined in the previous section allow us to prove various properties of eigenvalues and eigenfunctions of the exterior Steklov problem.
1.3.1 Courant’s theorem and multiplicity bounds
Recall that a nodal domain of a continuous function is a connected component of the complement of its nodal set (that is, the set of zero values of the function). One of the fundamental properties of the Laplace and the (interior) Steklov eigenfunctions is Courant’s domain theorem. It is easy to show that it holds for the exterior Steklov eigenfunctions as well.
Theorem 1.8.
Let $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$, be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. An eigenfunction $u_{k}$ associated to $\sigma_{k}\left(\Omega^{\mathrm{ext}}\right)$ has at most $k$ nodal domains.
The result in two dimensions follows immediately from the conformal approach and Courant’s theorem for the interior problem, and in dimensions $n\geq 3$ one can use the finite energy approach and apply the same proof as in the interior case, see [KuSi69]. The proof of Theorem 1.8 is given in §6.1.1.
Courant’s theorem implies that the first exterior Steklov eigenfunction does not change sign in $\Omega^{\mathrm{ext}}$. In dimension two this is clear, since the first eigenfunction is constant, however in higher dimensions it is a priori not obvious. Moreover, we have the following
Corollary 1.9.
The first exterior Steklov eigenvalue $\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)$ is simple.
Once again, this is immediate in dimension two, and the proof in dimension $n\geq 3$ is presented in §6.1.1.
Remark 1.10.
Similarly to the interior problem, there is no direct analogue of Courant’s theorem for the nodal domains of the exterior Dirichlet-to-Neumann map (i.e., for the boundary nodal domains of Steklov eigenfunctions). In two dimensions, Theorem 1.8 and elementary topological considerations yield a bound on the number of boundary sign changes, cf. [GirPo17, Section 6]. Also note that using the results on the exterior Dirichlet-to-Neumann semigroup obtained in [AretE15, Section 5] one can deduce that the first exterior Steklov eigenfunction on any Lipschitz domain can be chosen to be strictly positive on the boundary, cf. [AretEGl20]. ∎
1.3.2 Escobar-type lower bounds for the first eigenvalue
For the interior problem, it was shown by Payne [Pa70] that the first nonzero Steklov eigenvalue on a planar convex bounded domain is estimated from below by the minimal curvature and from above by the maximal curvature of the boundary. Note that both estimates are sharp and attained on a disk. Escobar found a non-sharp extension of Payne’s lower bound to two-dimensional compact manifolds with boundary [Es97]. He also conjectured a sharp bound that in the case of Euclidean domains of arbitrary dimension was proved in [XiaXio24]. It states that if all the principal curvatures of the boundary $\partial\Omega$ are bounded below by a constant $c$ then the first nonzero Steklov eigenvalue $\sigma_{2}(\Omega)\geq c>0$ with equality if and only if $\Omega$ is a ball.
For the exterior Steklov problem in dimension $n\geq 3$ we use the finite energy space approach to prove a stronger Escobar-type bound. We show that the first exterior Steklov eigenvalue is bounded from below by the minimum of the logarithmic (and, hence, the geometric) mean of the curvatures, which itself is bounded from below by the minimum of the principal curvatures.
Let $k\in\mathbb{N}$, and $\alpha_{1},\dots,\alpha_{k}$ be distinct positive real numbers. Their logarithmic mean $L(\alpha_{1},\ldots,\alpha_{k})$ is defined as [Pi85]
$L(\alpha_{1},\ldots,\alpha_{k})=\frac{1}{\sum_{j=1}^{k}\frac{\alpha_{j}^{k-2}}{\prod\limits_{i=1,\,i\neq j}^{k}(\alpha_{j}-\alpha_{i})}\log\alpha_{j}}.$
If the $\alpha_{i}$ are not distinct, $L$ is defined as the limit of the right-hand side.
For $j=1,\ldots,n-1$ and $s\in\partial\Omega$, let $\kappa_{j}(s)$ denote the $j$th principal curvature of $\partial\Omega$, where defined. We adopt the convention that the principal curvatures are nonnegative if $\Omega$ is convex.
Theorem 1.11.
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded, convex domain with $\partial\Omega\in C^{1,1}$. Then,
$\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)\geq(n-2)\inf_{s\in\partial\Omega}L\left(\kappa_{1}(s),\ldots,\kappa_{n-1}(s)\right)=:\beta(\partial\Omega).$ (1.3)
For $\partial\Omega\in C^{1,1}$, the principal curvatures are defined almost everywhere. Accordingly, the infimum in (1.3) and in (1.4) is taken over points where the curvatures are defined.
Additionally, Theorem 1.11 immediately yields a lower bound expressed in terms of the geometric mean of the curvatures. We expect that equality in (1.3) holds if and only if $\Omega$ is a ball, although we were only able to establish this in the setting of Corollary 1.12.
Corollary 1.12.
Suppose that $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded, convex domain with $\partial\Omega\in C^{1,1}$. Then,
$\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)\geq(n-2)\inf_{s\in\partial\Omega}\sqrt[n-1]{\prod_{j=1}^{n-1}\kappa_{j}(s)},$ (1.4)
with equality if and only if $\Omega$ is a ball.
The proofs of Theorem 1.11 and Corollary 1.12 are given in §6.2.
Remark 1.13.
A different lower bound for the first exterior Steklov eigenvalue was obtained in [Xio23, Theorem 1]: for any bounded open set $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, with $C^{1}$ boundary and with $0\in\Omega$, the first exterior Steklov eigenvalue satisfies
$\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)\geq(n-2)\min_{x\in\partial\Omega}\frac{\left\langle x,-\nu\right\rangle}{|x|^{2}}=:\beta_{\mathrm{X}}\left(\partial\Omega\right)$ (1.5)
(the bound is non-trivial only for star-shaped domains $\Omega$). In Example 6.3, we compare (1.5) with our bound from Theorem 1.11 for various spheroids, showing that neither inequality implies the other one. ∎
Remark 1.14.
Note that one cannot replace the minimum of principal curvatures by the minimum of their geometric mean in the original Escobar’s conjecture for the first nonzero interior Steklov eigenvalue. Indeed, consider a prolate spheroid $\mathscr{p}_{a}:=\left\{\frac{x_{1}^{2}}{a^{2}}+\frac{x_{2}^{2}}{a^{2}}+x_{3}^{2}=1\right\}$, with $0<a<1$, which divides $\mathbb{R}^{3}$ into a bounded domain $\mathcal{P}_{a}$ and an unbounded domain $\mathcal{P}^{\mathrm{ext}}_{a}$. One can check that the minimum of the geometric mean of the principal curvatures of $\mathscr{p}_{a}$ (i.e., the square root of the Gaussian curvature) is equal to one for any $a>0$. At the same time, taking the trial function $x_{3}$ in the Rayleigh quotient for $\sigma_{2}\left(\mathcal{P}_{a}\right)$ (note that $x_{3}$ is orthogonal to constants on $\mathscr{p}_{a}$) one can show that $\sigma_{2}\left(\mathcal{P}_{a}\right)\to 0$ as $a\searrow 0$. Therefore, for the interior problem, the geometric mean of the curvatures cannot be used to bound $\sigma_{2}$ from below. ∎
Remark 1.15.
The upper bound in [Pa70] follows directly from the Gauss–Bonnet formula for curves together with Weinstock’s inequality. Since an analogue of Weinstock’s inequality also holds for the exterior problem in two dimensions (see Theorem 1.16 below), the same upper bound extends to the first nonzero exterior Steklov eigenvalue. In contrast, the proof of the lower bound established in [Pa70], does not carry over immediately to the exterior case. It would be interesting to check whether such a lower bound remains valid for the exterior problem. ∎
1.3.3 Isoperimetric upper bounds for the first eigenvalue
Let us start with an analogue of the Weinstock inequality for the first nonzero exterior Steklov eigenvalue $\sigma_{2}$ in dimension two.
Theorem 1.16.
Suppose $\Omega\subset\mathbb{R}^{2}$ is a bounded simply connected Lipschitz domain. Then
$\sigma_{2}\left(\Omega^{\mathrm{ext}}\right)|\partial\Omega|\leq 2\pi.$ (1.6)
with equality if and only if  $\Omega$ is a disk.
Moreover, combining (1.6) with the isoperimetric inequality we obtain
$\sigma_{2}\left(\Omega^{\mathrm{ext}}\right)|\Omega|^{1/2}\leq\sqrt{\pi},$
with equality if and only if $\Omega$ is a disk.
Inequality (1.6) follows from the conformal mapping approach and the usual Weinstock inequality, see the proof in §6.3. The proof of the equality case contains additional steps compared to the interior Steklov problem, because one needs to keep track of the image of the infinity under the conformal mapping.
Remark 1.17.
In fact, the conformal approach essentially allows to extend any result that is valid for the weighted interior Steklov problem in two dimensions to the exterior Steklov problem. In particular, assuming that $\Omega\subset\mathbb{R}^{2}$ is simply connected we obtain the analogues of the Hersch–Payne–Schiffer inequalities [GirPo10a],
$\sigma_{k+1}\left(\Omega^{\mathrm{ext}}\right)|\partial\Omega|\leq 2\pi k,\qquad k\in\mathbb{N},$
as well as the eigenvalue asymptotics (1.10) below. ∎
Consider now the case $n\geq 3$. In the interior case, Brock’s inequality [Bro01] implies that the first nonzero Steklov eigenvalue $\sigma_{2}$ attains its maximum on a ball among all domains of given volume. Moreover, there is a higher-dimensional version of Weinstock’s inequality [BucFNT21] stating that among all convex domains with given surface area, $\sigma_{2}$ is maximised by a ball. It turns out that neither of these results holds in the exterior case.
Theorem 1.18.
Let $n\geq 3$. There exists a sequence of convex smooth bounded domains $\Omega_{m}\subset\mathbb{R}^{n}$ of fixed volume such that $\sigma_{1}\left(\Omega^{\mathrm{ext}}_{m}\right)\to+\infty$ as $m\to\infty$.
The proof of this theorem, see §6.3, is obtained by applying Theorem 1.11 to prolate spheroids. Moreover, by the classical isoperimetric inequality, the sequence in Theorem 1.18 can also be normalised by surface area instead of by volume.
1.3.4 Exterior Robin problem
As a consequence of Theorem 1.18, we obtain the following result for the exterior Robin problem. For $\alpha\in\mathbb{R}$ and a bounded Lipschitz domain $\Omega\subset\mathbb{R}^{n}$ with connected $\Omega^{\mathrm{ext}}$, the lowest point of the spectrum of the Robin Laplacian $\Delta^{\mathrm{Rob},\alpha}$ in $\Omega^{\mathrm{ext}}$ with the parameter $\alpha$ in the Robin condition is given by
$\lambda_{1}^{\alpha}\left(\Omega^{\mathrm{ext}}\right):=\min\operatorname{Spec}\left(-\Delta^{\mathrm{Rob},\alpha}\right)=\inf_{u\in H^{1}\left(\Omega^{\mathrm{ext}}\right)}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x+\alpha\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}{\int_{\Omega^{\mathrm{ext}}}|u|^{2}\,\mathrm{d}x},$
see [KrLo18, KrLo20]. For $n\geq 3$, it was shown in [Bun25, Theorem 1] that
$\lambda_{1}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)<0\qquad\text{if and only if}\qquad\alpha<-\sigma_{1}\left(\Omega^{\mathrm{ext}}\right),$ (1.7)
in which case it is an eigenvalue. Otherwise $\lambda_{1}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)=0=\min\operatorname{Spec}_{\mathrm{ess}}\left(-\Delta^{\mathrm{Rob},\alpha}\right)$ is the bottom of the essential spectrum. Note that in [Bun25], $\Omega$ is assumed to be a domain (and hence connected), but the proof of (1.7) carries over verbatim to any open bounded set $\Omega$ with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$.
Thus, for given $\alpha\in\mathbb{R}$ and $R>0$, Theorem 1.18 ensures the existence of a smooth convex domain $\Omega\subset\mathbb{R}^{n}$ such that $|\Omega|=|B_{R}|$ and $\sigma_{1}(\Omega^{\mathrm{ext}})\geq-\alpha$. Then by (1.7), $\lambda_{1}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)=0$. If we now only consider the values $\alpha\in\left(-\infty,-\sigma_{1}\left(B_{R}^{\mathrm{ext}}\right)\right)$, then we have $\lambda_{1}^{\alpha}\left(B_{R}^{\mathrm{ext}}\right)<0=\lambda_{1}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)$. The preceding argument gives the following
Corollary 1.19.
Let $n\geq 3$. For any $R>0$ and $\alpha<-\sigma_{1}\left(B_{R}^{\mathrm{ext}}\right)=-\frac{n-2}{R}$, there exists a smooth convex domain $\Omega\subset\mathbb{R}^{n}$ with $|\Omega|=|B_{R}|$ satisfying $\lambda_{1}^{\alpha}\left(B_{R}^{\mathrm{ext}}\right)<\lambda_{1}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)$.
It was already observed in [KrLo18] that the ball does not maximise $\lambda_{1}^{\alpha}$ as $\alpha\to$-$\infty$. Corollary 1.19 extends this observation to all relevant values of $\alpha$. The duality between the Robin and Steklov eigenvalues in (1.7) extends to higher eigenvalues and also holds in dimension $n=2$ [Bun25, Corollary 1]. Specifically, for $n\geq 2$, for any bounded open Lipschitz set $\Omega\subset\mathbb{R}^{n}$ with connected $\Omega^{\mathrm{ext}}$, and for $k\in\mathbb{N}$,
$\lambda_{k}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)<0\quad\text{ if and only if }\quad\alpha<-\sigma_{k}\left(\Omega^{\mathrm{ext}}\right).$ (1.8)
In dimension two, $\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)=0$ for any $\Omega$ and therefore $\lambda_{1}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)<0$ whenever $\alpha<0$. This was also shown directly for the first Robin eigenvalue in [KrLo18, Proposition 2]. Moreover, in dimension $n=2$, for any $\alpha<0$ the disk maximises $\lambda_{1}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)$ among smooth bounded simply connected domains $\Omega$ with fixed perimeter or area [KrLo20, Corollary 5], see also [KrLo20, Theorem 4] for results regarding disconnected $\Omega$.
In [KrLo24], Krejčiřík and Lotoreichik study the second Robin eigenvalue $\lambda_{2}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)$ in dimension $n=2$. They observe the existence of a threshold value $\alpha_{*}\left(\Omega^{\mathrm{ext}}\right)<0$ such that $\lambda_{2}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)<0$ if and only if $\alpha<\alpha_{*}\left(\Omega^{\mathrm{ext}}\right)$. In view of (1.8), this threshold satisfies $\alpha_{*}\left(\Omega^{\mathrm{ext}}\right)=-\sigma_{2}\left(\Omega^{\mathrm{ext}}\right)$. They conjecture [KrLo24, Conjecture 1.2] that $\lambda_{2}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)<\lambda_{2}^{\alpha}(B_{R}^{\mathrm{ext}})$ for all bounded simply connected open sets $\Omega$ with the same perimeter or area as $B_{R}$. However, this conjecture remains open and instead they prove that the disk maximises $\lambda_{2}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)$ among convex domains with prescribed elastic energy, given by $\frac{1}{2}\int_{\partial\Omega}\kappa(s)^{2}\,\mathrm{d}S$, and they derive an inequality for $\alpha_{*}\left(\Omega^{\mathrm{ext}}\right)$ in terms of the elastic energy of $\partial\Omega$. While Theorem 1.16 does not settle [KrLo24, Conjecture 1.2], it constitutes a step towards this conjecture. Namely, it implies, for a bounded simply connected domain $\Omega\subset\mathbb{R}^{2}$, with $|\Omega|=|B_{R}|$ and $\Omega\neq B_{R}$, and with $\alpha=-\sigma_{2}\left(B_{R}^{\mathrm{ext}}\right)$, that
$\lambda_{2}^{\alpha}\left(\Omega^{\mathrm{ext}}\right)<0=\lambda_{2}^{\alpha}\left(B_{R}^{\mathrm{ext}}\right).$
It appears reasonable to expect that this inequality also holds for $\alpha$ in a neighbourhood of $\sigma_{2}\left(B_{R}^{\mathrm{ext}}\right)$. A rigorous justification of this, however, is beyond the scope of the present discussion.
1.3.5 Eigenvalue asymptotics
Denote by
$\mathcal{N}_{\Omega^{\mathrm{ext}}}(\sigma)=\#\left\{k\in\mathbb{N}:\sigma_{k}\left(\Omega^{\mathrm{ext}}\right)\leq\sigma\right\}$
the eigenvalue counting function for the exterior problem. Extending the methods of [GirKLP22] to the exterior setting one can prove the following result.
Proposition 1.20.
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded open set with $C^{2,\alpha}$ boundary for some $\alpha>0$, and assume that $\Omega^{\mathrm{ext}}$ is connected. Then
$\mathcal{N}_{\Omega^{\mathrm{ext}}}(\sigma)=\frac{\omega_{n-1}}{(2\pi)^{n-1}}|\partial\Omega|\sigma^{n-1}+O\left(\sigma^{n-2}\right)\quad\text{ as }\sigma\to\infty,$ (1.9)
where $\omega_{n-1}$ denotes the volume of the unit ball in $\mathbb{R}^{n-1}$.
As in the interior case, the remainder estimate in (1.9) is sharp and attained for a ball.
Remark 1.21.
If $\Omega$ is smooth, the layer potential approach yields in a standard way that the exterior Dirichlet-to-Neumann map on $\partial\Omega$ is an elliptic pseudodifferential operator, and the result then follows from the general theory of such operators [Hö68]. Further improvements can be obtained for smooth domains in two dimensions using the conformal approach and the results of [Ro86, GirPPS14]. In particular, if $\Omega$ is a smooth simply connected planar domain, we have
$\sigma_{2k+1}\left(\Omega^{\mathrm{ext}}\right)=\sigma_{2k}\left(\Omega^{\mathrm{ext}}\right)+O\left(k^{-\infty}\right)=\frac{2\pi k}{|\partial\Omega|}+O\left(k^{-\infty}\right)\qquad\text{as }k\to\infty,$ (1.10)
where $O\left(k^{-\infty}\right)$ means that the remainder decays faster than any negative power of $k$. ∎
For domains with Lipschitz boundary, asymptotics (1.9) is expected to hold with the remainder term of order $o\left(\sigma^{n-1}\right)$. We prove this in dimension two using the conformal approach and the results of [KarLaPo23]. It is likely that it could be proved in arbitrary dimension using the methods of [Ro23].
Proposition 1.22.
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded open set with Lipschitz boundary such that $\Omega^{\mathrm{ext}}$ is connected. Then
$\mathcal{N}_{\Omega^{\mathrm{ext}}}(\sigma)=\frac{|\partial\Omega|}{\pi}\sigma+o\left(\sigma\right)\quad\text{ as }\sigma\to\infty.$ (1.11)
Propositions 1.20 and 1.22 are proved in §6.3.3.
§1.4.  Outline of the paper
We begin in §2 with examples where separation of variables allows the explicit computation of exterior Steklov eigenvalues. These are useful in guiding us towards the various formulations of the exterior Steklov problem, some already found in the existing literature, that we present in §3. For planar domains we present an approach based on conformal mappings. In §4 we develop a formulation using boundary layer potentials. The equivalence of the different formulations is established in §5; see Figures 2 and 3. In §6, we study geometric properties of the eigenvalues and eigenfunctions using the various formulations. This section also contains the proofs of the results stated in §1.3, notably a Weinstock-type upper bound in two dimensions and an Escobar-type lower bound in higher dimensions, showing that isoperimetric inequalities fail in the latter setting.
§1.5.  Acknowledgements
This project originated from questions that have been raised in conversations with Bruno Colbois. The authors are very grateful to him for numerous stimulating exchanges related to this work. We would also like to thank Dorin Bucur, Yakar Kannai, Yuri Latushkin, and Nilima Nigam for useful discussions.
Research of AG and IP was partially supported by NSERC and FRQNT. Research of ML was partially supported by the EPSRC. DSG acknowledges the support of the Simons Foundation for his sabbatical stay in 2024 at the CRM in Montréal, Canada.
§2.  Examples
In this section we discuss the exterior Steklov problem on domains for which one can use separation of variables, such as balls and spherical shells (see also [BanWa23, Section 3.2]). As above, we consider the dimensions $n=2$ and $n\geq 3$ separately.
§2.1.  Exterior problem for a disk
We begin by stating a basic but useful property of harmonic functions, which, for example, follows from [Ax86].
Lemma 2.1.
Let $\rho>0$, and consider the disk $B_{\rho}\subset\mathbb{R}^{2}$. Any harmonic function $u:B_{\rho}^{\mathrm{ext}}\to\mathbb{R}$ admits the representation
$u(r,\theta)=a_{0}+b_{0}\log r+\sum_{k=1}^{\infty}\left(a_{k}r^{k}+b_{k}r^{-k}\right)\cos(k\theta)+\sum_{k=1}^{\infty}\left(c_{k}r^{k}+d_{k}r^{-k}\right)\sin(k\theta),$
where $r\in(\rho,\infty)$, $\theta\in(-\pi,\pi]$. This series converges absolutely in $B_{\rho}^{\mathrm{ext}}$ and uniformly on any compact subset.
For $\Omega^{\mathrm{ext}}=B_{\rho}^{\mathrm{ext}}=\{(r,\theta):\theta\in\mathbb{S}^{1},r>\rho\}$, it is easy to check, using Lemma 2.1, that the functions
$u_{(0,1)}(r,\theta):=1,\quad u_{(\ell,1)}(r,\theta):=r^{-\ell}\sin(\ell\theta),\quad u_{(\ell,2)}(r,\theta):=r^{-\ell}\cos(\ell\theta),\qquad\ell\in\mathbb{N},$ (2.1)
are bounded solutions of the exterior Steklov problem ($\mathrm{ES}$) which satisfy the condition (1.2), and their boundary traces form a basis in $L^{2}(\partial B_{\rho})$. The corresponding eigenvalues are given by
$\sigma_{(0)}\left(B_{\rho}^{\mathrm{ext}}\right)=0,\qquad\sigma_{(\ell)}\left(B_{\rho}^{\mathrm{ext}}\right)=\frac{\ell}{\rho},\qquad\ell\in\mathbb{N},$ (2.2)
and have multiplicities
$d_{2,\ell}=\begin{cases}1,\qquad&\text{if }\ell=0,\\ 2,\qquad&\text{if }\ell\in\mathbb{N}.\end{cases}$
They can be re-ordered to form the single non-decreasing sequence $\sigma_{k}\left(B_{\rho}^{\mathrm{ext}}\right)$ of eigenvalues enumerated with multiplicities via
$\sigma_{1}:=\sigma_{(0)},\qquad\sigma_{2\ell}=\sigma_{2\ell+1}:=\sigma_{(\ell)},\qquad\ell\in\mathbb{N}.$
We reiterate that the functions $u_{(\ell,\cdot)}(r,\theta)$ are not in $L^{2}\left(\Omega^{\mathrm{ext}}\right)$ for $\ell\in\{0,1\}$, and thus the usual Sobolev space $H^{1}\left(\Omega^{\mathrm{ext}}\right)$ does not provide a suitable functional setting for the exterior problem.
Let us now illustrate how the truncated domains approach works in this case. In fact, we will do it for both Dirichlet and Neumann truncations, which, as mentioned in Remark 1.6, yield equivalent formulations in dimension two.
For $R>\rho$, consider two mixed problems in the truncated domain $B_{\rho,R}^{\mathrm{ext}}=B_{R}\setminus\overline{B_{\rho}}$ with the Steklov condition on $\partial B_{\rho}$.
We start by imposing the Dirichlet condition on $\partial B_{R}$. Solving the mixed Steklov–Dirichlet problem in $B_{\rho,R}^{\mathrm{ext}}$ using the separation of variables, we get the eigenfunctions
$u_{(0,1)}^{\mathrm{D}}(r,\theta):=1-\frac{\log r}{\log R},$
$u_{(\ell,1)}^{\mathrm{D}}(r,\theta):=\left(r^{-\ell}-\frac{r^{\ell}}{R^{2\ell}}\right)\sin(\ell\theta),\qquad u_{(\ell,2)}^{\mathrm{D}}(r,\theta):=\left(r^{-\ell}-\frac{r^{\ell}}{R^{2\ell}}\right)\cos(\ell\theta),\qquad\ell\in\mathbb{N},$
with the corresponding eigenvalues
$\sigma_{(0)}^{\mathrm{D}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=\frac{1}{\rho(\log R-\log\rho)},\qquad\sigma_{(\ell)}^{\mathrm{D}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=\frac{\ell}{\rho}\frac{1+\left(\frac{\rho}{R}\right)^{2\ell}}{1-\left(\frac{\rho}{R}\right)^{2\ell}},\qquad\ell\in\mathbb{N}$ (2.3)
of multiplicity $d_{2,\ell}$.
On the other hand, imposing the Neumann condition on $\partial B_{R}$ and solving the mixed Steklov–Neumann problem in $B_{\rho,R}^{\mathrm{ext}}$, we obtain the eigenfunctions
$u_{(0,1)}^{\mathrm{N}}(r,\theta):=1,$
$u_{(\ell,1)}^{\mathrm{N}}(r,\theta):=\left(r^{-\ell}+\frac{r^{\ell}}{R^{2\ell}}\right)\sin(\ell\theta),\qquad u_{(\ell,2)}^{\mathrm{N}}(r,\theta):=\left(r^{-\ell}+\frac{r^{\ell}}{R^{2\ell}}\right)\cos(\ell\theta),\qquad\ell\in\mathbb{N},$
and the eigenvalues
$\sigma_{(0)}^{\mathrm{N}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=0,\qquad\sigma_{(\ell)}^{\mathrm{N}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=\frac{\ell}{\rho}\frac{1-\left(\frac{\rho}{R}\right)^{2\ell}}{1+\left(\frac{\rho}{R}\right)^{2\ell}},\qquad\ell\in\mathbb{N}.$ (2.4)
It is easily seen from (2.2), (2.3), and (2.4), that for all $R>\rho$ and all $\ell\in\{0\}\cup\mathbb{N}$, we have
$\sigma_{(\ell)}^{\mathrm{D}}\left(B_{\rho,R}^{\mathrm{ext}}\right)>\sigma_{(\ell)}\left(B_{\rho}^{\mathrm{ext}}\right)>\sigma_{(\ell)}^{\mathrm{N}}\left(B_{\rho,R}^{\mathrm{ext}}\right),$
and that
$\lim_{R\to\infty}\sigma_{(\ell)}^{\mathrm{D}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=\lim_{R\to\infty}\sigma_{(\ell)}^{\mathrm{N}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=\sigma_{(\ell)}\left(B_{\rho}^{\mathrm{ext}}\right).$
Moreover, the eigenfunctions $u_{(\ell,j)}^{\mathrm{D}}$ and $u_{(\ell,j)}^{\mathrm{N}}$ converge pointwise to eigenfunctions (2.1) as $R\to\infty$.
Note that the exterior Steklov spectrum of a disk (2.2) coincides with its interior spectrum (see §3.4 for an explanation using the conformal approach).
Open problem 2.2.
Let $\Omega$ be a bounded planar domain such that its exterior and interior Steklov spectra coincide with the account of multiplicities. Show that $\Omega$ is a disk.
Note that in dimensions $n\geq 3$ the exterior and the interior Steklov spectra never coincide in view of Remark 1.5.
We additionally have
Proposition 2.3.
Consider a disk $B_{\rho}\in\mathbb{R}^{2}$, $\rho>0$, and let $f\in H^{1/2}(\partial B_{\rho})$. Then $\nabla\left(\mathcal{H}^{\mathrm{ext}}f\right)\in L^{2}\left(B_{\rho}^{\mathrm{ext}}\right)$.
Proof.
We expand $f$ in the Fourier series in the boundary traces of eigenfunctions (2.1) on $\partial B_{\rho}$, re-writing them in the complex form, to get
$f=\sum_{m\in\mathbb{Z}}\hat{f}_{m}\rho^{-|m|}\mathrm{e}^{\mathrm{i}m\theta},$
where the condition $f\in H^{1/2}(\partial B_{\rho})$ is equivalent to
$\sum_{m\in\mathbb{Z}}\sqrt{1+|m|^{2}}\rho^{-2|m|}\left|\hat{f}_{m}\right|^{2}<\infty.$ (2.5)
The harmonic extension $u:=\mathcal{H}^{\mathrm{ext}}f$ is then
$u(r,\theta)=\sum_{m\in\mathbb{Z}}\hat{f}_{m}r^{-|m|}\mathrm{e}^{\mathrm{i}m\theta},$
and explicit integration gives
$\left\|\nabla u\right\|^{2}_{L^{2}\left(B_{\rho}^{\mathrm{ext}}\right)}=\sum_{m\in\mathbb{Z}}|m|\rho^{-2|m|}\left|\hat{f}_{m}\right|^{2},$
which is finite by (2.5). ∎
A similar result holds in higher dimensions as well, see Proposition 2.5.
§$2.$  Exterior problem for Euclidean balls in dimensions $n\geq 3$
In order to fix the terminology, let us recall that the spectrum of the Laplace–Beltrami operator $\Delta_{\mathbb{S}^{n-1}}$ on the unit sphere $\mathbb{S}^{n-1}\subset\mathbb{R}^{n}$, $n\geq 3$, consists of eigenvalues $\alpha_{\ell}:=\ell(\ell+n-2)$, $\ell\in\{0\}\cup\mathbb{N}$, of multiplicity
$d_{n,\ell}:=\binom{\ell+n-1}{n-1}-\binom{\ell+n-3}{n-1}$.
The elements of the eigenspace of $\Delta_{\mathbb{S}^{n-1}}$ corresponding to $\alpha_{\ell}$ are known as spherical harmonics of degree $\ell$, and one can choose an orthonormal basis $\left\{Y_{\ell,i}\right\}_{i=1}^{d_{n,\ell}}$ in each such eigenspace.
We solve the exterior Steklov problem ($\mathrm{ES}$) in $B_{\rho}^{\mathrm{ext}}\subset\mathbb{R}^{n}$, $n\geq 3$, by separation of variables in spherical coordinates using spherical harmonics, to obtain the eigenvalues
$\sigma_{(\ell)}\left(B_{\rho}^{\mathrm{ext}}\right):=\frac{n+\ell-2}{\rho},\qquad\ell\in\{0\}\cup\mathbb{N},$ (2.6)
of multiplicity $d_{n,\ell}$, with the corresponding eigenfunctions
$u_{(\ell,i)}(r,\theta)=r^{2-n-\ell}\,{Y}_{\ell,i}(\theta),\qquad\ell\in\{0\}\cup\mathbb{N},\quad i=1,\ldots,d_{n,\ell},$ (2.7)
where $r\in[\rho,\infty)$ and $\theta\in\mathbb{S}^{n-1}$.
Note that, although all the functions $u_{(\ell,i)}$ decay at infinity, the radial eigenfunction $u_{(0,1)}(r,\theta)=r^{2-n}\notin L^{2}\left(\Omega^{\mathrm{ext}}\right)$ for $n=3$ and $n=4$. Once again, this reflects the need for a space larger than $H^{1}\left(\Omega^{\mathrm{ext}}\right)$ in order to provide a functional setting for the exterior Steklov problem.
Consider now the truncated domains approach, see also [CoVe21]. For a mixed Steklov–Dirichlet problem in the spherical shell $B^{\mathrm{ext}}_{\rho,R}$, with the Steklov condition imposed on $\partial B_{\rho}$ and the Dirichlet condition on $\partial B_{R}$, we get the eigenvalues
$\sigma_{(\ell)}^{\mathrm{D}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=\frac{n+\ell-2+\ell\left(\frac{\rho}{R}\right)^{n+2\ell-2}}{\rho\left(1-\left(\frac{\rho}{R}\right)^{n+2\ell-2}\right)},\qquad\ell\in\{0\}\cup\mathbb{N},$
of multiplicity $d_{n,\ell}$, with the corresponding eigenfunctions
$u_{(\ell,i)}^{\mathrm{D}}(r,\theta)=\left(r^{2-n-\ell}-R^{2-n-2\ell}r^{\ell}\right){Y}_{\ell,i}(\theta),\qquad\ell\in\{0\}\cup\mathbb{N},\quad i=1,\ldots,d_{n,\ell},$
Recalling that $n\geq 3$ and taking the limit as $R\to\infty$, we observe the convergence of eigenvalues,
$\lim_{R\to\infty}\sigma_{(\ell)}^{\mathrm{D}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=\sigma_{(\ell)}\left(B_{\rho}^{\mathrm{ext}}\right),\qquad\ell\in\{0\}\cup\mathbb{N},$
and pointwise convergence of eigenfunctions,
$\lim_{R\to\infty}u_{(\ell,i)}^{\mathrm{D}}(r,\theta)=u_{(\ell,i)}(r,\theta),\qquad\ell\in\{0\}\cup\mathbb{N},\quad i=1,\dots,d_{n,\ell},\quad r\in[\rho,\infty),\quad\theta\in\mathbb{S}^{n-1},$
to those of the exterior Steklov problem obtained above.
For a mixed Steklov–Neumann problem in the spherical shell $B^{\mathrm{ext}}_{\rho,R}$ with the Neumann condition on $\partial B_{R}$, we get the eigenvalues
$\sigma_{(0)}^{\mathrm{N}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=0,\qquad\sigma_{(\ell)}^{\mathrm{N}}\left(B_{\rho,R}^{\mathrm{ext}}\right)=\frac{\ell(n+\ell-2)\left(1-\left(\frac{\rho}{R}\right)^{2\ell+n-2}\right)}{\rho\left(\ell+(n+\ell-2)\left(\frac{\rho}{R}\right)^{2\ell}\right)},\qquad\ell\in\mathbb{N}.$
of multiplicity $d_{n,\ell}$, with the corresponding eigenfunctions
$u_{(0,1)}^{\mathrm{N}}(r,\theta)=1,$
$u_{(\ell,i)}^{\mathrm{N}}(r,\theta)=\left(r^{2-n-\ell}+\frac{n+\ell-2}{\ell}R^{2-n-2\ell}r^{\ell}\right)\,{Y}_{\ell,i}(\theta),\qquad\ell\in\mathbb{N},\quad i=1,\ldots,d_{n,\ell}.$
For $\ell\geq 1$ we observe, as in the Dirichlet case in dimensions $n\geq 2$ and in the Neumann case in dimension $n=2$, the convergence of eigenvalues and the pointwise convergence of eigenfunctions, as $R\to\infty$, to those of the exterior Steklov problem. However, it is not the case for the first eigenvalue and eigenfunction (corresponding to $\ell=0$).
Remark 2.4.
For arbitrary domains, we consider eigenfunction convergence in a different sense, see §5.1.1 for details. We note also that although the higher Steklov–Neumann eigenvalues for the ball $B^{\mathrm{ext}}_{\rho,R}$ converge to the corresponding ones for the exterior Steklov problem in $B^{\mathrm{ext}}_{\rho}$ as $R\to\infty$, this property is specific to balls; in general, Neumann truncation is expected to yield different higher eigenvalues in the limit. ∎
We also have the direct analogue of Proposition 2.3, which is easily proved in a similar manner.
Proposition 2.5.
Consider a ball $B_{\rho}\in\mathbb{R}^{n}$, $\rho>0$, $n\geq 3$, and let $f\in H^{1/2}(\partial B_{\rho})$. Then $\nabla\left(\mathcal{H}^{\mathrm{ext}}f\right)\in L^{2}\left(B_{\rho}^{\mathrm{ext}}\right)$.
§2.3.  Numerical examples
Let $\mathcal{K}\subset\mathbb{R}^{2}$ be an asymmetric bounded “kite” domain whose boundary is given parametrically by $\partial\mathcal{K}:=\left\{\left(1.5\cos t+0.7\cos 2t-0.4,1.5\sin t-0.3\cos t\right),t\in[0,2\pi)\right\}$. We show, in Figure 4, numerically computed eigenfunctions of the exterior Steklov problem in $\mathcal{K}^{\mathrm{ext}}$ corresponding to its third and seventh eigenvalues.
Figure 4: Density plots of eigenfunctions of the Steklov problem in the exterior $\mathcal{K}^{\mathrm{ext}}$ of the kite.
To illustrate the differences between interior and exterior Steklov problems, we compare some numerically computed eigenvalues of $\mathcal{K}$ and $\mathcal{K}^{\mathrm{ext}}$ in Table 1; note that, as always in the planar case, we have $\sigma_{1}\left(\mathcal{K}\right)=\sigma_{1}\left(\mathcal{K}^{\mathrm{ext}}\right)=0$.
$k$ $\sigma_{k}\left(\mathcal{K}\right)$ $\sigma_{k}\left(\mathcal{K}^{\mathrm{ext}}\right)$
22 $0.403$ $0.545$
33 $0.524$ $0.571$
44 $1.183$ $1.130$
55 $1.384$ $1.309$
66 $1.721$ $1.746$
77 $2.018$ $1.821$
88 $2.201$ $2.293$
99 $2.706$ $2.450$
$10$ $2.785$ $2.903$
Table 1: Numerically computed Steklov eigenvalues of $\mathcal{K}$ and $\mathcal{K}^{\mathrm{ext}}$.
Additionally, the numerically computed eigenfunctions of the Steklov problem (corresponding to its fourth and eighth eigenvalues) in the exterior of the disjoint union $\mathcal{T}\subset\mathbb{R}^{2}$ of the unit disk centred at the origin, the disk of radius $\frac{2}{3}$ centred at (2,0)(-2,0), and the disk of radius $\frac{3}{2}$ centred at (2,2)(2,-2), are shown in Figure 5.
Figure 5: Density plots of eigenfunctions of the Steklov problem in the exterior $\mathcal{T}^{\mathrm{ext}}$ of the disjoint union of three disks.
For both geometries, the exterior eigenvalues and eigenfunctions are computed using the conformal transform method of §3.4 and the finite element calculations in FreeFEM [Hec12] for the corresponding interior domain problem, see scripts cited at the front page. For an alternative FEM approach to the exterior Steklov problem, see also [GrCh25] and [ChaGr24].
§3.  Different approaches to the exterior Steklov problem
In this section we introduce some of the existing approaches to the exterior Steklov problem.
§3.1.  The Steklov problem for finite-energy functions, $n\geq 3$
We start with the global approach introduced by Arendt and ter Elst [AretE15] for the Dirichlet-to-Neumann operators, as well as by Auchmuty and Han [AuHa14b] for the weak formulation of the Steklov problem. We work directly in function spaces over the full exterior domain. It is in this setting that we will present the spectral theorem for the exterior problem under consideration. It will be considered as a reference problem, that will later be compared to various other methods. The planar case is different and will be discussed separately.
3.1.1 Dirichlet-to-Neumann operators
For $n\geq 3$, Arendt and ter Elst [AretE15] introduced two exterior Dirichlet-to-Neumann operators $A^{\mathrm{N}}$ and $A^{\mathrm{D}}$ on $L^{2}(\partial\Omega)$ using the following two function spaces,
$W\left(\Omega^{\mathrm{ext}}\right):=\left\{u\in H^{1}_{\mathrm{loc}}\left(\Omega^{\mathrm{ext}}\right):\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x<\infty\right\}$
and
$W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right):=W\left(\Omega^{\mathrm{ext}}\right)\cap L^{p}\left(\Omega^{\mathrm{ext}}\right),$
where $p:=\frac{2n}{n-2}>2$, and therefore $\frac{1}{p}=\frac{1}{2}-\frac{1}{n}$.
By [AretE15, Proposition 2.6],
$W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)=\left\{u\in W\left(\Omega^{\mathrm{ext}}\right)\,:\,\langle u\rangle=0\right\},$
where
$\langle u\rangle=\lim_{R\to\infty}\frac{1}{\left|\Omega^{\mathrm{ext}}_{R}\right|}\int_{\Omega^{\mathrm{ext}}_{R}}u\,\mathrm{d}x$ (3.1)
is the average of $u$ over $\Omega^{\mathrm{ext}}$. Because $p>2$, it follows from the Sobolev embedding theorem that $H^{1}\left(\Omega^{\mathrm{ext}}\right)\subset W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)$. This inclusion is strict. Indeed, there are functions in $W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)$ that are not in $L^{2}\left(\Omega^{\mathrm{ext}}\right)$ as their decay at infinity is only fast enough to guarantee that they belong to $L^{p}\left(\Omega^{\mathrm{ext}}\right)$.
The space $W\left(\Omega^{\mathrm{ext}}\right)$, provided with the norm
$\|u\|_{W\left(\Omega^{\mathrm{ext}}\right)}:=\left(\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x+|\langle u\rangle|^{2}\right)^{1/2},$
is complete, the trace operator $\boldsymbol{\tau}:W\left(\Omega^{\mathrm{ext}}\right)\to L^{2}(\partial\Omega)$, $\boldsymbol{\tau}:u\mapsto u|_{\partial\Omega}$, is compact, and the norm defined above is equivalent to $u\mapsto\left(\|\nabla u\|_{L^{2}\left(\Omega^{\mathrm{ext}}\right)}^{2}+\|u\|_{L^{2}(\partial\Omega)}^{2}\right)^{1/2}$, see [AretE15, Proposition 2.5].
Definition 3.1.
The Dirichlet-to-Neumann operator with Neumann condition at infinity ,
$A^{\mathrm{N}}:\operatorname{Dom}\left(A^{\mathrm{N}}\right)\subset L^{2}(\partial\Omega)\to L^{2}(\partial\Omega),$
is the self-adjoint unbounded operator associated with $\left(W\left(\Omega^{\mathrm{ext}}\right),Q,\boldsymbol{\tau}\right)$, where the quadratic form
$Q:W\left(\Omega^{\mathrm{ext}}\right)\to\mathbb{R}$ is given by
$Q[u]:=\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x.$
Similarly, the Dirichlet-to-Neumann operator with Dirichlet condition at infinity,
$A^{\mathrm{D}}:\operatorname{Dom}\left(A^{\mathrm{D}}\right)\subset L^{2}(\partial\Omega)\to L^{2}(\partial\Omega),$
is the self-adjoint operator on $L^{2}(\partial\Omega)$ associated with $\left(W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right),Q,\boldsymbol{\tau}\right)$, see [AretE15]. ∎
The following result follows from [AretE15, Theorem 5.9 and the remark before Proposition 5.3].
Theorem 3.2.
The self-adjoint operators $A^{\mathrm{N}}$ and $A^{\mathrm{D}}$ have the same domain $\operatorname{Dom}\left(A^{\mathrm{N}}\right)=\operatorname{Dom}\left(A^{\mathrm{D}}\right)\subset L^{2}(\partial\Omega)$, and
$A^{\mathrm{D}}h=A^{\mathrm{N}}h+\frac{1}{b}\left(v,A^{\mathrm{D}}1\right)_{L^{2}(\partial\Omega)}(A^{\mathrm{D}}1),$
where $b:=\int_{\partial\Omega}A^{\mathrm{D}}1>0$. Thus, $A^{\mathrm{D}}$ is a bounded rank-one perturbation of $A^{\mathrm{N}}$. Both operators have compact resolvents and their spectra are purely discrete and non-negative, $0=\lambda_{1}\left(A^{\mathrm{N}}\right)\leq\lambda_{2}\left(A^{\mathrm{N}}\right)\leq\cdots\nearrow+\infty,$ and $0<\lambda_{1}\left(A^{\mathrm{D}}\right)\leq\lambda_{2}\left(A^{\mathrm{D}}\right)\leq\cdots\nearrow+\infty$. In both cases, the corresponding eigenfunctions can be chosen to form a complete orthonormal set in $L^{2}(\partial\Omega)$.
We additionally have the following interlacing result.
Theorem 3.3.
For any $k\in\mathbb{N}$,
$\lambda_{k}\left(A^{\mathrm{N}}\right)\leq\lambda_{k}\left(A^{\mathrm{D}}\right)\leq\lambda_{k+1}\left(A^{\mathrm{N}}\right).$
The proof of Theorem 3.3 (presented below) relies on the following abstract result generalising the finite-dimensional Cauchy interlacing theorem [Be97, Chapter 7]. Although the result is elementary, we provide a proof, as we have been unable to locate a reference for unbounded operators.
Proposition 3.4.
Let $\mathcal{A}$ and $\mathcal{B}$ be two self-adjoint semi-bounded below operators in a Hilbert space $H$ with the same domain $D:=\operatorname{Dom}(\mathcal{A})=\operatorname{Dom}(\mathcal{B})$ and with discrete spectra, and suppose that there exists ${h}_{0}\in D$ such that
$\mathcal{B}{h}=\mathcal{A}{h}+\left({h},{h}_{0}\right)_{H}{h}_{0}\qquad\text{for all }{h}\in D.$
Then the eigenvalues of $\mathcal{A}$ and $\mathcal{B}$ interlace,
$\lambda_{k}(\mathcal{A})\leq\lambda_{k}(\mathcal{B})\leq\lambda_{k+1}(\mathcal{A})\qquad\text{for all }k\in\mathbb{N}.$ (3.2)
Proof of Proposition 3.4.
First, for any ${h}\in D$,
$\left(\mathcal{B}{h},{h}\right)_{H}=\left(\mathcal{A}{h},{h}\right)_{H}+\left({h},{h}_{0}\right)_{H}^{2}\geq\left(\mathcal{A}{h},{h}\right)_{H},$
immediately establishing the first inequality in (3.2).
Let $P_{0}^{\perp}$ denote the orthogonal projection on the orthogonal complement of $\operatorname{Span}\{{h}_{0}\}$ in $H$, and let $\mathcal{R}_{\mathcal{A}}[{h}]$ and $\mathcal{R}_{\mathcal{B}}[{h}]$ be the Rayleigh quotients of $\mathcal{A}$ and $\mathcal{B}$ evaluated at ${h}\in D\setminus\{0\}$. Fix $k\in\mathbb{N}$, and recall the variational principle
$\lambda_{k+1}(\mathcal{A})=\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{L}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{A}}[{h}],\qquad\lambda_{k}(\mathcal{B})=\min_{\begin{subarray}{c}\mathcal{M}\subset D\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{M}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{B}}[{h}].$
Take any subspace $\mathcal{L}\subset D$ with $\dim\mathcal{L}=k+1$. Then $\dim P_{0}^{\perp}\mathcal{L}\geq k$. Choose now any subspace $\mathcal{M}\subset P_{0}^{\perp}\mathcal{L}$, $\dim\mathcal{M}=k$. Then restrictions of the operators $\mathcal{A}$ and $\mathcal{B}$ to $\mathcal{M}$ coincide, and therefore, as $\mathcal{M}\subset\mathcal{L}$,
$\max_{\begin{subarray}{c}{h}\in\mathcal{M}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{B}}[{h}]=\max_{\begin{subarray}{c}{h}\in\mathcal{M}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{A}}[{h}]\leq\max_{\begin{subarray}{c}{h}\in\mathcal{L}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{A}}[{h}].$ (3.3)
Let us now take, in the right- and left-hand sides of (3.3), the minimum over all subspaces $\mathcal{L}\subset D$, $\dim\mathcal{L}=k+1$, and all $\mathcal{M}\subset P_{0}^{\perp}\mathcal{L}$, $\dim\mathcal{M}=k$. As the right-hand side of (3.3) is independent of $\mathcal{M}$, we get
$\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\\ \mathcal{M}\subset P_{0}^{\perp}\mathcal{L}\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{L}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{A}}[{h}]=\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{L}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{A}}[{h}]=\lambda_{k+1}(\mathcal{A}).$
At the same time, minimising in the left-hand side over a wider choice of all subspaces $\mathcal{M}\subset D$ with $\dim\mathcal{M}=k$, we get
$\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\\ \mathcal{M}\subset P_{0}^{\perp}\mathcal{L}\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{M}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{B}}[{h}]\geq\min_{\begin{subarray}{c}\mathcal{M}\subset D\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{M}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{B}}[{h}]=\lambda_{k}(\mathcal{B}),$
which together imply the second inequality in (3.2). ∎
Proof of Theorem 3.3.
We apply Proposition 3.4 with $\mathcal{A}=A^{\mathrm{N}}$, $\mathcal{B}=A^{\mathrm{D}}$, $H=L^{2}(\partial\Omega)$, and ${h}_{0}=\frac{A^{\mathrm{D}}1}{\sqrt{b}}$, which we can do since $b=\left(A^{\mathrm{D}}1,1\right)_{L^{2}(\partial\Omega)}\geq\lambda_{1}\left(A^{\mathrm{D}}\right)>0$, and
$A^{\mathrm{D}}{h}=A^{\mathrm{N}}{h}+\left({h},{h}_{0}\right)_{L^{2}(\partial\Omega)}{h}_{0}$
by Theorem 3.2. ∎
3.1.2 Weak formulation of the Steklov problem
It is often convenient to study the Steklov eigenvalue problem directly at the level of quadratic forms, in their weak formulation. For exterior domains, this approach was used by Auchmuty and Han in [AuHa14b], where they introduced a seemingly different function space $E^{1}\left(\Omega^{\mathrm{ext}}\right)$ of real-valued finite-energy functions that decay at infinity [AuHa14b, Section 3]. It is the natural analogue, for exterior domains, of the space $D^{1}(\mathbb{R}^{n})$ studied in [LiLo01, Chapter 8], and consists of Lebesgue measurable functions $u:\Omega^{\mathrm{ext}}\to\mathbb{R}$ satisfying the following three conditions:
  1. (i)
    $u\in L^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ for any $R>R_{0}(\Omega)$,
  2. (ii)
    $|\nabla u|\in L^{2}\left(\Omega^{\mathrm{ext}}\right)$,
  3. (iii)
    $\left\{x\in\Omega^{\mathrm{ext}}:|u(x)|\geq c\right\}$ has finite Lebesgue measure for any $c>0$.
However, as a consequence of [AuHa14b, Corollary 3.4], this is the real version of $W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)$,
$E^{1}\left(\Omega^{\mathrm{ext}}\right)=\left\{f\in W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)\,:\,f\text{ is real valued }\right\},$ (3.4)
for any bounded $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ with Lipschitz boundary and connected $\Omega^{\mathrm{ext}}=\mathbb{R}^{n}\setminus\overline{\Omega}$. Equipped with the inner product introduced in [AuHa14b, formula (3.5)],
$\left(f,g\right)_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}:=\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla f,\nabla g\right\rangle\,\mathrm{d}x+\frac{1}{|\partial\Omega|}\int_{\partial\Omega}fg\,\mathrm{d}S,$
$E^{1}\left(\Omega^{\mathrm{ext}}\right)$ becomes a Hilbert space and the associated norm is equivalent to the norm $\|\cdot\|_{W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)}$.
The following theorem follows from the characterisation [AretE15, Proposition 5.1] of the operator domain $\operatorname{Dom}\left(A^{\mathrm{D}}\right)$ and [AuHa14b, Section 9].
Theorem 3.5.
Let $n\geq 3$. Suppose $\Omega\subset\mathbb{R}^{n}$ is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. A function $f\in L^{2}(\partial\Omega)$ satisfies $A^{\mathrm{D}}f=\lambda f$ if and only $f$ is the trace $\boldsymbol{\tau}(u)$ of a function $u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ such that
$\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla u,\nabla v\right\rangle\,\mathrm{d}x=\lambda\int_{\partial\Omega}uv\,\mathrm{d}S\qquad\text{for all }v\in E^{1}\left(\Omega^{\mathrm{ext}}\right).$ ($\mathrm{FE}$)
Moreover for each $k\geq 1$,
$\lambda_{k}\left(A^{\mathrm{D}}\right)=\min_{\begin{subarray}{c}M\subset E^{1}\left(\Omega^{\mathrm{ext}}\right)\\ \dim(M)=k\end{subarray}}\max_{\begin{subarray}{c}u\in M\\ u\neq 0\end{subarray}}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$ (3.5)
Additionally, the corresponding eigenfunctions $u_{k}$ can be chosen to form a complete orthonormal basis of the subspace $\left\{u\in E^{1}\left(\Omega^{\mathrm{ext}}\right):\Delta u=0\right\}$ and the restrictions of these eigenfunctions to $\partial\Omega$ form a complete orthonormal basis of $L^{2}(\partial\Omega)$.
The positivity of the first eigenvalue $\lambda_{1}$ that was claimed in Theorem 3.2 also immediately follows from the fact that the constant function is not an element of $E^{1}\left(\Omega^{\mathrm{ext}}\right)$.
Remark 3.6.
The approach presented in this section does not extend to dimension $n=2$. In this case, the space $E^{1}\left(\Omega^{\mathrm{ext}}\right)$ is no longer complete with respect to the $\|\cdot\|_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}$-norm; see [AuHa14a, Appendix A].
Even if the decay condition is weakened to $u\in L^{\infty}\left(\Omega^{\mathrm{ext}}\right)$, the resulting space is not complete in dimension two. For example, consider the sequence $\left(f_{m}\right)_{m\in\mathbb{N}}\subset E^{1}\left(B_{1}^{\mathrm{ext}}\right)$ defined by
$f_{m}(x):=\begin{cases}0\qquad&\text{if }|x|<\mathrm{e},\\ \log\log|x|\qquad&\text{if }\mathrm{e}\leq|x|<m,\\ \log\log m\qquad&\text{if }|x|\geq m,\end{cases}$
which defines a Cauchy sequence in $E^{1}\left(B_{1}^{\mathrm{ext}}\right)$, whose limit is not in $L^{\infty}\left(\Omega^{\mathrm{ext}}\right)$. ∎
Remark 3.7.
Another potential approach would be to use weighted Sobolev spaces to enforce some uniqueness of harmonic extensions, see, for instance, [AmGiGi94]. ∎
§3.2.  Truncated domains in dimensions $n\geq 2$
Arendt and ter Elst [AretE15] proposed two ways to approximate the Dirichlet-to-Neumann operator on exterior domains by truncated domains. For any $R>R_{0}(\Omega)$, recalling that $\Omega^{\mathrm{ext}}_{R}=\Omega^{\mathrm{ext}}\cap B_{R}$, we consider the Dirichlet approximation,
$\begin{cases}\Delta u_{R}=0\qquad&\text{ in }\Omega^{\mathrm{ext}}_{R},\\ \partial_{\nu}u_{R}=\sigma^{\mathrm{D}}u_{R}\qquad&\text{ on }\partial\Omega,\\ u_{R}=0\qquad&\text{ on }\partial B_{R},\end{cases}$ ($\mathrm{SD}_{R}$)
with the spectral parameter $\sigma^{\mathrm{D}}=\sigma^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$, and the Neumann approximation,
$\begin{cases}\Delta u_{R}=0\qquad&\text{ in }\Omega^{\mathrm{ext}}_{R},\\ \partial_{\nu}u_{R}=\sigma^{\mathrm{N}}u_{R}\qquad&\text{ on }\partial\Omega,\\ \partial_{\nu}u_{R}=0\qquad&\text{ on }\partial B_{R}.\end{cases}$ ($\mathrm{SN}_{R}$)
with the spectral parameter $\sigma^{\mathrm{N}}=\sigma^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$.
We can interpret ($\mathrm{SD}_{R}$) as a spectral problem for the self-adjoint operator $A_{R}^{\mathrm{D}}$ in $L^{2}(\partial\Omega)$ that is associated to the quadratic form
$Q_{R}^{\mathrm{D}}[u]=\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u|^{2}\,\mathrm{d}x,$
defined on
$\operatorname{Dom}\left(Q_{R}^{\mathrm{D}}\right)=H_{D}^{1}\left(\Omega^{\mathrm{ext}}_{R}\right):=\left\{u\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right):u|_{\partial B_{R}}=0\right\},$
and to the compact trace operator $\boldsymbol{\tau}:H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)\to L^{2}(\partial\Omega)$. Similarly, we understand ($\mathrm{SN}_{R}$) as the spectral problem for the self-adjoint operator $A_{R}^{\mathrm{N}}$ in $L^{2}(\partial\Omega)$ associated to the quadratic form
$Q_{R}[u]=\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u|^{2}\,\mathrm{d}x,\quad\operatorname{Dom}(Q_{R})=H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$
and the trace operator $\boldsymbol{\tau}$. These are well-known interior mixed problems: the spectra of both operators are purely discrete and their eigenvalues can be characterised by
$\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ $=\inf_{\begin{subarray}{c}M\subset H^{1}_{D}\left(\Omega^{\mathrm{ext}}_{R}\right)\\ \dim(M)=k\end{subarray}}\ \sup_{\begin{subarray}{c}u\in M\\ u\neq 0\end{subarray}}\frac{\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S},$ (3.6)
$\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ $=\inf_{\begin{subarray}{c}M\subset H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)\\ \dim(M)=k\end{subarray}}\ \sup_{\begin{subarray}{c}u\in M\\ u\neq 0\end{subarray}}\frac{\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$ (3.7)
The following monotonicity results follow directly from (3.6) and (3.7).
Lemma 3.8.
Let $n\geq 2$. Suppose $\Omega\subset\mathbb{R}^{n}$ is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. For any $R>R_{0}(\Omega)$ we have
$\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)\leq\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right),\qquad k\in\mathbb{N},$
and for any $R_{1}>R>R_{0}(\Omega)$, we additionally have
$\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R_{1}}\right)\geq\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right),\qquad\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R_{1}}\right)\leq\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right),\qquad k\in\mathbb{N}.$
Therefore, both limits $\lim\limits_{R\to\infty}\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ and $\lim\limits_{R\to\infty}\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ exist.
For dimensions $n\geq 3$, it was already proved in [AretE15] that the truncated domain approaches are equivalent to the approaches from §3.1.1.
Theorem 3.9 ([AretE15, Theorems 5.5 and 5.6]).
For $n\geq 3$, the resolvent of $A_{R}^{\mathrm{D}}$ and the resolvent of $A_{R}^{\mathrm{N}}$ converge uniformly to the resolvents of the operators $A^{\mathrm{D}}$ and $A^{\mathrm{N}}$, respectively. Thus, $\lim\limits_{R\to\infty}\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)=\lambda_{k}\left(A^{\mathrm{D}}\right)$ and $\lim\limits_{R\to\infty}\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)=\lambda_{k}\left(A^{\mathrm{N}}\right)$.
Therefore, by Theorem 3.5, the truncated Dirichlet approach is equivalent to the energy space approach. For the analysis in the case of planar domains, we refer to §5.2.
§$3.$  Helmholtz equation in dimension $n\geq 2$
An alternative approach to obtaining a discrete spectrum, followed by Grebenkov and Chaigneau in [GrCh25], is to introduce a fixed parameter $\Lambda>0$ and consider the spectral problem
$\begin{cases}(\Lambda^{2}-\Delta)u=0\qquad&\text{ in }\Omega^{\mathrm{ext}},\\ \partial_{\nu}u=\mu u\qquad&\text{ on }\partial\Omega,\end{cases}$ ($\mathrm{H}_{\Lambda}$)
where $\mu=\mu\left(\Omega^{\mathrm{ext}},\Lambda\right)$ is the spectral parameter. Since $\Lambda^{2}>0$, the corresponding bilinear form is $H^{1}\left(\Omega^{\mathrm{ext}}\right)$-coercive. In [Au12, Sections 4 and 8], it is demonstrated that coercivity can be used to show that the spectrum of a Steklov problem on bounded domains is purely discrete. Similarly, by applying the same modifications to [Au12] as in, e.g., [AuHa13, Theorem 5.1], we deduce that the spectrum of ($\mathrm{H}_{\Lambda}$) is purely discrete, with the eigenvalues
$\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)=\min_{\begin{subarray}{c}M\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)\\ \dim(M)=k\end{subarray}}\max_{\begin{subarray}{c}u\in M\\ u\neq 0\end{subarray}}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x+\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S},$ (3.8)
and with the eigenfunctions belonging to $H^{1}\left(\Omega^{\mathrm{ext}}\right)$. In addition, the eigenfunctions are all exponentially decaying as $|x|\to\infty$ due to
Lemma 3.10 ([BarMe77]).
Let $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$, be a bounded domain with a Lipschitz boundary $\partial\Omega$ and $\Lambda>0$. For any function $u\in H^{1}\left(\Omega^{\mathrm{ext}}\right)$ which satisfies $(\Lambda^{2}-\Delta)u=0$ in $\Omega^{\mathrm{ext}}$, there exists a constant $C>0$ such that
$|u(x)|\leq C|x|^{\frac{n-1}{2}}\mathrm{e}^{-\Lambda|x|}$
for sufficiently large $|x|$.
Since $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$ is monotonically increasing in $\Lambda$ for $\Lambda>0$, the limit
$\mu_{k}\left(\Omega^{\mathrm{ext}}\right):=\lim_{\Lambda\searrow 0}\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$
exists. The asymptotic behaviour of $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$ has been studied in [GrCh25], providing both theoretical results on the convergence rate and numerical examples.
Example 3.11.
For a given $\Lambda>0$, functions satisfying $(\Lambda^{2}-\Delta)f=0$ in the exterior of $\Omega=B_{\rho}\subset\mathbb{R}^{n}$ with $n\geq 2$ are of the form
$f(r,\theta)=\sum_{\ell=0}^{\infty}\sum_{i=1}^{d_{n,\ell}}r^{-\frac{n-2}{2}}\left(b_{\ell,i}K_{\frac{n+2\ell-2}{2}}(\Lambda r)+c_{\ell,i}I_{\frac{n+2\ell-2}{2}}(\Lambda r)\right){Y}_{\ell,i}(\theta),$
where $I_{m}(\cdot)$ and $K_{m}(\cdot)$ are the modified Bessel functions of the first and second kind, respectively, of order $m$, and $b_{\ell,i},c_{\ell,i}$ are constants. In view of the asymptotic behaviour of the Bessel functions (see, e.g., [DLMF, Chapter 10]), we have to choose $c_{\ell,i}=0$ in order ensure the decay at infinity. Since $K_{m}(\Lambda r)$ decays exponentially for $r\to\infty$, the resulting function belongs to $H^{1}\left(\Omega^{\mathrm{ext}}\right)$. So, the Steklov eigenvalue problem ($\mathrm{H}_{\Lambda}$) in $B_{\rho}^{\mathrm{ext}}$ has the eigenfunctions (with arbitrary scaling factors $b_{\ell,i}$)
$u_{(\ell,i)}(r,\theta,\Lambda)=b_{\ell,i}r^{1-\frac{n}{2}}K_{\ell+\frac{n}{2}-1}(\Lambda r){Y}_{\ell,i}(\theta),\qquad\ell\in\{0\}\cup\mathbb{N},\qquad i=1,\dots,d_{n,\ell},$ (3.9)
associated to the eigenvalues
$\mu_{(\ell)}\left(B_{\rho}^{\mathrm{ext}},\Lambda\right)=\frac{\Lambda K_{\ell+\frac{n}{2}}(\Lambda\rho)}{K_{\ell+\frac{n}{2}-1}(\Lambda\rho)}-\frac{\ell}{\rho},\qquad\ell\in\{0\}\cup\mathbb{N},$
of multiplicity $d_{n,\ell}$.
We now consider the cases of the balls of dimension $n\geq 3$ and disks in the planar case $n=2$ separately.
We start with the case $n\geq 3$. Using $K_{m}(x)=\frac{\Gamma(m)2^{m}}{2x^{m}}+o\left(x^{-m}\right)$ as $x\to 0$ for $m>0$, we obtain
$\lim_{\Lambda\searrow 0}\mu_{(\ell)}\left(B_{\rho}^{\mathrm{ext}},\Lambda\right)=\frac{n+2\ell-2}{\rho}-\frac{\ell}{\rho}=\frac{n+\ell-2}{\rho},$
which are exactly the eigenvalues given by (2.6). Choosing the normalising constants $b_{\ell,i}=\frac{\left(\frac{\Lambda}{2}\right)^{\ell+\frac{n}{2}-1}}{2\Gamma(\ell+\frac{n}{2}-1)}$ in (3.9), we get
$\lim_{\Lambda\searrow 0}u_{(\ell,i)}(r,\theta,\Lambda)=r^{-(n+\ell-2)}{Y}_{\ell,i}(\theta),$
see (2.7). So, we recover the original eigenvalues and eigenfunctions as $\Lambda\searrow 0$.
For $n=2$, we have the same form of the eigenfunctions as in (3.9). For $\ell\geq 1$, we act as before. For $\ell=0$, we use additionally $K_{0}(x)=-\log\frac{x}{2}+O(1)$ as $x\searrow 0$. This yields
$\lim_{\Lambda\searrow 0}\mu_{(0)}\left(B_{\rho}^{\mathrm{ext}},\Lambda\right)=0.$
Furthermore, with $b_{0}=-\frac{1}{2\log\frac{\Lambda}{2}}$, we obtain
$\lim_{\Lambda\searrow 0}u_{0}(r,\theta)=1.$
Therefore, as in the case of higher-dimensional balls, the eigenvalues and eigenfunctions of ($\mathrm{H}_{\Lambda}$) in the exterior of a disk converge to those of ($\mathrm{ES}$).
We emphasise that in these examples the convergence of eigenfunctions is established only pointwise; in §5.2.2 we provide a stronger statement. ∎
§3.4.  Exterior problem for planar domains via conformal mappings
In this section, we present a new formulation of the Steklov eigenvalue problem in exterior domains in two dimensions, based on conformal mappings.
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. We assume that the origin $0\in\Omega$, which can be always achieved by a change of variables, and we identify $\mathbb{R}^{2}$ with $\mathbb{C}$ by interpreting $x=\left(x_{1},x_{2}\right)\in\mathbb{R}^{2}$ as $z=x_{1}+\mathrm{i}x_{2}\in\mathbb{C}$.
Let
$\phi:\mathbb{C}\setminus\{0\}\to\mathbb{C}\setminus\{0\},\qquad\phi(z)=\frac{1}{z},$ (3.10)
and set
$\Omega^{*}:=\phi\left(\Omega^{\mathrm{ext}}\right)\cup\{0\}.$
We remark that $\phi$ is a conformal diffeomorphism between $\Omega^{\mathrm{ext}}$ and $\Omega^{*}\setminus\{0\}$. The connectedness of $\Omega^{\mathrm{ext}}$ implies that $\Omega^{*}$ is connected; furthermore, if $\Omega$ is connected, then $\Omega^{*}$ is simply connected. See Figure 6 for an illustration.
Figure 6: The kite domain $\mathcal{K}$ from §2.3, its exterior $\mathcal{K}^{\mathrm{ext}}$, and the bounded domain $\mathcal{K}^{*}=\phi\left(\mathcal{K}^{\mathrm{ext}}\right)\cup\{0\}$.
As in [GirPo17], the key idea is to pull back the eigenvalue problem on $\Omega^{\mathrm{ext}}$ to the bounded domain $\Omega^{*}$. This yields the transformed eigenvalue problem
$\begin{cases}\Delta w=0&\text{ in }\Omega^{*},\\ \partial_{\nu}w=\xi|\phi^{\prime}(z)|w&\text{ on }\partial\Omega^{*},\end{cases}$ ($\mathrm{CT}$)
where $\xi$ is the new spectral parameter, and $\left|\phi^{\prime}(z)\right|=\frac{1}{|z|^{2}}.$
As always, we understand ($\mathrm{CT}$) in the weak sense. It may seem natural to formulate ($\mathrm{CT}$) on $\Omega^{*}\setminus\{0\}$, but any bounded harmonic function $u:\Omega^{\mathrm{ext}}\to\mathbb{R}$ admits a limit at infinity. Consequently, the transformed function $w=u\circ\phi$ extends smoothly to all of $\Omega^{*}$. Also, we note that inverting with respect to a point in $\Omega$ other than the origin, does not affect the eigenvalues.
Since $\Omega^{*}$ is a bounded Lipschitz domain, the eigenvalues of ($\mathrm{CT}$) admit the characterisation
$\xi_{k}\left(\Omega^{\mathrm{ext}}\right)=\inf_{\begin{subarray}{c}M\subset H^{1}(\Omega^{*})\\ \dim(M)=k\end{subarray}}\sup_{0\neq w\in M}\frac{\int_{\Omega^{*}}|\nabla w|^{2}\,\mathrm{d}z}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|w|^{2}\,\mathrm{d}S}.$ (3.11)
We will show later that (3.11) allows us to adapt classical arguments by Weinstock to derive isoperimetric-type inequalities for the first non-trivial Steklov eigenvalue. First, however, we establish the equivalence between the formulation ($\mathrm{CT}$) and the approaches described in §3.2 and §3.3.
Remark 3.12.
This conformal mapping approach is specific to two dimensions. While the Kelvin transform preserves harmonicity in $\mathbb{R}^{n}$ for $n\geq 3$, the boundary condition becomes significantly more complicated under this transformation. ∎
§4.  Boundary layer potentials
Numerical methods for solving boundary value problems often involve reformulating the boundary value problem on $\Omega$ as an integral equation over $\partial\Omega$, using single and double layer potentials. This approach is also applicable to boundary value problems on $\Omega^{\mathrm{ext}}$, provided that the solutions have a suitable behaviour at infinity, see [McL00, pp. 234–239], [St07, Section 7.5] and also [ChWGLS12].
Recently in [KacNaPa25], the Dirichlet-to-Neumann map in exterior domains has been studied using layer potentials. Building on these results, we formulate the exterior Steklov problem using layer potentials. While this is closely related to [KacNaPa25], we also represent the limit at infinity via layer potentials, which makes the resulting formulation especially suitable for the numerical computation of eigenvalues. For completeness, we briefly recall known results below.
In this section, differential equations and harmonic functions are again understood in the weak sense, unless stated otherwise.
§4.1.  Bounded domains
On bounded domains the Steklov eigenvalue problem on $\Omega$ can be reformulated in terms of single and double layer potentials; see, for instance, [McL00, p. 244–245], [St07, Section 6.6.3]. In what follows, we express the exterior Steklov problem using the same boundary layer potentials, defined below.
Definition 4.1.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ is a bounded open set with a Lipschitz boundary $\partial\Omega$. Let $\Phi$ be the fundamental solution for the Laplace equation,
[…]
where $|\mathbb{S}^{n-1}|=\frac{2\pi^{n/2}}{\Gamma(n/2)}$ is the measure of the unit sphere in $\mathbb{R}^{n}$.
We define the following layer potentials. Note that, compared to [St07], we might use a different sign convention.
  • The single layer potential $\operatorname{V}:H^{-\frac{1}{2}}(\partial\Omega)\to H^{\frac{1}{2}}(\partial\Omega)$, where $\operatorname{V}\eta:\partial\Omega\to\mathbb{R}$, is given by
    $(\operatorname{V}\eta)(x):=\int_{\partial\Omega}\Phi(x,y)\eta(y)\,\mathrm{d}S_{y}.$
  • The double layer potential $\operatorname{K}:H^{\frac{1}{2}}(\partial\Omega)\to H^{\frac{1}{2}}(\partial\Omega)$, where $\operatorname{K}\eta:\partial\Omega\to\mathbb{R}$, is given by
    $(\operatorname{K}\eta)(x):=\int_{\partial\Omega}(\partial_{\nu_{y}}\Phi)(x,y)\eta(y)\,\mathrm{d}S_{y},$
    and $\nu_{y}$ is the unit normal vector oriented outwards the domain $\Omega$ and
    $\partial_{\nu_{y}}\Phi(x,y)=\left\langle\nu_{y},\nabla_{y}\Phi(x,y)\right\rangle=\frac{\left\langle x-y,\nu_{y}\right\rangle}{|\mathbb{S}^{n-1}||x-y|^{n}}.$
  • The adjoint double layer potential $\operatorname{K}^{\prime}:H^{-\frac{1}{2}}(\partial\Omega)\to H^{-\frac{1}{2}}(\partial\Omega)$, where $\operatorname{K}^{\prime}\eta:\partial\Omega\to\mathbb{R}$, is given by
    $(\operatorname{K}^{\prime}\eta)(x):=\int_{\partial\Omega}(\partial_{\nu_{x}}\Phi)(x,y)\eta(y)\,\mathrm{d}S_{y}.$
  • The hypersingular boundary integral operator $\operatorname{D}:H^{\frac{1}{2}}(\partial\Omega)\to H^{-\frac{1}{2}}(\partial\Omega)$, where $\operatorname{D}\eta:\partial\Omega\to\mathbb{R}$, is given by
    $(\operatorname{D}\eta)(x):=-\partial_{\nu_{x}}\int_{\partial\Omega}(\partial_{\nu_{y}}\Phi)(x,y)\eta(y)\,\mathrm{d}S_{y}.$
Moving the normal derivative inside the integral in $(\operatorname{D}\eta)$ may lead to a non-integrable singularity. Also note that $\operatorname{V}$ and $\operatorname{D}$ are self-adjoint, and $\operatorname{K}$ and $\operatorname{K}^{\prime}$ are formally adjoint on $L^{2}(\partial\Omega)$. ∎
For connected $\Omega$, the classical (interior) Dirichlet-to-Neumann operator $\mathcal{D}:H^{\frac{1}{2}}(\partial\Omega)\to H^{-\frac{1}{2}}(\partial\Omega)$ is considered in [St07, Section 6.6.3] (note that $\sigma(x)$ in [St07] equals $\frac{1}{2}$ almost everywhere since $\partial\Omega$ is Lipschitz). Provided that the single layer operator $\operatorname{V}$ is invertible, $\mathcal{D}$ admits the representation
$\mathcal{D}=\operatorname{V}^{-1}\left(\frac{1}{2}\operatorname{I}-K\right)=\operatorname{D}+\left(\frac{1}{2}\operatorname{I}-\operatorname{K}^{\prime}\right)\operatorname{V}^{-1}\left(\frac{1}{2}\operatorname{I}-\operatorname{K}\right),$
where the second representation has the advantage of being symmetric. Note that the double and adjoint double layer operator defined in [St07] equal $\operatorname{K}$ and $\operatorname{K}^{\prime}$, respectively, because of the orientation of $\nu$. Consider a pair $(u,\tau)\in H^{\frac{1}{2}}(\partial\Omega)\times\mathbb{R}$ satisfying $\mathcal{D}u=\tau u$, or, equivalently
$\frac{1}{2}u-\operatorname{K}u=\tau\operatorname{V}u.$ (4.1)
Then the harmonic extension $U$ of $u$ into $\Omega$, given by
$\begin{split}U:\Omega\to\mathbb{R},\quad U(x)&:=\tau\int_{\partial\Omega}\Phi(x,y)u(y)\,\mathrm{d}S_{y}+\int_{\partial\Omega}(\partial_{\nu_{y}}\Phi)(x,y)u(y)\,\mathrm{d}S_{y}\end{split}$
is an eigenfunction of (1.1) corresponding to the eigenvalue $\tau$.
By Green’s third identity, for $x\in\partial\Omega$,
$\int\limits_{\partial\Omega}(\partial_{\nu_{y}}\Phi)(x,y)\,\mathrm{d}S_{y}=\frac{1}{2}\,$, (4.2)
so that $u\equiv 1$ solves (4.1) with $\tau=0$, as expected.
Finally, let us recall that the single layer operator is always invertible in dimension $n=3$, whereas for $n=2$ we have to assume a suitable scaling of the domain $\Omega$ [St07, Section 6.6]. If $n=2$ and the logarithmic capacity of $\Omega$ is equal to one, the single layer potential has a nontrivial kernel and one cannot use (4.1) to define the eigenvalues of the Dirichlet-to-Neumann operator.
§4.2.  Exterior domains in dimension two
While the boundary integral equation (4.1) remains valid for characterising those eigenfunctions whose harmonic extensions vanish at infinity (up to switching the sign), see §4.3.1, the situation differs in dimension two: as noted in §2.1, exterior Steklov eigenfunctions in two dimensions are only required to remain bounded and do not need to decay. The Dirichlet-to-Neumann operator in exterior domains, including dimension two, together with its Fredholm properties, has been analysed in [KacNaPa25, Section 3]. Similarly, we obtain a representation of the Dirichlet-to-Neumann operator in terms of boundary layer potentials. In contrast to [KacNaPa25], our formulation also incorporates the limit at infinity via layer potentials, yielding a representation that is particularly suitable for numerical computation of eigenvalues.
Let us consider a harmonic function $u\in H^{1}_{\mathrm{loc}}\left(\Omega^{\mathrm{ext}}\right)$ which satisfies the far field boundary condition
$\left|u(x)-u_{\infty}\right|=O\left(\frac{1}{|x|}\right)\qquad\text{ as }|x|\to\infty$ (4.3)
for a given $u_{\infty}\in\mathbb{R}$. According to [St07, Section 7.5] (note that $\gamma_{1}^{\mathrm{ext}}$ in [St07] is $\partial_{\nu}$ with our convention for $\nu$), we obtain, for $x\in\Omega^{\mathrm{ext}}$,
$u(x)=u_{\infty}+\int_{\partial\Omega}\Phi(x,y)(\partial_{\nu_{y}}u)(y)\,\mathrm{d}S_{y}-\int_{\partial\Omega}\left(\partial_{\nu_{y}}\Phi\right)(x,y)u(y)\,\mathrm{d}S_{y}.$
Similar to the interior case, applying the exterior trace operator yields as in [St07, p. 182],
$u(x)=u_{\infty}+(\operatorname{V}(\partial_{\nu}u))(x)+\frac{1}{2}u(x)-(\operatorname{K}u)(x)\quad\text{ for $x\in\partial\Omega$}.$ (4.4)
We aim to express the constant $u_{\infty}$ in terms of single and double layer operators. To this end, we write $u_{\infty}$ in terms of the harmonic measure density. Let $G_{\mathrm{D}}$ be the Dirichlet Green’s function “with pole at infinity” (see, e.g. [GaMa05, p. 41 and p. 74]) for the exterior problem. This means, $G_{\mathrm{D}}$ is the solution of
$\begin{cases}\Delta_{y}G_{\mathrm{D}}(y)=0&\text{ in }\Omega^{\mathrm{ext}},\\ G_{\mathrm{D}}(y)=0&\text{ on }\partial\Omega,\\ G_{\mathrm{D}}(y)-\frac{\log(|y|)}{2\pi}+\frac{\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}=O\left(\frac{1}{|y|}\right)&\text{ as }|y|\to\infty,\end{cases}$
where $\operatorname{Cap}_{\mathrm{log}}(\Omega)$ is the logarithmic capacity of $\Omega$. While [GaMa05] only states o(1){o}(1) decay, any harmonic function vanishing at infinity behaves as $O(1/|y|)$, c.f. Lemma 2.1.
Define the harmonic measure density $\omega:\partial\Omega\to\mathbb{R}$, by
$\omega(y)=-\partial_{\nu_{y}}G_{\mathrm{D}}(y).$
Then, the constant $u_{\infty}$ can be written as
$u_{\infty}=\int_{\partial\Omega}u(y)\omega(y)\,\mathrm{d}S_{y}.$
Thus, equation (4.4) becomes
$\frac{u(x)}{2}+(\operatorname{K}u)(x)-\int_{\partial\Omega}u(y)\omega(y)\,\mathrm{d}S_{y}=(\operatorname{V}(\partial_{\nu}u))(x).$ (4.5)
The main disadvantage of (4.5) is its dependence on the generally unknown harmonic measure density $\omega(y)$. In the following, we express the Dirichlet Green’s function $G_{\mathrm{D}}$ in terms of a single layer potential to overcome this issue.
As in [Ah73, p. 25] or [GaMa05, p. 41], the Dirichlet Green’s function can be expressed as
$G_{\mathrm{D}}(x)=-\Phi(x,0)-g_{0}+g(x),$
where $g_{0}=\frac{\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}$ and $g$ is the harmonic function satisfying
$\Delta g=0\quad\textrm{in}~\Omega^{\mathrm{ext}},\qquad g=\Phi(\cdot,0)+g_{0}\quad\textrm{on}~\partial\Omega,\qquad g(x)=O\left(\frac{1}{|x|}\right)\quad\text{as}~|x|\to{\infty}.$
Then,
$\omega(x)=-\partial_{\nu_{x}}G_{\mathrm{D}}(x)=(\partial_{\nu_{x}}\Phi)(x,0)-(\partial_{\nu_{x}}g)(x)=-\frac{\left\langle x,\nu(x)\right\rangle}{2\pi|x|^{2}}-(\partial_{\nu_{x}}g)(x).$
Since $u$ is harmonic, Green’s identities on $\Omega^{\mathrm{ext}}_{R}$ yield
$\int_{\partial\Omega}u(y)(\partial_{\nu_{y}}g)(y)\,\mathrm{d}S_{y}=\int_{\partial\Omega}g(y)(\partial_{\nu_{y}}u)(y)\,\mathrm{d}S_{y}+\int_{\partial B_{R}}\left(g(y)(\partial_{\nu_{y}}u)(y)-u(y)(\partial_{\nu_{y}}g)(y)\right)\,\mathrm{d}S_{y}$
for any $R>R_{0}(\Omega)$. According to Lemma 2.1, the behaviour of $g$ and $u$ at infinity implies that the last integral vanishes as $R\to\infty$ (note that $\partial_{\nu}u=\partial_{r}u$ on $\partial B_{R}$). As a consequence,
$u_{\infty}=\int_{\partial\Omega}u(y)\omega(y)\,\mathrm{d}S_{y}=-\int_{\partial\Omega}\left(u(y)\frac{\left\langle y,\nu_{y}\right\rangle}{2\pi|y|^{2}}+(\partial_{\nu_{y}}u)(y)g(y)\right)\,\mathrm{d}S_{y}.$
Since $g=\Phi(\cdot,0)+g_{0}$ on $\partial\Omega$, we have
$u_{\infty}=-\int_{\partial\Omega}u(y)\frac{\left\langle y,\nu_{y}\right\rangle}{2\pi|y|^{2}}+(\partial_{\nu_{y}}u)(y)\left(\frac{-\log(|y|)}{2\pi}+\frac{\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}\right)\,\mathrm{d}S_{y}.$ (4.6)
Hence, (4.5) becomes
$\frac{u(x)}{2}+\int_{\partial\Omega}u(y)\frac{\left\langle y,\nu_{y}\right\rangle}{2\pi|y|^{2}}\,\mathrm{d}S_{y}+(\operatorname{K}u)(x)=\int_{\partial\Omega}(\partial_{\nu_{y}}u)(y)\frac{\log(|y|)-\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}\,\mathrm{d}S_{y}+(\operatorname{V}(\partial_{\nu}u))(x).$
To derive a representation of the Dirichlet-to-Neumann map, we define the operator $\operatorname{V}_{0}:H^{-\frac{1}{2}}(\partial\Omega)\to H^{\frac{1}{2}}(\partial\Omega)$, where $\operatorname{V}_{0}\eta:\partial\Omega\to\mathbb{R}$ is given by
$(\operatorname{V}_{0}\eta)(x):=\int_{\partial\Omega}\left(\frac{\log(|y|)}{2\pi}-\frac{\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}+\Phi(x,y)\right)\eta(y)\,\mathrm{d}S_{y}.$
Note that also $\operatorname{V}_{0}u$ is well defined for $u\in H^{1}_{\text{loc}}(\Omega^{\mathrm{ext}})$, because the restriction of $u$ to $\partial\Omega$ is in $H^{\frac{1}{2}}(\partial\Omega)$ and $H^{\frac{1}{2}}(\partial\Omega)\hookrightarrow H^{-\frac{1}{2}}(\partial\Omega)$ continuously. Additionally, we define the functional $\operatorname{K}_{0}:H^{\frac{1}{2}}(\partial\Omega)\to\mathbb{R}$, as
$\operatorname{K}_{0}(\eta):=\int_{\partial\Omega}\eta(y)\frac{\left\langle-y,\nu_{y}\right\rangle}{2\pi|y|^{2}}\,\mathrm{d}S_{y}=(\operatorname{K}\eta)(0).$
If $\operatorname{V}_{0}$ is invertible (see Lemma 4.2), then the Dirichlet-to-Neumann map can be expressed as
$\mathcal{D}^{\mathrm{ext}}:=\operatorname{V}_{0}^{-1}\left(\frac{1}{2}\operatorname{I}-\operatorname{K}_{0}+\operatorname{K}\right):H^{\frac{1}{2}}(\partial\Omega)\to H^{-\frac{1}{2}}(\partial\Omega).$
Accordingly, the boundary condition, $\partial_{\nu}u=\tau u$ on $\partial\Omega$ is equivalent to
$\frac{u}{2}-\operatorname{K}_{0}(u)+(\operatorname{K}u)=\tau(\operatorname{V}_{0}u)\quad\text{ on }\partial\Omega.$ ($\mathrm{LP}_{2}$)
Writing ($\mathrm{LP}_{2}$) as an integral equation, we get
$\begin{split}\frac{u(x)}{2}&+\int_{\partial\Omega}\left(\frac{\left\langle x-y,\nu_{y}\right\rangle}{2\pi|x-y|^{2}}+\frac{\left\langle y,\nu_{y}\right\rangle}{2\pi|y|^{2}}\right)u(y)\,\mathrm{d}S_{y}\\ &=\,\tau\int_{\partial\Omega}\frac{-\log(|x-y|)+\log(|y|)-\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}u(y)\,\mathrm{d}S_{y}.\end{split}$ (4.7)
As a consistency check, consider the case $u=1$. Using the third Green’s identity (4.2) (with reversed orientation) yields $\operatorname{K}u=\frac{1}{2}$. Since the harmonic extension of $u=1$ is constant, we have $u_{\infty}=1$, which makes the left-hand side vanish. One therefore retrieves $\tau_{1}=0$, as it should be.
Finally, it remains to verify the invertibility of $\operatorname{V}_{0}$. The assumption made on the diameter of $\Omega$ can be satisfied by appropriately rescaling the domain.
Lemma 4.2.
Suppose $\Omega\subset\mathbb{R}^{2}$ is a bounded open set with Lipschitz boundary, with connected $\Omega^{\mathrm{ext}}$ and $\operatorname{diam}(\Omega)<1$. Then, $\operatorname{V}_{0}:H^{-\frac{1}{2}}(\partial\Omega)\to H^{\frac{1}{2}}(\partial\Omega)$ is invertible.
Proof.
In [St07, Theorem 6.23], it is shown that $\operatorname{V}:H^{-\frac{1}{2}}(\partial\Omega)\to H^{\frac{1}{2}}(\partial\Omega)$ is invertible if $\operatorname{diam}(\Omega)<1$. Since we build up on this proof, we briefly recall the relevant constructions. Define the subspace
$H_{*}^{-\frac{1}{2}}(\partial\Omega):=\left\{\eta\in H^{-\frac{1}{2}}(\partial\Omega):\left(\eta,1\right)_{\partial\Omega}=0\right\},$
where $\left(\cdot,\cdot\right)_{\partial\Omega}$ means the duality pairing $H^{-\frac{1}{2}}(\partial\Omega)\times H^{\frac{1}{2}}(\partial\Omega)\to\mathbb{R}$. Additionally, let $w_{\mathrm{eq}}\in H^{-\frac{1}{2}}(\partial\Omega)$ be the unique function satisfying
$\operatorname{V}(w_{\mathrm{eq}})(x)=\frac{-\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}\quad\text{ for $x\in\partial\Omega$ }\quad\text{and}\quad\left(w_{\mathrm{eq}},1\right)_{\partial\Omega}=1.$
Any $w\in H^{-\frac{1}{2}}(\partial\Omega)$ admits a unique decomposition
$w=w_{*}+\alpha w_{\mathrm{eq}},\qquad\text{with}\quad w_{*}\in H_{*}^{-\frac{1}{2}}(\partial\Omega),\quad\alpha=\left(w,1\right)_{L^{2}(\partial\Omega)}\in\mathbb{R}.$
The image of $H_{*}^{-\frac{1}{2}}(\partial\Omega)$ under $\operatorname{V}$, is given by
$\left\{v\in H^{\frac{1}{2}}(\partial\Omega):\left(v,w_{\mathrm{eq}}\right)_{\partial\Omega}=0\right\}.$
The preliminaries from [St07] being now in place, we turn to the proof for $\operatorname{V}_{0}$, which can be written as $\operatorname{V}_{0}(\eta)=\operatorname{V}(\eta)+\operatorname{W}(\eta)$, where $\operatorname{W}:H^{-\frac{1}{2}}(\partial\Omega)\to\mathbb{R}$, is defined by
$\operatorname{W}(\eta)=\frac{1}{2\pi}\int_{\partial\Omega}\left(\log(|y|)-\log(\operatorname{Cap}_{\log}(\Omega))\right)\eta(y)\,\mathrm{d}S_{y}=-(\operatorname{V}(\eta)(0)-\frac{\log(\operatorname{Cap}_{\log}(\Omega))}{2\pi}\left(\eta,1\right)_{\partial\Omega}.$
To show that $\operatorname{V}_{0}$ is invertible, we prove its kernel is trivial. Assume $\operatorname{V}_{0}(w)=0$. For $w=w_{*}+\alpha w_{\mathrm{eq}}$,
$\operatorname{V}_{0}(w)=\operatorname{V}(w)+\operatorname{W}(w)=\operatorname{V}(w_{*})+\alpha\operatorname{V}(w_{\mathrm{eq}})+\operatorname{W}(w)=\operatorname{V}(w_{*})+\alpha\frac{-\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}+\operatorname{W}(w).$
Since $\operatorname{W}(w)$ is a constant, $\operatorname{V}(w_{*})$ must be constant. However, $\operatorname{V}(w_{*})$ is orthogonal to $w_{\mathrm{eq}}$ and since $\left(w_{\mathrm{eq}},1\right)_{L^{2}(\partial\Omega)}=1$, it follows that $\operatorname{V}(w_{*})=0$. By the invertibility of $\operatorname{V}$, we conclude $w_{*}=0$. Moreover,
$\operatorname{W}(w_{\mathrm{eq}})=\operatorname{V}(w_{\mathrm{eq}})(0)+\frac{\log(\operatorname{Cap}_{\log}(\Omega))}{2\pi}\left(w_{\mathrm{eq}},1\right)_{\partial\Omega}=-\frac{-\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}-\frac{\log(\operatorname{Cap}_{\log}(\Omega))}{2\pi}=0.$
Thus, if $\operatorname{V}_{0}(\alpha w_{\mathrm{eq}})=0$, then $\operatorname{V}(\alpha w_{\mathrm{eq}})=0$ and therefore $\alpha=0$ because $\operatorname{V}$ is invertible. ∎
To summarise, we present the following
Theorem 4.3.
Suppose $\Omega\subset\mathbb{R}^{2}$ is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and $\operatorname{diam}(\Omega)<1$. The boundary integral equation ($\mathrm{LP}_{2}$) has a sequence of eigenvalues
$0=\tau_{1}\leq\tau_{2}\leq\ldots$
accumulating to $\infty$, with associated eigenfunctions $u_{k}\in H^{\frac{1}{2}}(\partial\Omega)$. For any eigenfunction $u_{k}$, we define $U_{k}:\Omega^{\mathrm{ext}}\to\mathbb{R}$ by
$U_{k}(x):=(u_{k})_{\infty}+\tau_{k}\int_{\partial\Omega}\Phi(x,y)u_{k}(y)\,\mathrm{d}S_{y}-\int_{\partial\Omega}\left(\partial_{\nu_{y}}\Phi\right)(x,y)u_{k}(y)\,\mathrm{d}S_{y},$ (4.8)
where $(u_{k})_{\infty}$ is the limit at infinity, c.f. (4.6). Then, $U_{k}$ is a weak solution of
$\begin{cases}\Delta U_{k}=0\qquad&\text{ in }\Omega^{\mathrm{ext}},\\ \partial_{\nu}U_{k}=\tau_{k}U_{k}\qquad&\text{ on }\partial\Omega,\\ \left|U_{k}(x)-(u_{k})_{\infty}\right|=O(|x|^{-1})\qquad&\text{ as }|x|\to\infty.\end{cases}$
Additionally, if $(U_{k},\tau_{k})$ is a solution of the above boundary-value problem, then $\tau_{k}$ and $U_{k}|_{\partial\Omega}$ satisfy ($\mathrm{LP}_{2}$).
Proof.
The main steps of the proof are given above. However, the asymptotic behaviour of $U_{k}$ may not be immediately apparent. From the definition of $\Phi$, it is clear that the last integral in (4.8) decays. For the first eigenfunction, which is constant, the first integral vanishes. In the case of functions orthogonal to a constant, the decay of the first integral is shown in [St07, Lemma 6.21]. ∎
Remark 4.4.
When restricting to functions with $\left(u,1\right)_{L^{2}(\partial\Omega)}=0$, the constant $\operatorname{Cap}_{\mathrm{log}}(\Omega)$ in (4.7) can be replaced by any other constant without affecting the value of the integral. This could be useful for numerical calculations, as determining the logarithmic capacity, despite the availability of various numerical techniques, introduces an additional computational step. ∎
§4.3.  Exterior domains in dimensions $n\geq 3$
4.3.1 Decay at infinity
As in [KacNaPa25] (see also [Sa15, McL00, St07]), the boundary integral equation (4.1) becomes
$\frac{1}{2}u+\operatorname{K}u=\tau\operatorname{V}u$ ($\mathrm{LP}_{n}$)
for the exterior Steklov problem in $\Omega^{\mathrm{ext}}$ when $n\geq 3$ and we assume decay at infinity. Once an eigenfunction $u$ is found on $\partial\Omega$ with eigenvalue $\tau$ , it can be harmonically extended into $\Omega^{\mathrm{ext}}$ as $U=\mathcal{H}u$, see §1.1. As before, this harmonic extension can also be written as
$U(x)=\tau\int_{\partial\Omega}\Phi(x,y)u(y)\,\mathrm{d}S_{y}-\int_{\partial\Omega}\left(\partial_{\nu_{y}}\Phi\right)(x,y)u(y)\,\mathrm{d}S_{y}.$
Since $\Phi(x,y)=\frac{1}{(n-2)|\mathbb{S}^{n-1}||x-y|^{n-2}}$, it is easy to see that $U(x)$ decays for $|x|\to\infty$.
In summary, we have the following
Theorem 4.5.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. The boundary integral equation ($\mathrm{LP}_{n}$) has a sequence of eigenvalues
$0<\tau_{1}\leq\tau_{2}\leq\ldots,$
accumulating to $\infty$. For any eigenfunction $u_{k}$, we define $U_{k}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ by
$U_{k}(x):=\tau_{k}\int_{\partial\Omega}\Phi(x,y)u_{k}(y)\,\mathrm{d}S_{y}-\int_{\partial\Omega}\left(\partial_{\nu_{y}}\Phi\right)(x,y)u_{k}(y)\,\mathrm{d}S_{y}.$
Then, $U_{k}$ is a weak solution of
[…]
As an immediate consequence, we have the equivalence to the finite energy approach.
Corollary 4.6.
For any bounded open set $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and any $k\in\mathbb{N}$, we have
$\lambda_{k}\left(A^{\mathrm{D}}\right)=\tau_{k}(\Omega^{\mathrm{ext}}).$
Moreover, if $u_{k}$ is an eigenfunction associated to $\lambda_{k}\left(A^{\mathrm{D}}\right)$, then $\left(u_{k},\lambda_{k}\left(A^{\mathrm{D}}\right)\right)$ is a solution of ($\mathrm{LP}_{n}$) and vice versa.
4.3.2 Vanishing flow at infinity
In [HenTW70], Henrici, Troesch and Wuytack investigate the exterior Steklov problem in the upper half space with circular or strip-like aperture in the special case $n=3$. Instead of a decay property, they impose the condition, which can be generalised to any $n\geq 2$ as
$|x|^{n-1}|\nabla u(x)|\to 0\qquad\text{as}\quad|x|\to\infty.$ (4.9)
To establish the well-posedness of the exterior Steklov problem under (4.9), they consider an equivalent boundary integral equation. However, their approach relies on the specific geometry of the domain (a disc or a strip) and is not directly applicable to a general domain $\Omega$. By adapting the calculations from §4.2 to higher dimensions (we only have to adjust Green’s function), we obtain a representation for the corresponding Dirichlet-to-Neumann operator.
Let $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ be a bounded open set with Lipschitz boundary and $u\in H^{1}_{\text{loc}}(\Omega^{\mathrm{ext}})$ a harmonic function satisfying (4.9). First, note that (4.9) implies that $|\nabla u|\in L^{2}\left(\Omega^{\mathrm{ext}}\right)$. Hence, the average of $u$ as defined in (3.1) and denoted by $\langle u\rangle$, is finite. Define $w:=u-\langle u\rangle$. Then, by definition, $w\in W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)$ (see §3.1.1) and therefore, by (3.4), we have $w\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. By Proposition 6.1, $w=O(|x|^{2-n})$ which shows that a higher-dimensional analogue the of far-field condition (4.3) is satisfied with $u_{\infty}:=\langle u\rangle$.
As before, $u_{\infty}$ can be expressed in terms of the Dirichlet Green’s function with pole at infinity,
$\begin{cases}\Delta_{y}G_{\mathrm{D}}(y)=0\qquad&\text{ in }\Omega^{\mathrm{ext}},\\ G_{\mathrm{D}}(y)=0\qquad&\text{ on }\partial\Omega,\\ G_{\mathrm{D}}(y)=a_{0}+\frac{a_{1}}{|y|^{n-2}}+O\left(|y|^{1-n}\right)\qquad&\text{ as }|y|\to\infty.\end{cases}$
The constants $a_{0}$ and $a_{1}$ are determined as follows. The function $P(y):=\frac{a_{0}-G_{\mathrm{D}}(y)}{a_{0}}$ (if $a_{0}=0$, then $G_{\mathrm{D}}\equiv 0$) is the capacitary potential of $\Omega$ and therefore, by definition of the Newtonian capacity,
$P(y)=\frac{\operatorname{Cap}(\Omega)}{(n-2)|\mathbb{S}^{n-1}||y|^{n-2}}+o(|y|^{n-2})\qquad\text{as }|y|\to\infty,$
where $\operatorname{Cap}(\Omega)$ is the Newtonian capacity. This implies $a_{1}=-a_{0}\frac{\operatorname{Cap}(\Omega)}{(n-2)|\mathbb{S}^{n-1}|}$. On the other hand, we normalise $G_{\mathrm{D}}$ by requiring
$\int_{\partial\Omega}(-\partial_{\nu_{y}}G_{\mathrm{D}})(y)\,\mathrm{d}S_{y}=1.$
Using Green’s identities, this implies $1=a_{1}(2-n)|\mathbb{S}^{n-1}|$. Hence, we choose
$a_{0}=\frac{1}{\operatorname{Cap}(\Omega)}\qquad\text{ and }\qquad a_{1}=\frac{1}{|\mathbb{S}^{n-1}|(2-n)}.$
Then, as before,
$u_{\infty}=\int_{\partial\Omega}u(y)\omega(y)\,\mathrm{d}S_{y}=\int_{\partial\Omega}u(y)\left(-\partial_{\nu_{y}}G_{\mathrm{D}}\right)(y)\,\mathrm{d}S_{y}.$
Moreover, the Dirichlet Green’s function can be written as
$G_{\mathrm{D}}(y)=\frac{1}{\operatorname{Cap}(\Omega)}+\frac{1}{|\mathbb{S}^{n-1}|(2-n)|y|^{n-2}}+g(y),$
where $g$ is the harmonic function satisfying
$\begin{cases}\Delta g=0=0\qquad&\text{ in }\Omega^{\mathrm{ext}},\\ g=-\frac{1}{\operatorname{Cap}(\Omega)}-\frac{1}{|\mathbb{S}^{n-1}|(2-n)|y|^{n-2}}\qquad&\text{ on }\partial\Omega,\\ g(y)=O\left(|y|^{1-n}\right)\qquad&\text{ as }|y|\to\infty.\end{cases}$
If $u$ is harmonic, integration by parts yields
$\begin{split}u_{\infty}=\int_{\partial\Omega}u(y)(-\partial_{\nu_{y}}G_{\mathrm{D}})(y)\,\mathrm{d}S_{y}=&\int_{\partial\Omega}-u(y)\frac{\langle y,\nu_{y}\rangle}{|\mathbb{S}^{n-1}||y|^{n}}\,\mathrm{d}S_{y}-\int_{\partial\Omega}g(y)(\partial_{\nu_{y}}u)(y)\,\mathrm{d}S_{y}\\ &-\int_{\partial B_{R}}\left(g(y)(\partial_{\nu_{y}}u)(y)-u(y)(\partial_{\nu_{y}}g)(y)\right)\,\mathrm{d}S_{y}\end{split}$
for any $R>R_{0}(\Omega)$. Again, the last integral vanishes because $u=O(1)$, $\partial_{r}u=O(r^{1-n})$, $g=O(r^{1-n})$ and $\partial_{r}g=O(r^{-n})$ (c.f. the proof of Proposition 6.2). Hence, (4.4) becomes
$\frac{u}{2}=\widehat{\operatorname{V}}_{0}\left(\partial_{\nu}u\right)-\widehat{\operatorname{K}}_{0}u\quad\text{ on $\partial\Omega$},$
where
$\left(\widehat{\operatorname{V}}_{0}\eta\right)(x):=\left({\operatorname{V}}\eta\right)(x)+\int_{\partial\Omega}\eta(y)\left(\frac{1}{\operatorname{Cap}(\Omega)}+\frac{1}{|\mathbb{S}^{n-1}|(2-n)|y|^{n-2}}\right)\,\mathrm{d}S_{y},$
and
$\left(\widehat{\operatorname{K}}_{0}\eta\right)(x):=\left({\operatorname{K}}\eta\right)(x)+\int_{\partial\Omega}\eta(y)\frac{\langle y,\nu_{y}\rangle}{|\mathbb{S}^{n-1}||y|^{n}}\,\mathrm{d}S_{y}.$
As in Lemma 4.2, we can show that $\widehat{\operatorname{V}}_{0}$ is invertible. The only difference is that the logarithmic Green function and the logarithmic equilibrium density $w_{\mathrm{eq}}$ are replaced by their Newtonian counterparts; see, for instance, [St07, formula (6.35)].
Lemma 4.7.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. Then, $\widehat{\operatorname{V}}_{0}:H^{-\frac{1}{2}}(\partial\Omega)\to H^{\frac{1}{2}}(\partial\Omega)$ is invertible.
Thus, we have the following representation for the Dirichlet-to-Neumann operator
$\mathcal{D}:=\widehat{\operatorname{V}}_{0}^{-1}\left(\frac{1}{2}\operatorname{I}+\widehat{\operatorname{K}}_{0}\right):H^{\frac{1}{2}}(\partial\Omega)\to H^{-\frac{1}{2}}(\partial\Omega).$
Proposition 4.8.
Let $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$.
  1. (a)
    Let $u\in H^{-\frac{1}{2}}(\partial\Omega)$ , $\chi\in\mathbb{R}$ solve $\frac{u}{2}+\widehat{\operatorname{K}}_{0}(u)=\chi\widehat{\operatorname{V}}_{0}(u)$, then $U\in H^{1}_{\mathrm{loc}}\left(\Omega^{\mathrm{ext}}\right)$, defined as
    $U(x):=(u)_{\infty}+\chi\int_{\partial\Omega}\Phi(x,y)u(y)\,\mathrm{d}S_{y}-\int_{\partial\Omega}\left(\partial_{\nu_{y}}\Phi\right)(x,y)u(y)\,\mathrm{d}S_{y}$
    solves
    $\begin{cases}\Delta U=0\qquad&\text{in }\Omega^{\mathrm{ext}},\\ \partial_{\nu}U=\chi U\qquad&\text{on }\partial\Omega,\\ |x|^{n-1}|\nabla U(x)|\to 0\qquad&\text{as }|x|\to\infty.\end{cases}$ ($\mathrm{VF}$)
  2. (b)
    Let $\chi\in\mathbb{R}$ and $U\in H^{1}_{\mathrm{loc}}\left(\Omega^{\mathrm{ext}}\right)$ be a weak solution of ($\mathrm{VF}$). Then, $u=U|_{\partial\Omega}\in H^{-\frac{1}{2}}(\partial\Omega)$ solves $\frac{u}{2}+\widehat{\operatorname{K}}_{0}(u)=\chi\widehat{\operatorname{V}}_{0}(u)$.
Since the first eigenvalue vanishes and the first eigenfunction is constant, this formulation is not equivalent to ($\mathrm{ES}$), ($\mathrm{SD}_{R}$), or ($\mathrm{H}_{\Lambda}$), but to ($\mathrm{SN}_{R}$), see §5.1.3.
As in Remark 4.4, when restricting to functions with $\left(u,1\right)_{L^{2}(\partial\Omega)}=0$, the constant in $\hat{\operatorname{V}}_{0}$ can be replaced by any other constant without affecting the value of the integral.
Remark 4.9.
In dimension $n=2$, the vanishing flow condition does not lead to a new problem, since the condition $|x|\cdot|\nabla u(x)|\to 0$ as $|x|\to\infty$ is equivalent to condition (1.2) for harmonic functions, which can easily be seen from Lemma 2.1. ∎
§5.  Equivalence of Different Approaches
§5.1.  Equivalence in dimensions $n\geq 3$
In this section, we prove that the approaches introduced in §3 are equivalent for dimensions $n\geq 3$. The arguments are largely straightforward and rely on approximating the eigenfunctions of one formulation by functions from the corresponding alternative function space.
As before, all differential equations are understood in the weak sense.
5.1.1 Finite-energy functions and Dirichlet truncation
One could use the resolvent convergence of $A_{R}^{\mathrm{D}}$, shown in [AretE15], to infer that the spectral problems ($\mathrm{ES}$) and ($\mathrm{SD}_{R}$) yield the same spectrum as $R\to\infty$. Note that the Neumann formulation ($\mathrm{SN}_{R}$) leads to a different spectrum. We instead proceed directly via the variational characterisation. This approach yields a more direct and unified framework which is also applicable to  ($\mathrm{H}_{\Lambda}$).
We use that smooth, compactly supported functions are dense in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$. Specifically, we use the following approximation property of functions in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$. Let $\psi\in C_{0}^{\infty}(\mathbb{R}^{n})$ be a cut-off function satisfying
[…]
For each $R>0$, define the scaled function $\psi_{R}(x):=\psi\left(\frac{x}{R}\right)$. Since every $u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ has zero average, i.e. $\langle u\rangle=0$, the following lemma is an immediate consequence of [LuOu05, Proposition 2.2]; see also [Bun25, Proof of Theorem 1].
Lemma 5.1.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. For any function $u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ and any $\varepsilon>0$, there exists ${R^{*}}(u,\varepsilon)>0$ such that
$\int_{\Omega^{\mathrm{ext}}}\left|\nabla u-\nabla(\psi_{R}u)\right|^{2}\,\mathrm{d}x<\varepsilon\qquad\text{for all }R>{R^{*}}(u,\varepsilon).$
To establish the convergence of $\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$, we separately prove an upper and a lower bound, both of which converge to $\lambda_{k}\left(A^{\mathrm{D}}\right)$.
We start with the easier inequality. Extending the eigenfunctions corresponding to $\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ by zero yields the following extension of Lemma 3.8.
Lemma 5.2.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. For any $R_{2}>R_{1}>R_{0}(\Omega)$ and any $k\in\mathbb{N}$,
$\lambda_{k}\left(A^{\mathrm{D}}\right)\leq\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R_{1}}\right)\leq\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R_{1}}\right).$
In particular, the limit $\lim_{R\to\infty}\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ exists.
Lemma 5.3.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. For any $\varepsilon>0$ and any $k\in\mathbb{N}$, there exists a constant ${R_{*}}(\varepsilon,k)>0$ such that
$\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{{R_{*}}}\right)\leq\lambda_{k}\left(A^{\mathrm{D}}\right)\left(1+\varepsilon(2+\varepsilon)(1+k)\right).$
Proof.
For each $1\leq m\leq k$, let $f_{m}$ be an eigenfunction associated with $\lambda_{m}\left(A^{\mathrm{D}}\right)$, normalised such that $\int_{\Omega^{\mathrm{ext}}}|\nabla f_{m}|^{2}\,\mathrm{d}x=1$. By Lemma 5.1, for any $\varepsilon>0$, there exists ${R^{*}}>0$ such that the cut-off function satisfies $\operatorname{supp}(\psi_{{R^{*}}}f_{m})\subset B_{2{{R^{*}}}}$, $\psi_{{R^{*}}}(x)=1$ on $\Omega^{\mathrm{ext}}_{{{R^{*}}}}$, and
$\int_{\Omega^{\mathrm{ext}}}\left|\nabla f_{m}-\nabla(\psi_{{R^{*}}}f_{m})\right|^{2}\,\mathrm{d}x<\varepsilon^{2}.$
Note that we can choose a single cut-off function $\psi_{{R^{*}}}$ that works simultaneously for all $m\leq k$. Define $g_{m}:=\psi_{{R^{*}}}f_{m}$ and consider
$M:=\operatorname{Span}\left\{g_{1},\ldots,g_{k}\right\}\subset H^{1}_{D}(\Omega^{\mathrm{ext}}_{2{R^{*}}}).$
Since $g_{m}$ vanishes outside of $\Omega^{\mathrm{ext}}_{2{R^{*}}}$, we use the same notation for the restriction of $g_{m}$ to $\Omega^{\mathrm{ext}}_{2{R^{*}}}$ as for the “original” function. Since $\|\cdot\|_{H^{1}\left(\Omega^{\mathrm{ext}}_{2{R^{*}}}\right)}$ and $\|\cdot\|_{E^{1}\left(\Omega^{\mathrm{ext}}_{2{R^{*}}}\right)}$ are equivalent, [AuHa14b, Theorem A.4], $M$ is $k$-dimensional. Consequently,
$\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}_{2{R^{*}}})\leq\sup_{u\in M,\,u\neq 0}\frac{\int_{\Omega^{\mathrm{ext}}_{2{R^{*}}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$ (5.1)
We now estimate the Rayleigh quotient for each $g_{m}$ using the triangle inequality, the normalisation of $f_{m}$, and that $g_{m}=f_{m}$ on $\partial\Omega$,
$\frac{\int_{\Omega^{\mathrm{ext}}_{2{R^{*}}}}|\nabla g_{m}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|g_{m}|^{2}\,\mathrm{d}S}\leq\frac{\left(\|\nabla(f_{m}-g_{m})\|_{L^{2}\left(\Omega^{\mathrm{ext}}_{2{R^{*}}}\right)}+\|\nabla f_{m}\|_{L^{2}\left(\Omega^{\mathrm{ext}}_{2{R^{*}}}\right)}\right)^{2}}{\int_{\partial\Omega}|f_{m}|^{2}\,\mathrm{d}S}\leq\frac{(\varepsilon+1)^{2}}{\int_{\partial\Omega}|f_{m}|^{2}\,\mathrm{d}S}=(1+\varepsilon)^{2}\lambda_{m}\left(A^{\mathrm{D}}\right).$
Moreover, the functions $f_{1},\ldots,f_{k}$ are orthogonal on $\partial\Omega$, hence
$\int_{\partial\Omega}g_{m}g_{j}\,\mathrm{d}S=\int_{\partial\Omega}f_{m}f_{j}\,\mathrm{d}S=0\qquad\text{ for }m\neq j.$
For the gradients, we estimate using Hölder’s inequality, for $m\neq j$,
$\begin{split}&\left|\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla g_{m},\nabla g_{j}\right\rangle\,\mathrm{d}x\right|\leq\int_{\Omega^{\mathrm{ext}}}|\nabla f_{m}-\nabla g_{m}||\nabla g_{j}|+|\nabla f_{m}||\nabla f_{j}-\nabla g_{j}|\,\mathrm{d}x\\ \leq&\left(\varepsilon^{2}\int_{\Omega^{\mathrm{ext}}}|\nabla g_{j}|^{2}\,\mathrm{d}x\right)^{\frac{1}{2}}+\left(\varepsilon^{2}\int_{\Omega^{\mathrm{ext}}}|\nabla f_{m}|^{2}\,\mathrm{d}x\right)^{\frac{1}{2}}\leq\varepsilon(1+\varepsilon)+\varepsilon.\end{split}$
Now, inserting $u=\sum_{m=1}^{k}c_{m}g_{m}$ into (5.1),
$\begin{split}\frac{\int_{\Omega^{\mathrm{ext}}_{2{R^{*}}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}u^{2}\,\mathrm{d}S}&=\frac{\int_{\Omega^{\mathrm{ext}}_{2{R^{*}}}}\left(\sum_{m=1}^{k}c_{m}^{2}|\nabla g_{m}|^{2}+\sum_{j,m=1,\,j\neq m}^{k}c_{m}c_{j}\left\langle\nabla g_{j},\nabla g_{m}\right\rangle\right)\,\mathrm{d}x}{\int_{\partial\Omega}\left(\sum_{m=1}^{k}c_{m}^{2}g_{m}^{2}+\sum_{j,m=1,\,j\neq m}^{k}c_{m}c_{j}g_{j}g_{m}\right)\,\mathrm{d}S}\\ &\leq\frac{(1+\varepsilon)^{2}\sum_{m=1}^{k}c_{m}^{2}+\varepsilon(\varepsilon+2)\sum_{j,m=1,\,j\neq m}^{k}|c_{m}c_{j}|}{\sum_{m=1}^{k}\frac{c_{m}^{2}}{\lambda_{m}\left(A^{\mathrm{D}}\right)}}\leq\lambda_{k}\left(A^{\mathrm{D}}\right)\left(1+\varepsilon(2+\varepsilon)(1+k)\right),\end{split}$
where in the last step we used the Cauchy—Schwarz inequality, and the fact that $\lambda_{m}\left(A^{\mathrm{D}}\right)\leq\lambda_{k}\left(A^{\mathrm{D}}\right)$ for $1\leq m\leq k$. Hence, with $R_{*}=2R^{*}$, this completes the proof. ∎
Combining Lemma 5.2 and Lemma 5.3, we obtain the following convergence result.
Theorem 5.4.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. Then, for every $k\in\mathbb{N}$,
$\lim_{R\to\infty}\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)=\lambda_{k}\left(A^{\mathrm{D}}\right).$
For the Dirichlet-type problem, convergence of eigenfunctions holds in a suitable sense. The proof follows a similar strategy to that used for the eigenvalues. Note that convergence cannot be expected for an arbitrary sequence of eigenfunctions. For instance, if $\Omega=B_{1}\subset\mathbb{R}^{3}$, then $\sigma_{2}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ has multiplicity three, so a sequence of eigenfunctions $\left(u_{k,R_{m}}\right)_{m\in\mathbb{N}}$ associated with $\sigma_{2}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R_{m}}\right)$ could ”jump” between different eigenmodes as $m$ varies.
Corollary 5.5.
Suppose $f_{k}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ is an eigenfunction associated with $\lambda_{k}\left(A^{\mathrm{D}}\right)$. Then, there exists a sequence $\left(R_{m}\right)_{m\in\mathbb{N}}$ with $\lim_{m\to\infty}R_{m}=\infty$ and a sequence $\left(g_{k,R_{m}}\right)_{m\in\mathbb{N}}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)$, such that
  • $g_{k,R_{m}}(x)=0$ for $|x|>R_{m}$ and $g_{k,R_{m}}|_{\Omega^{\mathrm{ext}}_{R_{m}}}$ is an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}_{R_{m}})$,
  • $\lim_{m\to\infty}\|g_{k,R_{m}}-f_{k}\|_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}=0$.
Proof.
We assume, without loss of generality, that $\int_{\Omega^{\mathrm{ext}}}|\nabla f_{k}|^{2}\,\mathrm{d}x=1$. For any $R>R_{0}(\Omega)$, let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ denote an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ (respectively its extension by zero outside $\Omega^{\mathrm{ext}}_{R}$), normalised to
$\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u_{k,R}|^{2}\,\mathrm{d}x=\int_{\Omega^{\mathrm{ext}}}|\nabla{u_{k,R}}|^{2}\,\mathrm{d}x=1.$
By Lemma 5.3,
$\lim_{R\to\infty}\int_{\partial\Omega}|u_{k,R}|^{2}\,\mathrm{d}S=\lim_{R\to\infty}\frac{1}{\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)}=\frac{1}{\lambda_{k}\left(A^{\mathrm{D}}\right)}=\int_{\partial\Omega}|f_{k}|^{2}\,\mathrm{d}S.$
Hence, the norm $\|{u_{k,R}}\|_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}$ is bounded. Thus, there exists a weakly convergent subsequence in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$, which we again denote by $u_{k,R}$, with weak limit $\widehat{u}_{k}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. The compactness of the trace operator $\boldsymbol{\tau}:E^{1}(\Omega)\to L^{2}(\partial\Omega)$ (see [AuHa14b, Corollary 3.4]) and the convergence of the eigenvalues imply that the convergence of $u_{k,R}$} is strong. Thus, $\widehat{u}_{k}$ is a Steklov eigenfunction with associated eigenvalue $\lambda_{k}\left(A^{\mathrm{D}}\right)$.
In the second part of the proof, we show that every eigenfunction of $\lambda_{k}\left(A^{\mathrm{D}}\right)$ arises as the limit of such a sequence.
Using a diagonal argument, we may choose a sequence $(R_{m})_{m\in\mathbb{N}}$ such that each $u_{j,R_{m}}$ converges in $E^{1}(\Omega^{\mathrm{ext}})$ for $1\leq j\leq k$. Specifically, we first extract a subsequence along which $u_{1,R}$ converges; from this subsequence we then extract one along which $u_{2,R}$ converges, and we continue this process up to $u_{k,R}$. Let $h_{j}\in E^{1}(\Omega^{\mathrm{ext}})$ denote the corresponding limits.
We now claim that the limits $h_{1},\ldots,h_{k}$ are linearly independent. Suppose for contradiction that $\dim\left(\operatorname{Span}\left\{h_{1},\ldots,h_{k}\right\}\right)\leq k-1$. Then there exists an eigenfunction $\widehat{h}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ associated with $\lambda_{\ell}\left(A^{\mathrm{D}}\right)$, $\ell\leq k$ that is linearly independent of $h_{1},\ldots,h_{k}$. As in the proof of Lemma 5.3, we approximate $\widehat{h}$ by a smooth function $\mathcal{U}$ with $\operatorname{supp}(\mathcal{U})\subset B_{R_{m}}$ and consider
$M:=\operatorname{Span}\left\{u_{1,R_{m}},\ldots,u_{k,R_{m}},\mathcal{U}\right\}.$
For sufficiently large $m$, we have $\dim(M)=k+1$. Hence, using the variational characterisation,
$\sigma_{k+1}^{\mathrm{D}}(\Omega^{\mathrm{ext}}_{R_{m}})\leq\sup_{u\in M,\,u\neq 0}\frac{\int_{\Omega^{\mathrm{ext}}_{R_{m}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$
As in the proof of Lemma 5.3, the right-hand side converges to $\lambda_{k}\left(A^{\mathrm{D}}\right)$ as $m\to\infty$. If we consider a $k$ where $\lambda_{k}\left(A^{\mathrm{D}}\right)\neq\lambda_{k+1}\left(A^{\mathrm{D}}\right)$, this forms a contradiction. Consequently, for any eigenfunction $f_{q}$ associated with $\lambda_{q}\left(A^{\mathrm{D}}\right)$, we can set $k=\max\left\{p\in\mathbb{N}:\lambda_{p}\left(A^{\mathrm{D}}\right)\leq\lambda_{q}\left(A^{\mathrm{D}}\right)\right\}$ and obtain the existence of an $r\leq k$ with $h_{r}=f_{q}$. By definition, the sequence $u_{r,R_{m}}$} satisfies the required properties. ∎
5.1.2 Finite-energy functions and Helmholtz equation
In this section, similarly to §5.1.1, we show that as $\Lambda\searrow 0$, the eigenvalues $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$ and their associated eigenfunctions converge to $\lambda_{k}\left(A^{\mathrm{D}}\right)$ and their corresponding eigenfunctions respectively.
Since $H^{1}\left(\Omega^{\mathrm{ext}}\right)\subset E^{1}\left(\Omega^{\mathrm{ext}}\right)$, the variational characterisations (3.5) and (3.8) immediately imply
$\lambda_{k}\left(A^{\mathrm{D}}\right)<\mu_{k}(\Omega^{\mathrm{ext}},\Lambda_{1})<\mu_{k}(\Omega^{\mathrm{ext}},\Lambda_{2}).$
for any $k\in\mathbb{N}$ and any $0<\Lambda_{1}<\Lambda_{2}$. Thus, to conclude the convergence of $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$, it suffices to establish an upper bound which converges to $\lambda_{k}\left(A^{\mathrm{D}}\right)$ as $\Lambda\searrow 0$.
Theorem 5.6.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and let $k\in\mathbb{N}$. Then,
$\lim_{\Lambda\searrow 0}\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)=\lambda_{k}\left(A^{\mathrm{D}}\right).$
Proof.
For $1\leq m\leq k$, let $f_{m}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ be the eigenfunction associated with $\lambda_{m}\left(A^{\mathrm{D}}\right)$ normalised to $\int_{\Omega^{\mathrm{ext}}}|\nabla f_{m}|^{2}\,\mathrm{d}x=1$. By Lemma 5.1, for any $\varepsilon>0$, there exist a constant ${R^{*}}(\varepsilon,m)>0$ and functions $g_{m}$ with $\operatorname{supp}(g_{m})\subset B_{2{R^{*}}}$ such that $g_{m}(x)=f_{m}(x)$ in $B_{{R^{*}}}\setminus{\Omega}$ and
$\int_{\Omega^{\mathrm{ext}}}|\nabla f_{m}-\nabla g_{m}|^{2}\,\mathrm{d}x<\varepsilon^{2}.$
Define the $k$-dimensional subspace $M:=\operatorname{Span}\left\{g_{1},\ldots,g_{k}\right\}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)$. Using the variational characterisation of $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$, we obtain
$\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)\leq\sup_{\begin{subarray}{c}u\in M\\ u\neq 0\end{subarray}}\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u|^{2}\,\mathrm{d}x+\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$
We now show that the right-hand side tends to $\lambda_{k}\left(A^{\mathrm{D}}\right)$ as $\Lambda\searrow 0$. Since $\|\nabla f_{m}-\nabla g_{m}\|_{L^{2}\left(\Omega^{\mathrm{ext}}\right)}<\varepsilon$ and $g_{m}=f_{m}$ on $\partial\Omega$, we estimate
$\begin{split}\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|g_{m}|^{2}\,\mathrm{d}x+\int_{\Omega^{\mathrm{ext}}}|\nabla g_{m}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|g_{m}|^{2}\,\mathrm{d}S}&\leq\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|g_{m}|^{2}\,\mathrm{d}x+\left(\|\nabla f_{m}\|_{L^{2}\left(\Omega^{\mathrm{ext}}\right)}+\varepsilon\right)^{2}}{\int_{\partial\Omega}|f_{m}|^{2}\,\mathrm{d}S}\\ &=\lambda_{m}\left(A^{\mathrm{D}}\right)\left(\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|g_{m}|^{2}\,\mathrm{d}x+\left(1+\varepsilon\right)^{2}\right).\end{split}$
Since the $g_{m}$ are independent of $\Lambda$, it follows that
$\lim_{\Lambda\searrow 0}\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|g_{m}|^{2}\,\mathrm{d}x+\int_{\Omega^{\mathrm{ext}}}|\nabla g_{m}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|g_{m}|^{2}\,\mathrm{d}S}\leq\lambda_{m}\left(A^{\mathrm{D}}\right)(1+\varepsilon)^{2}.$
As in the proof of Lemma 5.3,
$\int_{\partial\Omega}g_{i}g_{j}\,\mathrm{d}S=0\quad\text{ and }\quad\left|\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla g_{i},\nabla g_{j}\right\rangle\,\mathrm{d}x\right|<3\varepsilon\qquad\text{for}\,i\neq j.$
Now let $u=\sum_{m=1}^{k}c_{m}g_{m}\in M$. Then, as $\Lambda\searrow 0$,
$\begin{split}&\lim_{\Lambda\searrow 0}\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u|^{2}\,\mathrm{d}x+\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}u^{2}\,\mathrm{d}S}=\frac{\sum_{m,j=1}^{k}c_{m}c_{j}\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla g_{m},\nabla g_{j}\right\rangle\,\mathrm{d}S}{\sum_{m,j=1}^{k}c_{m}c_{j}\int_{\partial\Omega}g_{m}g_{j}\,\mathrm{d}S}\\ \leq&\frac{\sum_{m=1}^{k}c_{m}^{2}(1+\varepsilon)^{2}+3\varepsilon\sum_{\begin{subarray}{c}m,j=1,\,m\neq j\end{subarray}}^{k}|c_{m}c_{j}|}{\sum_{m=1}^{k}c_{m}^{2}\lambda_{m}\left(A^{\mathrm{D}}\right)^{-1}}\leq\lambda_{k}\left(A^{\mathrm{D}}\right)\left((1+\varepsilon)^{2}+3k\varepsilon\right).\end{split}$
Since $\varepsilon>0$ was arbitrary, the claim follows. ∎
Using the convergence of $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$, we can now establish the convergence of the associated eigenfunctions, similar to Corollary 5.5.
Corollary 5.7.
Suppose $f_{k}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ is an eigenfunction associated with $\lambda_{k}\left(A^{\mathrm{D}}\right)$. Then, there exist a sequence $\left(\Lambda_{m}\right)_{m\in\mathbb{N}}\subset(0,\infty)$ with $\lim_{m\to\infty}\Lambda_{m}=0$ and a sequence $\left(g_{k,\Lambda_{m}}\right)_{m\in\mathbb{N}}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)$, such that
  • $g_{k,\Lambda_{m}}$ is an eigenfunction associated with $\mu_{k}(\Omega^{\mathrm{ext}},\Lambda_{m})$ and $\lim_{m\to\infty}\mu_{k}(\Omega^{\mathrm{ext}},\Lambda_{m})=\lambda_{k}\left(A^{\mathrm{D}}\right)$,
  • $\lim_{m\to\infty}\|g_{k,\Lambda_{m}}-f_{k}\|_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}=0$.
Proof.
Let $u_{k,\Lambda}\in H^{1}\left(\Omega^{\mathrm{ext}}\right)$ denote an eigenfunction associated with $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$, and we may assume without restriction
$\int_{\Omega^{\mathrm{ext}}}|\nabla f_{k}|^{2}\,\mathrm{d}x=1\quad\text{ and }\quad\int_{\Omega^{\mathrm{ext}}}|\nabla u_{k,\Lambda}|^{2}\,\mathrm{d}x=1\quad\text{for all}\quad\Lambda>0.$
We first prove that the sequence $u_{k,\Lambda}$ converges strongly in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$}. Observe that
$\frac{1}{\int_{\partial\Omega}|u_{k,\Lambda}|^{2}\,\mathrm{d}S}\leq\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u_{k,\Lambda}|^{2}\,\mathrm{d}x+1}{\int_{\partial\Omega}|u_{k,\Lambda}|^{2}\,\mathrm{d}S}=\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)\xrightarrow{\Lambda\searrow 0}\lambda_{k}\left(A^{\mathrm{D}}\right).$ (5.2)
On the other hand, since $H^{1}\left(\Omega^{\mathrm{ext}}\right)\subset E^{1}\left(\Omega^{\mathrm{ext}}\right)$, the space
$M:=\operatorname{Span}\left\{u_{1,\Lambda},\ldots,u_{k,\Lambda}\right\}\subset E^{1}\left(\Omega^{\mathrm{ext}}\right)$
has dimension $k$. Hence, the min-max principle for $\lambda_{k}\left(A^{\mathrm{D}}\right)$ yields
$\lambda_{k}\left(A^{\mathrm{D}}\right)\leq\sup_{\begin{subarray}{c}u\in M\\ u\neq 0\end{subarray}}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$ (5.3)
In the case $k=1$, inequality (5.3) immediately yields $\int_{\partial\Omega}|u_{1,\Lambda}|^{2}\,\mathrm{d}S\leq\frac{1}{\lambda_{1}\left(A^{\mathrm{D}}\right)}$. Combining this with (5.2), we conclude
$\lim_{\Lambda\searrow 0}\int_{\partial\Omega}|u_{1,\Lambda}|^{2}\,\mathrm{d}S=\frac{1}{\lambda_{1}\left(A^{\mathrm{D}}\right)}.$
Since $\|u_{1,\Lambda}\|_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}$ is bounded, there exists a weakly convergent subsequence with limit $v\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. The compactness of the trace operator implies that $\int_{\partial\Omega}|v|^{2}\,\mathrm{d}S=\lambda_{1}\left(A^{\mathrm{D}}\right)^{-1}$, and by lower semicontinuity, we have $\int_{\Omega^{\mathrm{ext}}}|\nabla v|^{2}\,\mathrm{d}x=1$ (if $\int_{\Omega^{\mathrm{ext}}}|\nabla v|^{2}\,\mathrm{d}x<1$, we would have $\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla v|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|v|^{2}\,\mathrm{d}S}<\lambda_{1}\left(A^{\mathrm{D}}\right)$ which is not possible). In particular, the limit has to be an eigenfunction associated with $\lambda_{1}\left(A^{\mathrm{D}}\right)$, i.e $v=f_{1}$. Since $\lambda_{1}\left(A^{\mathrm{D}}\right)$ is simple and all $u_{1,\Lambda}$ are normalised, the whole sequence converges in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$ to $f_{1}$.
We now proceed by induction on $k$. Let $m(k)$ denote the multiplicity of $\lambda_{k}\left(A^{\mathrm{D}}\right)$, and define $M(k):=\sum_{j=1}^{k}m(k)$. Then
$\lambda_{M(k)}\left(A^{\mathrm{D}}\right)<\lambda_{M(k)+1}\left(A^{\mathrm{D}}\right)=\lambda_{M(k)+2}\left(A^{\mathrm{D}}\right)=\ldots=\lambda_{M(k+1)}\left(A^{\mathrm{D}}\right).$
Assume that for a given $k\in\mathbb{N}$, there exists a sequence $\Lambda_{m}$ with $\Lambda_{m}\to 0$, such that for all $1\leq j\leq M(k)$,
$\lim_{m\to\infty}\|u_{j,\Lambda_{m}}-f_{j}\|_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}=0.$
For each $M(k)<j\leq M(k+1)$, we will show that $(u_{j,\Lambda_{m}})_{m}$ has a convergent subsequence. Consider the space
$M_{j}:=\operatorname{Span}\left\{u_{1,\Lambda_{m}},\ldots,u_{M(k),\Lambda_{m}},u_{j,\Lambda_{m}}\right\},$
which has dimension $M(k)+1$. Since $u_{j,\Lambda_{m}}$ is an eigenfunction associated with a different eigenvalue than $u_{1,\Lambda_{m}},\ldots,u_{M(k),\Lambda_{m}}$, it is orthogonal to any $u_{1,\Lambda_{m}},\ldots,u_{M(k),\Lambda_{m}}$ in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$ and in $L^{2}(\partial\Omega)$. Thus, analogously to (5.3), we obtain
$\lambda_{M(k)+1}\left(A^{\mathrm{D}}\right)\leq\sup_{\begin{subarray}{c}u\in M_{j}\\ u\neq 0\end{subarray}}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}\leq\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u_{j,\Lambda_{m}}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u_{j,\Lambda_{m}}|^{2}\,\mathrm{d}S}=\frac{1}{\int_{\partial\Omega}|u_{j,\Lambda_{m}}|^{2}\,\mathrm{d}S}.$
Combined with (5.2), we again obtain boundedness of $\|u_{j,\Lambda_{m}}\|_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}$ and hence existence of a convergent subsequence where the limit, denoted by $h_{j}$, is an eigenfunction associated with $\lambda_{M(k+1)}\left(A^{\mathrm{D}}\right)$. By passing to a diagonal subsequence, i.e.we first extract a subsequence along which $u_{1,\Lambda_{m}}$ converges; from this subsequence we then extract one along which $u_{2,\Lambda_{m}}$ converges and so on, we may assume that $u_{j,\Lambda_{m}}\to h_{j}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ for all $j\leq M(k+1)$.
We now show that the limits $h_{1},\dots,h_{M(k+1)}$ are linearly independent. Assume to the contrary that they are linearly dependent. Then, there exists $\widehat{h}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ which is an eigenfunction associated with $\sigma_{j}\left(\Omega^{\mathrm{ext}}\right)$, $j\leq M(k+1)$ such that
$\widehat{h}\notin\operatorname{Span}\left\{h_{i}\mid 1\leq i\leq M(k+1)\right\}.$
Choose $\mathcal{U}\in H^{1}\left(\Omega^{\mathrm{ext}}\right)$ with $\|\nabla\widehat{h}-\nabla\mathcal{U}\|_{L^{2}\left(\Omega^{\mathrm{ext}}\right)}<\varepsilon$ (as in Lemma 5.5). Then for $m$ large, the space
$M:=\operatorname{Span}\left\{u_{1,\Lambda_{m}},\ldots,u_{M(k+1),\Lambda_{m}},\mathcal{U}\right\}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)$
has dimension $M(k+1)+1$. Therefore,
$\mu_{M(k+1)+1}\left(\Omega^{\mathrm{ext}},\Lambda\right)\leq\inf_{u\in M}\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u|^{2}\,\mathrm{d}x+\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$
Analogously to the proof of Theorem 5.6, the right-hand side tends to $\lambda_{M(k+1)}\left(A^{\mathrm{D}}\right)$ as $m\to\infty$. However, $\lim_{m\to\infty}\mu_{M(k+1)+1}\left(\Omega^{\mathrm{ext}},\Lambda\right)=\lambda_{M(k+1)+1}\left(A^{\mathrm{D}}\right)>\lambda_{M(k+1)}\left(A^{\mathrm{D}}\right)$, which forms a contradiction. Hence, $h_{1},\ldots,h_{M(k+1)}$ must be linearly independent. In particular, for any eigenfunction $f$ associated with $\lambda_{k}\left(A^{\mathrm{D}}\right)$, there exists $r$ such that $f=h_{r}$ and $u_{r,\Lambda_{m}}\to f$ in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$. ∎
In summary, the four different available approaches to the exterior problem, introduced in §3.1, §3.2, §3.3, and §4.3.1, are equivalent in dimensions $n\geq 3$.
5.1.3 Vanishing flow and Neumann truncation
In §5.1.1 we have seen that the finite energy approach is equivalent to approximating the exterior domain by truncated domains equipped with Dirichlet boundary conditions on the outer boundary. By contrast, if Neumann boundary condition is imposed on the truncation boundary, one obtains a different limiting spectrum in dimensions $n\geq 3$, see Remark 1.6. In dimension $n=2$, however, the Dirichlet and Neumann truncations lead to the same spectrum in the limit; this will be shown in §5.2.1.
In §4.3.2 we introduced, via layer potentials, an alternative formulation of the exterior Steklov problem. In this setting, functions are not required to vanish at infinity, but their gradients must decay faster than $|x|^{1-n}$. We now prove that this formulation is equivalent to the Neumann truncation.
Theorem 5.8.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and let $k\in\mathbb{N}$. Denote by $\chi_{k}\left(\Omega^{\mathrm{ext}}\right)$ the eigenvalues of ($\mathrm{VF}$), and recall that $\sigma^{\mathrm{N}}_{k}\left(\Omega^{\mathrm{ext}}_{R}\right)$ are the eigenvalues of ($\mathrm{SN}_{R}$) and. Then
$\lim_{R\to\infty}\sigma^{\mathrm{N}}_{k}\left(\Omega^{\mathrm{ext}}_{R}\right)=\chi_{k}\left(\Omega^{\mathrm{ext}}\right).$
Proof.
By Theorem 3.9 (see also [AretE15, Theorem 5.6]), the resolvent of the Dirichlet-to-Neumann operator with Neumann condition on $\partial B_{R}$ converges uniformly to the resolvent of $A^{\mathrm{N}}$ from Definition 3.1. Thus it suffices to show the equivalence between $A^{\mathrm{N}}$ and ($\mathrm{VF}$).
Let $\left(\sigma^{\mathrm{N}},u^{\mathrm{N}}\right)\in\mathbb{R}\times W\left(\Omega^{\mathrm{ext}}\right)$ be an eigenpair of $A^{\mathrm{N}}$. Let $\langle u\rangle$ denote the average of $u$, as defined in (3.1). Then, by definition, $w:=u^{\mathrm{N}}-\left\langle u^{\mathrm{N}}\right\rangle$ is in $W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)$, and therefore, by (3.4), $w\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. Without loss of generality, assume $\Omega\subset B_{1}$. As in the proof of Proposition 6.2, which will be presented later, we expand $w$ in spherical harmonics
$w(x)=\sum_{k=0}^{\infty}\sum_{i=1}^{d_{n,k}}\left(w,{u}_{(k,i)}\right)_{L^{2}\left(\partial B_{1}\right)}{u}_{(k,i)}(x),$
where ${u}_{(k,i)}\in E^{1}\left(B_{1}^{\mathrm{ext}}\right)$, given by ${u}_{(k,i)}(r,\theta)=r^{2-n-k}Y_{k,i}(\theta)$, are Steklov eigenfunctions in the exterior of the unit ball forming an orthonormal basis of $L^{2}\left(\partial B_{1}\right)$, see §2.2. This expansion converges uniformly in $\Omega^{\mathrm{ext}}$, see the proof of Proposition 6.2. From the explicit form of ${u}_{(k,i)}$ it follows that $w$ (and thus $u^{\mathrm{N}}$) satisfies the vanishing flow condition (4.9) if and only if $\left(w,{u}_{(0,1)}\right)_{L^{2}\left(\partial B_{1}\right)}=0$.
To establish this, note first that since $w$ is harmonic and $\partial_{\nu}w=\partial_{\nu}u^{\mathrm{N}}=\sigma^{\mathrm{N}}u^{\mathrm{N}}$ on $\partial\Omega$, we have
$0=\int_{\Omega^{\mathrm{ext}}\cap B_{1}}\Delta w\,\mathrm{d}x=\int_{\partial\Omega}\partial_{\nu}w\,\mathrm{d}S+\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S=\sigma^{\mathrm{N}}\int_{\partial\Omega}u^{\mathrm{N}}\,\mathrm{d}S+\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S=\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S.$
In the last step we used that the first eigenvalue vanishes and the associated eigenfunction is a constant and therefore, any other eigenfunction is orthogonal to a constant. Furthermore, $\partial_{\nu}$ coincides with the radial derivative on $\partial B_{1}$. Hence,
$0=\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S=\sum_{k=0}^{\infty}\sum_{i=1}^{d_{n,k}}\left(w,{u}_{(k,i)}\right)_{L^{2}\left(\partial B_{1}\right)}(2-n-k)\int_{\partial B_{1}}Y_{k,i}(\theta)\,\mathrm{d}S.$
Since the spherical harmonics of degree $k\neq 0$ integrate to zero, only the $k=0$ term remains. Therefore $0=\left(w,{u}_{0,1}\right)_{L^{2}\left(\partial B_{1}\right)}c_{0,1}(2-n)\int_{\partial B_{1}}Y_{0,1}(\theta)\,\mathrm{d}S$. Hence, $\left(w,{u}_{0,1}\right)_{L^{2}\left(\partial B_{1}\right)}=0$. Thus $u^{\mathrm{N}}$ satisfies the vanishing flow condition (4.9), and $\left(\sigma^{\mathrm{N}},u^{\mathrm{N}}\right)$ is an eigenpair of ($\mathrm{VF}$).
Conversely, if $(\chi,u)$ is an eigenpair of ($\mathrm{VF}$), then $u\in W\left(\Omega^{\mathrm{ext}}\right)$ by definition, hence it is also an eigenpair of $A^{\mathrm{N}}$. ∎
§5.2.  Equivalence in dimension two
We show that formulation ($\mathrm{CT}$) is equivalent to the approach via the Helmholtz equation, the truncated domain method and the approach via layer potentials. Recall that the approach of Auchmuty and Han does not extend to two dimensions, and Arendt and ter Elst did not study that case either, see Remark 3.6.
Moreover, the conformal mapping framework leads to several results analogous to those known for bounded domains.
In the following, when discussing ($\mathrm{CT}$), we switch from the complex formulation to real variables. We now view $\Omega^{*}$ as a subset of $\mathbb{R}^{2}$ and represent points as $x=(x_{1},x_{2})$ instead of $z=x_{1}+\mathrm{i}x_{2}$, so that $\phi:\left(x_{1},x_{2}\right)\mapsto\left(\frac{x_{1}}{|x|^{2}},-\frac{x_{2}}{|x|^{2}}\right)$, and $\left|\phi^{\prime}(x)\right|=\frac{1}{|x|^{2}}$. Accordingly, the derivatives and normal vectors are interpreted in the standard real sense.
5.2.1 Conformal mapping and truncated domains
In contrast to higher dimensions, the choice of boundary condition at the outer boundary of the truncated domain plays no role in two dimensions when $R\to\infty$. Loosely speaking, this is because under $\phi$, defined in (3.10), the outer boundary $\partial B_{R}$ becomes a small hole in $\Omega^{*}$ whose influence vanishes as $R\to\infty$.
Theorem 5.9.
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and such that $\Omega$ contains the origin. Then, for any $k\in\mathbb{N}$,
$\lim_{R\to\infty}\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)=\xi_{k}\left(\Omega^{\mathrm{ext}}\right),$
where $\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ is defined as in §3.2.
Proof.
Let $u_{k,R}\in H^{1}_{D}\left(\Omega^{\mathrm{ext}}_{R}\right)$ denote an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$. By slightly abusing the notation, $u_{k,R}$ also denotes the extension to $\Omega^{\mathrm{ext}}$ by zero. In addition, we define $f_{k,R}\in H^{1}(\Omega^{*})$ by
$f_{k,R}:\Omega^{*}\to\mathbb{R},\qquad f_{k,R}(x)=\begin{cases}u_{k,R}\left(\phi(x)\right)\qquad&\text{ if }|x|\geq\frac{1}{R},\\ 0\qquad&\text{ if }|x|<\frac{1}{R}.\end{cases}$
Since $\phi$ is a conformal map,
$\frac{\int_{\Omega^{*}}|\nabla f_{k,R}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|f_{k,R}|^{2}\,\mathrm{d}S}=\frac{\int_{\Omega^{*}\setminus\overline{B_{1/R}}}|\nabla f_{k,R}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|f_{k,R}|^{2}\,\mathrm{d}S}=\frac{\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u_{k,R}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u_{k,R}|^{2}\,\mathrm{d}S}.$
For fixed $R>R_{0}(\Omega)$, define the $j$-dimensional subspaces
$M(j):=\operatorname{Span}\left\{f_{k,R}\mid 1\leq k\leq j\right\}\subset H^{1}(\Omega^{*})\quad\text{ and }\quad N(j):=\operatorname{Span}\left\{u_{k,R}\mid 1\leq k\leq j\right\}\subset H^{1}_{D}\left(\Omega^{\mathrm{ext}}_{R}\right).$
Then,
$\xi_{j}\left(\Omega^{\mathrm{ext}}\right)\leq\sup_{0\neq u\in M(j)}\frac{\int_{\Omega^{*}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|u|^{2}\,\mathrm{d}S}=\sup_{0\neq u\in N(j)}\frac{\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}=\sigma_{j}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right).$
For the reverse inequality, let $f_{k}\in H^{1}(\Omega^{*})$ denote eigenfunctions associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$, and define
$g_{k,R}:\Omega^{\mathrm{ext}}\to\mathbb{R},\qquad g_{k,R}(x)=\begin{cases}f_{k}\left(\phi^{-1}(x)\right)\qquad&\text{ if }x\in\Omega^{\mathrm{ext}}_{R},\\ f_{k}\left(\phi^{-1}(x)\right)\left(\frac{\log(R^{2})-\log(|x|)}{\log(R^{2})-\log(R)}\right)\qquad&\text{ if }R<|x|\leq R^{2},\\ 0\qquad&\text{ if }|x|>R^{2}.\end{cases}$
Since $|\nabla f_{k}|\in L^{2}(\Omega^{*})$, a straightforward calculation shows $\lim_{R\to\infty}\int_{B_{R^{2}}\setminus\overline{B_{R}}}|\nabla g_{k,R}|^{2}\,\mathrm{d}x=0$. Furthermore,
$\int_{\partial\Omega}|g_{k,R}|^{2}\,\mathrm{d}S=\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|f_{k}|^{2}\,\mathrm{d}S\qquad\text{ and }\qquad\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla g_{k,R}|^{2}\,\mathrm{d}x=\int_{\Omega^{*}\setminus\overline{B_{1/R}}}|\nabla f_{k}|^{2}\,\mathrm{d}x.$
Hence,
$\lim_{R\to\infty}\frac{\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla g_{k,R}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|g_{k,R}|^{2}\,\mathrm{d}S}=\frac{\int_{\Omega^{*}}|\nabla f_{k}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|f_{k}|^{2}\,\mathrm{d}S}=\xi_{k}\left(\Omega^{\mathrm{ext}}\right).$
So, in order to find an upper bound for $\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}_{R^{2}})$, we consider
$L(j):=\operatorname{Span}\left\{g_{k,R},1\leq k\leq j\right\}\subset H^{1}_{D}(\Omega^{\mathrm{ext}}_{R^{2}}).$
For any $u=\sum_{p=1}^{j}c_{p}g_{p,R}$, one estimates
$\lim_{R\to\infty}\sqrt{\int_{\Omega^{\mathrm{ext}}_{R^{2}}\setminus\overline{B_{R}}}|\nabla u|^{2}\,\mathrm{d}x}\leq\lim_{R\to\infty}\sum_{p=1}^{j}|c_{p}|\sqrt{\int_{\Omega^{\mathrm{ext}}_{R^{2}}\setminus\overline{B_{R}}}|\nabla g_{p,R}|^{2}\,\mathrm{d}x}=0$
and
$\lim_{R\to\infty}\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u|^{2}\,\mathrm{d}x=\int_{\Omega^{*}}\left|\nabla\left(\sum_{p=1}^{j}c_{p}f_{p}\right)\right|^{2}\mathrm{d}x,\qquad\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S=\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot\left|\sum_{p=1}^{j}c_{p}f_{p}\right|^{2}\mathrm{d}S.$
Therefore,
$\lim_{R\to\infty}\sigma_{j}^{\mathrm{D}}(\Omega^{\mathrm{ext}}_{R^{2}})\leq\lim_{R\to\infty}\sup_{0\neq u\in L(j)}\frac{\int_{\Omega^{\mathrm{ext}}_{R^{2}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}\leq\sup_{0\neq u\in\operatorname{Span}\{f_{1},\ldots,f_{j}\}}\frac{\int_{\Omega^{*}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|u|^{2}\,\mathrm{d}S}=\xi_{j}\left(\Omega^{\mathrm{ext}}\right).$
Combining both inequalities gives the desired limit. ∎
While $\xi_{j}\left(\Omega^{\mathrm{ext}}\right)\leq\sigma_{j}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ for any $R>R_{0}(\Omega)$, the Neumann-truncated eigenvalues will approximate $\xi_{j}\left(\Omega^{\mathrm{ext}}\right)$ from below.
Theorem 5.10.
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and such that $\Omega$ contains the origin. Then, for any $k\in\mathbb{N}$,
$\lim_{R\to\infty}\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)=\xi_{k}\left(\Omega^{\mathrm{ext}}\right),$
where $\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ is defined as in §3.2.
Proof.
It is straightforward to verify that $\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ is monotonically increasing in $R$ and bounded above by $\xi_{k}(\Omega^{\mathrm{ext}})$; see Lemma 3.8.
For the reverse inequality, let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ denote an eigenfunction associated to $\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ normalised to $\|\nabla u_{k,R}\|_{L^{2}\left(\Omega^{\mathrm{ext}}_{R}\right)}=1$. Then, we define
$f_{k,R}:\Omega^{*}\setminus\overline{B_{1/R}}\to\mathbb{R},\qquad f_{k,R}(x)=u_{k,R}\left(\phi(x)\right)$
Since $\phi$ is a conformal mapping,
$\int_{\Omega^{*}\setminus\overline{B_{1/R}}}|\nabla f_{k,R}|^{2}\,\mathrm{d}x=\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u_{k,R}|^{2}\,\mathrm{d}x=1.$
By [AnPo21, Lemma 4.5], there exists a constant $C_{1}>0$ which depends only on the dimension such that any $f_{k,R}$ admits a harmonic extension $\widetilde{f}_{k,R}\in H^{1}(\Omega^{*})$ satisfying
$\int_{B_{1/R}}|\nabla\widetilde{f}_{k,R}|^{2}\,\mathrm{d}x\leq C_{1}\int_{B_{2/R}\setminus B_{1/R}}|\nabla{f}_{k,R}|^{2}\,\mathrm{d}x=C_{1}\int_{B_{R}\setminus B_{R/2}}|\nabla{u}_{k,R}|^{2}\,\mathrm{d}x.$ (5.4)
We claim that the right-hand side tends to zero as $R\to\infty$. For $k=1$, this is immediate since the corresponding eigenfunction is constant. For $k\geq 2$, we have $\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)>0$. Suppose, for contradiction, that there exists $\delta>0$ and a sequence $R_{n}$ and such that $\int_{B_{R_{n}}\setminus B_{R_{n}/2}}|\nabla{u}_{k,R_{n}}|^{2}\,\mathrm{d}x\geq\delta$. Then, using the normalisation,
$\int_{\Omega^{\mathrm{ext}}_{R_{n}/2}}|\nabla{u}_{k,R_{n}}|^{2}\,\mathrm{d}x\leq 1-\delta.$
Without loss of generality, we may assume that the sequence is chosen so that $R_{n}\geq 2R_{n-1}$. Together with Lemma 3.8, this implies
$\sigma_{k}^{\mathrm{N}}(\Omega^{\mathrm{ext}}_{R_{n-1}})\leq\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R_{n}/2}\right)\leq(1-\delta)\sigma_{k}^{\mathrm{N}}(\Omega^{\mathrm{ext}}_{R_{n}})\leq(1-\delta)\xi_{k}\left(\Omega^{\mathrm{ext}}\right).$
Iterating this inequality yields $\sigma_{k}^{\mathrm{N}}(\Omega^{\mathrm{ext}}_{R_{1}})\leq(1-\delta)^{n}\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$ which vanishes as $n\to\infty$. This contradicts $\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)>0$ for $k\geq 2$. Therefore, (5.4) yields $\lim_{R\to\infty}\int_{B_{1/R}}|\nabla\widetilde{f}_{k,R}|^{2}\,\mathrm{d}x=0$.
To use the $\widetilde{f}_{k,R}$’s as trial functions, we define the $k$-dimensional space
$M(k,R):=\operatorname{Span}\left\{\widetilde{f}_{1,R},\ldots,\widetilde{f}_{k,R}\right\}\subset H^{1}(\Omega^{*}).$
Let $w=\sum_{j=1}^{k}c_{j}\widetilde{f}_{j,R}\in M(k,R)$. Using that the $u_{j,R}$’s are orthogonal with respect to $\left(\nabla f,\nabla g\right)_{L^{2}\left(\Omega^{\mathrm{ext}}_{R}\right)}$ and with respect to $\left(f,g\right)_{L^{2}(\partial\Omega)}$, we obtain, using Cauchy–Schwarz inequality,
$\int_{\Omega^{*}}|\nabla w|^{2}\,\mathrm{d}x\leq\sum_{j=1}^{k}kc_{j}^{2}\int_{B_{1/R}}|\nabla\widetilde{f}_{j,R}|^{2}\,\mathrm{d}x+\sum_{j=1}^{k}c_{j}^{2}\quad\text{and}\quad\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|w|^{2}\,\mathrm{d}S=\sum_{j=1}^{k}c_{j}^{2}\int_{\partial\Omega}|u_{j,R}|^{2}\,\mathrm{d}S.$
With $\varepsilon(R):=\max_{1\leq j\leq k}\int_{B_{1/R}}|\nabla\widetilde{f}_{j,R}|^{2}\,\mathrm{d}x$, we obtain
$\frac{\int_{\Omega^{*}}|\nabla w|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|w|^{2}\,\mathrm{d}S}\leq\frac{\sum_{j=1}^{k}c_{j}^{2}+\sum_{j=1}^{k}c_{j}^{2}k\varepsilon(R)}{\sum_{j=1}^{k}c_{j}^{2}\int_{\partial\Omega}|u_{j,R}|^{2}\,\mathrm{d}S}=\frac{\left(1+k\varepsilon(R)\right)\sum_{j=1}^{k}c_{j}^{2}}{\sum_{j=1}^{k}c_{j}^{2}\left(\sigma_{j}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)\right)^{-1}}\leq\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)\left(1+k\varepsilon(R)\right).$
Thus, $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)\leq(1+k\varepsilon(R))\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ and taking the limit $R\to\infty$ yields the desired convergence. ∎
In analogy with Corollary 5.5, the convergence of the eigenvalues also implies convergence of the associated eigenfunctions in the following sense.
Corollary 5.11.
Suppose $f_{k}\in H^{1}(\Omega^{*})$ is an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$. Then, there exists a sequence $\left(R_{m}\right)_{m\in\mathbb{N}}$ with $\lim_{m\to\infty}R_{m}=\infty$ and a sequence $\left(g_{k,R_{m}}\right)_{m\in\mathbb{N}}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)$, such that
  • $g_{k,R_{m}}(x)=0$ for $|x|>R_{m}$ and $g_{k,R_{m}}|_{\Omega^{\mathrm{ext}}_{R_{m}}}$ is an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}_{R_{m}})$,
  • $\lim_{m\to\infty}\|g_{k,R_{m}}\circ\phi-f_{k}\|_{H^{1}(\Omega^{*})}=0$.
Proof.
We distinguish two cases. For the first part, we assume $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)\neq 0$. Let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ be an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}{R})$, extended by zero to $\Omega^{\mathrm{ext}}$ and define
$f_{k,R}:\Omega^{*}\to\mathbb{R},\qquad f_{k,R}(x)=\begin{cases}u_{k,R}\left(\phi(x)\right)\qquad&\text{if }\,|x|\geq\frac{1}{R},\\ 0\qquad&\text{if }\,|x|<\frac{1}{R}.\end{cases}$
We may normalise so that
$\int_{\Omega^{*}}|\nabla f_{k}|^{2}\,\mathrm{d}x=1\qquad\text{ and }\qquad\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u_{k,R}|^{2}\,\mathrm{d}x=1\quad\text{ for all }R.$
Since $\phi$ is conformal, we also have $\int_{\Omega^{*}}|\nabla f_{k,R}|^{2}\,\mathrm{d}x=1$. From Theorem 5.9, we know
$\xi_{k}\left(\Omega^{\mathrm{ext}}\right)=\lim_{R\to\infty}\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}{R})=\lim_{R\to\infty}\frac{1}{\int_{\partial\Omega}|u_{k,R}|^{2}\,\mathrm{d}S}=\lim_{R\to\infty}\frac{1}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|f_{k,R}|^{2}\,\mathrm{d}S}.$
Moreover, as $\inf_{x\in\partial\Omega^{*}}|\phi^{\prime}(x)|>0$, the norms $\|f_{k,R}\|_{H^{1}(\Omega^{*})}$ are bounded. Consequently, there exists a subsequence, again denoted by $f_{k,R}$} converging weakly in $H^{1}(\Omega^{*})$ to some function ${g_{k}}\in H^{1}(\Omega^{*})$. As before, the convergence of the eigenvalue implies that this convergence is strong and
$\xi_{k}\left(\Omega^{\mathrm{ext}}\right)=\frac{\int_{\Omega^{*}}|\nabla g_{k}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|g_{k}|^{2}\,\mathrm{d}S},$
which means that $g_{k}$ is an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$. We can now proceed as in the proof of Corollary 5.5 to obtain the claimed statement.
If $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)=0$, any eigenfunction is constant. Let $u_{k,R}$ be an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}{R})$ normalised to $\|u_{k,R}\|_{L^{2}(\partial\Omega)}=1$. The convergence of the eigenvalue implies
$\lim_{R\to\infty}\int_{\Omega^{*}}|\nabla u_{k,R}\circ\phi|^{2}\,\mathrm{d}x=0.$
By the Poincaré inequality, this implies that $\left(u_{k,R}\circ\phi\right)-\frac{\int_{\Omega^{*}}u_{k,R}\circ\phi\,\mathrm{d}x}{|\Omega^{*}|}$ converges strongly to zero in $H^{1}(\Omega^{*})$. Moreover, since $\int_{\partial\Omega^{*}}|u_{k,R}\circ\phi|^{2}\,\mathrm{d}S\leq\frac{1}{\inf_{x\in\partial\Omega^{*}}|\phi^{\prime}(x)|}<\infty$ the average value of $u_{k,R}\circ\phi$ remains bounded. Therefore, there exists a subsequence that converges strongly in $H^{1}(\Omega^{*})$ to a constant $c\in\mathbb{R}$. ∎
As in the Dirichlet case, we next prove that the eigenfunctions associated with $\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$, when pulled back via the conformal map $\phi$, converge to the eigenfunctions of the limiting problem on $\Omega^{*}$.
Corollary 5.12.
Suppose $f_{k}\in H^{1}(\Omega^{*})$ is an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$. Then, there exists a sequence $\left(R_{m}\right)_{m\in\mathbb{N}}$ with $\lim_{m\to\infty}R_{m}=\infty$ and a sequence $\left(g_{k,R_{m}}\right)_{m\in\mathbb{N}}$ with $g_{k,R_{m}}\in H^{1}(\Omega^{*})$, such that
  • $g_{k,R_{m}}\circ\phi:\Omega^{\mathrm{ext}}\to\mathbb{R}$, restricted to $\Omega^{\mathrm{ext}}_{R_{m}}$ is an eigenfunction associated with $\sigma_{k}^{\mathrm{N}}(\Omega^{\mathrm{ext}}_{R_{m}})$,
  • $\lim_{m\to\infty}\|g_{k,R_{m}}-f_{k}\|_{H^{1}(\Omega^{*})}=0$.
Proof.
The case $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)=0$ can be handled as in Corollary 5.11. For $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)\neq 0$, we follow the proof of Theorem 5.10. Let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ denote an eigenfunction associated to $\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ normalised to $\|\nabla u_{k,R}\|_{L^{2}\left(\Omega^{\mathrm{ext}}_{R}\right)}=1$. Define
$f_{k,R}:\Omega^{*}\setminus\overline{B_{1/R}}\to\mathbb{R},\quad f_{k,R}(x)=u_{k,R}\left(\phi(x)\right)$
and let $\widetilde{f}_{k,R}$ denote its harmonic extension to $\Omega^{*}$ as described in (5.4). As in the proof of Theorem 5.10,
$\lim_{R\to\infty}\int_{\Omega^{*}}|\nabla\widetilde{f}_{k,R}|^{2}\,\mathrm{d}x=\lim_{R\to\infty}\|\nabla u_{k,R}\|^{2}_{L^{2}\left(\Omega^{\mathrm{ext}}_{R}\right)}+\int_{B_{1/R}}|\nabla\widetilde{f}_{k,R}|^{2}\,\mathrm{d}x=1.$
Moreover, the convergence of the eigenvalues yields $\lim_{R\to\infty}\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|\widetilde{f}_{k,R}|^{2}\,\mathrm{d}S=\frac{1}{\xi_{k}\left(\Omega^{\mathrm{ext}}\right)}$. As in Corollary 5.11, it follows that the sequence $\widetilde{f}_{k,R}$ admits a strongly convergent subsequence whose limit is an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$. The claim then follows by repeating the final step of the proof of Corollary 5.5. ∎
5.2.2 Conformal mapping and Helmholtz equation
In this subsection, we establish the equivalence between the formulation with conformal mappings and the formulation introduced in §3.3, based on the Helmholtz equation. Theorem  5.13 mirrors the convergence result in Theorem 5.9 for the Dirichlet problem.
Theorem 5.13.
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and such that $\Omega$ contains the origin. Then, for any $k\in\mathbb{N}$,
$\lim_{\Lambda\searrow 0}\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)=\xi_{k}\left(\Omega^{\mathrm{ext}}\right),$
where $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$ is defined as in §3.3.
Proof.
Let $u_{k,\Lambda}\in H^{1}\left(\Omega^{\mathrm{ext}}\right)$ be an eigenfunction associated with $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$. Define
$f_{k,\Lambda}:\Omega^{*}\to\mathbb{R},\qquad f_{k,\Lambda}(x)=\begin{cases}u_{k,\Lambda}\left(\phi(x)\right)\qquad&\text{ if }x\neq 0,\\ 0,\qquad&\text{ if }x=0,\end{cases}$
where $\Omega^{*}$ and $\phi$ are defined as in §3.4. Since $u_{k,\Lambda}$ vanishes at infinity for any $\Lambda>0$, it follows that $f_{k,\Lambda}\in H^{1}(\Omega^{*})$. For any given $\Lambda>0$ we define the $j$-dimensional subspaces
$M(j):=\operatorname{Span}\left\{f_{k,\Lambda}|1\leq k\leq j\right\}\subset H^{1}(\Omega^{*})\quad\text{ and }\quad N(j):=\operatorname{Span}\left\{u_{k,\Lambda}|1\leq k\leq j\right\}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right).$
Since $\frac{\int_{\Omega^{*}}|\nabla f_{k,\Lambda}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|f_{k,\Lambda}|^{2}\,\mathrm{d}S}=\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u_{k,\Lambda}|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u_{k,\Lambda}|^{2}\,\mathrm{d}S}$ and $\Lambda>0$, we have
$\xi_{j}\left(\Omega^{\mathrm{ext}}\right)\leq\sup_{0\neq u\in M(j)}\frac{\int_{\Omega^{*}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|u|^{2}\,\mathrm{d}S}\leq\sup_{0\neq u\in N(j)}\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u|^{2}\,\mathrm{d}x+\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}=\mu_{j}\left(\Omega^{\mathrm{ext}},\Lambda\right).$
For the reverse inequality, let $f_{k}\in H^{1}(\Omega^{*})$ be an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$, normalised so that $\int_{\partial\Omega^{*}}|\phi|f_{j}f_{k}\,\mathrm{d}S=\delta_{j,k}$. We construct a cut-off function by
$g_{k,R}:\Omega^{\mathrm{ext}}\to\mathbb{R},\qquad g_{k,R}(x)=\begin{cases}f_{k}\left(\phi^{-1}(x)\right)\quad&\text{ if }x\in\Omega^{\mathrm{ext}}_{R},\\ f_{k}\left(\phi^{-1}(x)\right)\left(\frac{\log(R^{2})-\log(|x|)}{\log(R^{2})-\log(R)}\right)\quad&\text{ if }R<|x|\leq R^{2},\\ 0\quad&\text{ if }|x|>R^{2}.\end{cases}$
Now define the subspaces
$L(j,R):=\operatorname{Span}\left\{g_{k,R},1\leq k\leq j\right\}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)\quad\text{ and }\quad K(j):=\operatorname{Span}\left\{f_{k}|1\leq k\leq j\right\}\subset H^{1}(\Omega^{*}).$
Since $L(j,R)$ is independent of $\Lambda$, we may pass to the limit and obtain
$\mu_{j}\left(\Omega^{\mathrm{ext}},\Lambda\right)\leq\sup_{0\neq u\in L(j,R)}\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u|^{2}\,\mathrm{d}x+\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}\xrightarrow{\Lambda\searrow 0}\sup_{0\neq u\in L(j,R)}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$ (5.5)
Let $u=\sum_{k=1}^{j}c_{k}g_{k,R}$ and $v=\sum_{k=1}^{j}c_{k}f_{k}$. For any $R>R_{0}(\Omega)$, we have $\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S=\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|v|^{2}\,\mathrm{d}S$ by construction. Moreover, using that
$\int_{B_{R^{2}}\setminus B_{R}}\left|\nabla\left(\frac{\log(R^{2})-\log(|x|)}{\log(R^{2})-\log(R)}\right)\right|^{2}\,\mathrm{d}x=\int_{B_{R^{2}}\setminus B_{R}}\frac{1}{|x|^{2}\log(R)^{2}}\,\mathrm{d}x=\int_{R}^{R^{2}}\frac{1}{r\log(R)^{2}}\,\mathrm{d}x=\frac{1}{\log(R)},$
a straightforward computation yields
$\lim_{R\to\infty}\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla g_{k,R},\nabla g_{i,R}\right\rangle\,\mathrm{d}x=\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla(f_{k}\circ\phi^{-1}),\nabla(f_{i}\circ\phi^{-1})\right\rangle\,\mathrm{d}x=\int_{\Omega^{*}}\left\langle\nabla f_{k},\nabla f_{i}\right\rangle\,\mathrm{d}x.$
Hence,
$\lim_{R\to\infty}\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x=\int_{\Omega^{*}}|\nabla v|^{2}\,\mathrm{d}x.$
Taking the limit in (5.5), we conclude
$\lim_{\Lambda\searrow 0}\mu_{j}\left(\Omega^{\mathrm{ext}},\Lambda\right)\leq\sup_{0\neq u\in K(j)}\frac{\int_{\Omega^{*}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|u|^{2}\,\mathrm{d}S}=\xi_{j}\left(\Omega^{\mathrm{ext}}\right).$
Corollary 5.14.
Suppose $f_{k}\in H^{1}(\Omega^{*})$ is an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$. Then, there exists a sequence $\left(\Lambda_{m}\right)_{m\in\mathbb{N}}\subset(0,1)$ with $\lim_{m\to\infty}\Lambda_{m}=0$ and a sequence $\left(g_{k,\Lambda_{m}}\right)_{m\in\mathbb{N}}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)$, such that
  • $g_{k,\Lambda_{m}}$ is an eigenfunction associated with $\mu_{k}(\Omega^{\mathrm{ext}},\Lambda_{m})$ and $\lim_{m\to\infty}\mu_{k}(\Omega^{\mathrm{ext}},\Lambda_{m})=\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$,
  • $\lim_{m\to\infty}\|g_{k,\Lambda_{m}}\circ\phi-f_{k}\|_{H^{1}(\Omega^{*})}=0$.
Proof.
Without restriction we may assume $\|\nabla f_{k}\|_{L^{2}(\Omega^{*})}=1$. Let $u_{k,\Lambda}\in H^{1}\left(\Omega^{\mathrm{ext}}\right)$ denote the eigenfunctions associated with $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$ such that $\|\nabla u_{k,\Lambda}\|_{L^{2}\left(\Omega^{\mathrm{ext}}\right)}=1$ and define
$f_{k,\Lambda}:\Omega^{*}\to\mathbb{R},\qquad f_{k,\Lambda}(x)=\begin{cases}u_{k,\Lambda}\left(\phi(x)\right)\qquad&\text{ if }x\neq 0,\\ 0\qquad&\text{ if }x=0.\end{cases}$
Then,
$\frac{1}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|f_{k,\Lambda}|^{2}\,\mathrm{d}S}=\frac{1}{\int_{\partial\Omega}|u_{k,\Lambda}|^{2}\,\mathrm{d}S}\leq\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u_{k,\Lambda}|^{2}\,\mathrm{d}x+1}{\int_{\partial\Omega}|u_{k,\Lambda}|^{2}\,\mathrm{d}S}=\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right).$
To obtain a reverse inequality, define the $k$-dimensional spaces
$N(\Lambda):=\operatorname{Span}\left\{u_{1,\Lambda},\ldots,u_{k,\Lambda}\right\}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)\quad\,\text{ and }\,\quad M(\Lambda):=\operatorname{Span}\left\{f_{1,\Lambda},\ldots,f_{k,\Lambda}\right\}\subset H^{1}(\Omega^{*})$
and observe that
$\xi_{k}\left(\Omega^{\mathrm{ext}}\right)\leq\sup_{0\neq u\in M(\Lambda)}\frac{\int_{\Omega^{*}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|u|^{2}\,\mathrm{d}S}=\sup_{0\neq u\in N(\Lambda)}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$
In particular, for $k=1$, we get $\lim_{\Lambda\searrow 0}\int_{\partial\Omega^{*}}|\phi^{\prime}||f_{1,\Lambda}|^{2}\,\mathrm{d}S=\lim_{\Lambda\searrow 0}\int_{\partial\Omega}|u_{1,\Lambda}|^{2}\,\mathrm{d}S=\frac{1}{\xi_{1}\left(\Omega^{\mathrm{ext}}\right)}$. As before, this implies that a subsequence $f_{1,\Lambda_{m}}$ converges in $H^{1}(\Omega^{*})$ to a limit $g_{1}$ that is an eigenfunction of $\xi_{1}\left(\Omega^{\mathrm{ext}}\right)$. Since $\xi_{1}\left(\Omega^{\mathrm{ext}}\right)$ is simple, the normalisation implies $f_{1}=g_{1}$.
The general case follows by induction, using the same argument as in the proof of Corollary 5.7, with straightforward modifications, and thus completes the proof. ∎
5.2.3 Conformal mapping and layer potentials
In §4.2, we introduced a formulation of the exterior Steklov eigenvalue problem in two dimensions using boundary layer potentials. Using Theorem 4.3, we can readily verify the equivalence with the conformal mapping formulation.
Theorem 5.15.
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and such that $\Omega$ contains the origin. Let $\tau_{k}(\Omega^{\mathrm{ext}})$ be defined as in Theorem 4.3. Then, $\tau_{k}(\Omega^{\mathrm{ext}})=\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$ for any $k\in\mathbb{N}$.
Let $u_{k}\in H^{\frac{1}{2}}(\partial\Omega)$ be an eigenfunction associated with $\tau_{k}(\Omega^{\mathrm{ext}})$, and let $U_{k}$ be its harmonic extension. Then, $U_{k}\circ\phi$ coincides, up to a constant factor, with an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$.
Proof.
As $\tau_{1}=\xi_{1}=0$ with constant eigenfunctions, we only have to consider $k\geq 2$.
Let $w_{k}\in H^{1}(\Omega^{*})$ be an eigenfunction corresponding to $\xi_{k}\neq 0$. Define
$f_{k}:\Omega^{\mathrm{ext}}\to\mathbb{R},\quad f_{k}(x):=(w_{k}\circ\phi^{-1})(x).$
Since $\lim_{|x|\to\infty}f_{k}(x)=\lim_{x\to 0}w_{k}(x)=w_{k}(0)$, the mean value theorem yields
$\left|f_{k}(x)-w_{k}(0)\right|=\left|w_{k}\left(\frac{x}{|x|^{2}}\right)-w_{k}(0)\right|=\frac{|\nabla w_{k}(x_{0})|}{|x|^{2}}$
for some $x_{0}$ between 0 and $\frac{x}{|x|^{2}}$. By regularity of $w_{k}$ we know that $|\nabla w_{k}(x_{0})|$ is bounded, so
$\left|f_{k}(x)-w_{k}(0)\right|=O\left(\frac{1}{|x|}\right)\quad\text{ as }|x|\to\infty.$
Therefore, $f_{k}$ satisfies the far field condition (4.3) and by Theorem 4.3, the pair $(f_{k},\xi_{k})$ is also a solution of (4.7).
Conversely, let $u_{k}\in\left\{u\in H^{\frac{1}{2}}(\partial\Omega):\left(u,1\right)_{\partial\Omega}=0\right\}$ be an eigenfunction corresponding to $\tau_{k}\neq 0$. Define $(u_{k})_{\infty}\in\mathbb{R}$ as in (4.6) and let $U_{k}:\Omega^{\mathrm{ext}}\to\mathbb{R}$ be the harmonic extension of $u_{k}$, as given in Theorem 4.3. Then, $W_{k}\in H^{1}(\Omega^{*})$, given by
$W_{k}(x)=\begin{cases}U_{k}(\phi(x))\qquad&\text{ for }x\neq 0,\\ (u_{k})_{\infty}\qquad&\text{ for }x=0,\end{cases}$
is a solution of ($\mathrm{CT}$). That completes the proof. ∎
§6.  Properties of exterior Steklov eigenvalues and eigenfunctions
In the following, we no longer distinguish between the different formulations of the Steklov problem, as they are equivalent and denote the Steklov eigenvalue by $\sigma$.
§6.1.  Basic properties
Proof of Theorem 1.3.
Let $n\geq 3$, let $u$ be a solution of ($\mathrm{ES}$) and let $f$ be the restriction of $u$ to $\partial\Omega$. It follows from [AuHa14b, Section 5] that there is a unique $U\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ such that $U=f$ on $\partial\Omega$. Moreover, $U=O(|x|^{2-n})$ by [Xio23, Proposition 12] (see also Proposition 6.1). It follows that UuU-u is harmonic, vanishing on $\Gamma$ and decaying at $\infty$, whence $u=U\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. So, any solution of ($\mathrm{ES}$) is in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$. It then follows by integration by parts against test function that $u$ is a solution of the finite energy weak Steklov eigenvalue equation (3.5). On the other hand, Theorem 3.5 provides a sequence of weak Steklov eigenfunction $u_{k}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ whose restriction to $\partial\Omega$ form an orthogonal basis of $L^{2}(\partial\Omega)$. It follows from [Xio23, Proposition 12] that these functions satisfy (1.2), and hence they are solutions of ($\mathrm{ES}$).
Let $n=2$ and let $u$ be a solution of ($\mathrm{ES}$). Because $u$ is harmonic and bounded, it converges to a finite value at infinity, denoted by $u_{\infty}$. With $\phi$ as is (3.10), we consider the function $v:\Omega^{*}\to\mathbb{R}$, $v=u\circ\phi$ in $\Omega^{*}\setminus\{0\}$ and $v(0)=u_{\infty}$. Then, $v$ is harmonic and satisfies ($\mathrm{CT}$) and so it is a Steklov eigenfunction in the conformal sense. Similarly, any solution of ($\mathrm{CT}$) will give a solution of ($\mathrm{ES}$) once it is inverted through $\phi$. ∎
6.1.1 Courant-type bound for the nodal count
Proof of Theorem 1.8.
For $n\geq 3$, we can use the finite-energy approach and proceed verbatim as in the interior case, see, e.g., [KuSi69].
For $n=2$, if we show that the number of nodal domains in $\Omega^{\mathrm{ext}}$ is preserved under the conformal transformation $\phi$, defined in (3.10), then the claim follows from Courant’s nodal domain theorem for the bounded domain $\Omega^{*}$. Let $\Omega\subset\mathbb{R}^{2}$ be a bounded Lipschitz domain, and let $u$ be a solution of ($\mathrm{ES}$). Since $\phi$ is a homeomorphism between $\Omega^{\mathrm{ext}}$ and $\Omega^{*}\setminus\{0\}$, it preserves connected components, and hence the number of nodal domains of $u\in\Omega^{\mathrm{ext}}$ equals the number of nodal domains of $w:=u\circ\phi$ in $\Omega^{*}\setminus\{0\}$. Since $u$ is bounded, it has a limit at infinity, and therefore $w$ extends continuously to zero. If $w(0)=0$, then the origin is not part of any nodal domain, so extending $w$ to the origin does not merge or split any connected components. If $w(0)\neq 0$, there exists an $R>0$ such that $u(x)\neq 0$ for any $|x|>R$. Then, $w(x)\neq 0$ for any $|x|<\frac{1}{R}$. So again, extending $w$ to the origin does not merge or split any connected components. ∎
Proof of Corollary 1.9.
Suppose, for contradiction, that there are two linearly independent eigenfunctions $u_{1},u_{2}$ associated with $\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)$. By Theorem 1.8, we may choose them so that $u_{1}(x),u_{2}(x)\geq 0$ for any $x\in\Omega^{\mathrm{ext}}$. Take some $x_{0}\in\Omega^{\mathrm{ext}}$, and choose $c\leq 0$ such that $u_{1}(x_{0})+cu_{2}(x_{0})=0$. Then, $w(x):=u_{1}(x)+cu_{2}(x)$ is a nontrivial eigenfunction associated with $\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)$, and $w(x_{0})=0$. By Theorem 1.8 and the maximum principle, $w$ cannot vanish at an interior point unless it is identically zero. This contradiction implies that the first eigenvalue is simple. ∎
6.1.2 Decay rate of the eigenfunctions
First, we recall a result from [Xio23, Proposition 12], which describes a decay rate of the eigenfunctions.
Proposition 6.1.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded domain with Lipschitz boundary and connected $\Omega^{\mathrm{ext}}$. Any harmonic function $u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ satisfies
$u=O\left(|x|^{2-n}\right)\quad\text{as}\quad|x|\to\infty.$
In particular, any eigenfunction associated with $\sigma_{k}\left(\Omega^{\mathrm{ext}}\right)$, $k\in\mathbb{N}$, behaves asymptotically as $O\left(|x|^{2-n}\right)$.
To get a more precise understanding of the decay rate, we show below that the first eigenfunction can not decay faster than $|x|^{2-n}$.
Proposition 6.2.
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded domain with Lipschitz boundary and connected $\Omega^{\mathrm{ext}}$ and let $u_{1}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ be a eigenfunction associated with $\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)$. Then,
$\lim_{|x|\to\infty}\frac{u_{1}(x)}{|x|^{2-n}}\neq 0.$
For $n\in\{3,4\}$ this shows, in particular, that the first eigenfunction is not in $L^{2}\left(\Omega^{\mathrm{ext}}\right)$.
Proof.
Outside a ball, $u_{1}$ can be expanded in terms of Steklov eigenfunctions of the ball, whose explicit form is given in §2.2. The asymptotics then follow from the leading term of this expansion.
Without loss of generality, assume that $\Omega\subset B_{1}$. With the notation from §2.2, we define for $r\geq 1$, $\theta\in\mathbb{S}^{n-1}$, and $k\in\mathbb{N}_{0}$, $1\leq i\leq d_{n,k}$,
$u_{(k,i)}(r,\theta)=r^{2-n-k}Y_{k,i}(\theta).$
Then ${u}_{(k,i)}\in E^{1}\left(B_{1}^{\mathrm{ext}}\right)$ are Steklov eigenfunctions associated to $\sigma_{(k)}\left(B_{1}^{\mathrm{ext}}\right)=n+k-2$, and their restrictions to $\partial B_{1}$ form an orthonormal basis in $L^{2}\left(\partial B_{1}\right)$, see Theorem 1.3. Let $f:=u_{1}|_{\partial B_{1}}$. In $L^{2}\left(\partial B_{1}\right)$ we have the expansion
$f(s)=\sum_{k=0}^{\infty}\sum_{i=1}^{d_{n,k}}\left(f,{u}_{(k,i)}\right)_{L^{2}\left(\partial B_{1}\right)}{u}_{(k,i)}(s),\qquad s\in\partial B_{1}.$ (6.1)
In the first step, we show uniform convergence of (6.1). Since $u_{1}$ is harmonic, $u_{1}\in C^{\infty}\left(\Omega^{\mathrm{ext}}\right)$ and hence $f\in C^{\infty}\left(\partial B_{1}\right)$. Using the fact that the spherical harmonics $Y_{k,i}$ are eigenfunctions of the Laplace–Beltrami operator $\Delta_{\mathbb{S}}$ with eigenvalue k(k+n2)k(k+n-2), one obtains
$\left|\left(f,{u}_{(k,i)}\right)_{L^{2}\left(\partial B_{1}\right)}\right|=\frac{1}{k(k+n-2)}\left|\left(f,-\Delta_{\mathbb{S}}Y_{k,i}\right)_{L^{2}\left(\partial B_{1}\right)}\right|=\frac{1}{k(k+n-2)}\left|\left(-\Delta_{\mathbb{S}}f,Y_{k,i}\right)_{L^{2}\left(\partial B_{1}\right)}\right|.$
Repeating this process $m$ times and using the Cauchy–Schwarz inequality,
$\left|\left(f,{u}_{(k,i)}\right)_{L^{2}\left(\partial B_{1}\right)}\right|=\frac{1}{k^{m}(k+n-2)^{m}}\left|\left((-\Delta_{\mathbb{S}})^{m}f,Y_{k,i}\right)_{L^{2}\left(\partial B_{1}\right)}\right|\leq\frac{\left\|(-\Delta_{\mathbb{S}})^{m}f\right\|_{L^{2}\left(\partial B_{1}\right)}}{k^{m}(k+n-2)^{m}}.$
Because $f$ is smooth, $\left\|(-\Delta_{\mathbb{S}})^{m}f\right\|_{L^{2}\left(\partial B_{1}\right)}$ is bounded. Choosing $m$ large enough and noting that $d_{n,k}=O(k^{n-2})$, we see that (6.1) converges uniformly.
By uniqueness of the harmonic extension, the same coefficients give the expansion of $u_{1}$ in $\Omega^{\mathrm{ext}}$,
$u_{1}(x)=\sum_{k=0}^{\infty}\sum_{i=1}^{d_{n,k}}\left(f,{u}_{(k,i)}\right)_{L^{2}\left(\partial B_{1}\right)}{u}_{(k,i)}(x),\qquad x\in\Omega^{\mathrm{ext}}.$
Since ${u}_{(k,i)}(x)=O(|x|^{2-n-k})$, it remains to show that $\left(f,{u}_{0,1}\right)_{L^{2}\left(\partial B_{1}\right)}\neq 0$. Since spherical harmonics of degree 0 are constants, this is equivalent to $\int_{\partial B_{1}}f\,\mathrm{d}S\neq 0$.
If $\int_{\partial B_{1}}f\,\mathrm{d}S=0$, then $f=0$ on $\partial B_{1}$ since $u_{1}$ does not change sign by Theorem 1.8. Then, $u_{1}\equiv 0$ in $\Omega^{\mathrm{ext}}$. Therefore, the coefficient of degree $k=0$ in the spherical expansion of $u_{1}$ has to be nonzero. ∎
§6.2.  Lower bounds for the first eigenvalue
The only established lower bound so far, to the best of our knowledge, is given in [Xio23, Theorem 1]. In this section we prove a new lower bound depending on the mean curvatures of the boundary. In Figure 7, we compare the bound of [Xio23] with our result for different spheroids, showing that neither inequality implies the other.
Proof of Theorem 1.11.
If $\prod_{j=1}^{n-1}\kappa_{j}(s)=0$, the inequality becomes trivial. Thus, we assume $\kappa_{j}(s)>0$ for any $j\leq n-1$.
We adapt the method from [KoPa17, Proof of Theorem 1.1], see also [KrLo18, KrLo20], to derive a lower bound. Since $\Omega$ is convex, the mapping
$\Psi:\partial\Omega\times(0,\infty)\to\Omega^{\mathrm{ext}},\qquad(s,t)\mapsto s-t\nu(s)$
is bijective and locally bi-Lipschitz. For a discussion of the properties and further background on $\Psi$, we refer to [KoPa17, Chapter 3] and the references therein (note that we will switch the sign in front of $\kappa_{j}(s)$ in (6.2) compared to [KoPa17] because of the orientation of $\nu$). Any $u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ can be approximated by functions with compact support in $\Omega^{\mathrm{ext}}$, allowing us to perform a change of variables, yielding
$\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x=\int_{\partial\Omega\times(0,\infty)}|(\nabla u)\circ\Psi(s,t)|^{2}\operatorname{det}(D\Psi)\,\mathrm{d}S_{s}\,\mathrm{d}t,$
where the Jacobian determinant is given by
$\zeta(s,t):=\operatorname{det}(D\Psi)=\prod_{j=1}^{n-1}\left(1+\kappa_{j}(s)t\right).$ (6.2)
Define $w(s,t):=u(\Psi(s,t))$. By the chain rule, $\partial_{t}w(s,t)=-\left\langle(\nabla u)(\Psi(s,t)),\nu(s)\right\rangle$ and hence
$|(\nabla u)\circ\Psi(s,t)|\geq|\left\langle(\nabla u)\circ\Psi(s,t),\nu(s)\right\rangle|=|\partial_{t}w(s,t)|.$
It follows that
$\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x=\int_{\partial\Omega\times(0,\infty)}|(\nabla u)\circ\Psi(s,t)|^{2}\zeta(s,t)\,\mathrm{d}S_{s}\,\mathrm{d}t\geq\int_{\partial\Omega\times(0,\infty)}|\partial_{t}w(s,t)|^{2}\zeta(s,t)\,\mathrm{d}S_{s}\,\mathrm{d}t.$
Thus, we obtain the lower bound
$\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)=\inf_{0\neq u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}\geq\inf_{\begin{subarray}{c}0\neq u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)\\ w=u\circ\Psi\end{subarray}}\frac{\int_{\partial\Omega\times(0,\infty)}|\partial_{t}w(s,t)|^{2}\zeta(s,t)\,\mathrm{d}S_{s}\,\mathrm{d}t}{\int_{\partial\Omega}|w(s,0)|^{2}\,\mathrm{d}S_{s}}.$
Since $w=u\circ\Psi$ decays at infinity for any $u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$, we define, for each $s\in\partial\Omega$,
$K\left(s,\Omega^{\mathrm{ext}}\right):=\inf_{\begin{subarray}{c}f\in{H}_{\mathrm{loc}}^{1}\left((0,\infty)\right)\\ \lim_{x\to\infty}f(x)=0\end{subarray}}\frac{\int_{0}^{\infty}|f^{\prime}(t)|^{2}\zeta(s,t)\,\mathrm{d}t}{|f(0)|^{2}},$ (6.3)
which gives
$\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)\geq\inf_{\begin{subarray}{c}0\neq u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)\\ w=u\circ\Psi\end{subarray}}\frac{\int_{\partial\Omega}K(s,\Omega^{\mathrm{ext}})|w(s,0)|^{2}\,\mathrm{d}S_{s}}{\int_{\partial\Omega}|w(s,0)|^{2}\,\mathrm{d}S_{s}}\geq\inf_{s\in\partial\Omega}K\left(s,\Omega^{\mathrm{ext}}\right).$ (6.4)
To show that the infimum in (6.3) is attained, consider a minimising sequence $(f_{k})_{k\in\mathbb{N}}\subset H^{1}_{\mathrm{loc}}((0,\infty))$ with $f_{k}(0)=1$, $f_{k}(x)\to 0$ as $x\to\infty$ and $\lim_{k\to\infty}\int_{0}^{\infty}|f_{k}^{\prime}(t)|^{2}\zeta(s,t)\,\mathrm{d}t=K\left(s,\Omega^{\mathrm{ext}}\right)$. It is straightforward to show that, for each $R>0$, the sequence is bounded in $H^{1}((0,R))$. Hence, by standard compactness results, there exists a subsequence, again denoted by $f_{k}$, that converges weakly in $H^{1}((0,R))$ and strongly in $L^{2}((0,R))$ to a limit function $f\in H^{1}((0,R))$.
It remains to show that $f$ decays at infinity. Since $n\geq 3$ and $\prod_{j=1}^{n-1}\kappa_{j}(s)>0$, there exists $D_{1}(s)>0$ with $\zeta(s,t)\geq D_{1}(s)t^{n-1}$. Thus,
$\int_{x}^{\infty}\frac{1}{\zeta(s,t)}\,\mathrm{d}t<\frac{x^{2-n}}{(n-2)D_{1}(s)}.$
Then, there exists a constant $D_{2}(s)$ such that for any $x\in(0,\infty)$,
$|f_{k}(x)|=\left|\int_{x}^{\infty}f_{k}^{\prime}(t)\,\mathrm{d}t\right|\leq\int_{x}^{\infty}|f_{k}^{\prime}(t)|\,\mathrm{d}t\leq\sqrt{\int_{x}^{\infty}|f_{k}^{\prime}(t)|^{2}\zeta(s,t)\,\mathrm{d}t}\sqrt{\int_{x}^{\infty}\frac{1}{\zeta(s,t)}\,\mathrm{d}t}\leq D_{2}(s)x^{\frac{2-n}{2}},$
where we use that the first integral in the product is uniformly bounded for any minimising sequence. This uniform decay implies $f(x)\to 0$ as $x\to\infty$. Finally, by the lower semicontinuity of the functional, $f$ minimises (6.3). Moreover, $K\left(s,\Omega^{\mathrm{ext}}\right)>0$ since $f$ is nonconstant. The associated Euler–Lagrange equation for this minimisation problem reads
$\left(f^{\prime}(t)\zeta(s,t)\right)^{\prime}=0\,\text{ for }t\in(0,\infty),\qquad-f^{\prime}(0)=K\left(s,\Omega^{\mathrm{ext}}\right)f(0),\qquad f(t)\to 0\,\text{ as }t\to\infty.$ (6.5)
Assume now that the principal curvatures $\kappa_{j}(s)$ are distinct. Then the general solution of the differential equation in (6.5) is given by
$\begin{split}f(t)&=C_{1}+C_{2}\int_{0}^{t}\frac{1}{\zeta(s,r)}\mathrm{d}r=C_{1}+C_{2}\int_{0}^{t}\left(\sum_{j=1}^{n-1}\frac{c_{j}(s)}{1+\kappa_{j}(s)r}\right)\,\mathrm{d}r\\ &=C_{1}+C_{2}\sum_{j=1}^{n-1}\frac{c_{j}(s)\log(1+\kappa_{j}(s)t)}{\kappa_{j}(s)}=C_{1}+C_{2}\log\left(\prod_{j=1}^{n-1}(1+\kappa_{j}(s)t)^{\frac{c_{j}(s)}{\kappa_{j}(s)}}\right),\end{split}$
where $c_{j}(s)=\frac{\kappa_{j}(s)^{n-2}}{\prod_{\begin{subarray}{c}i=1,\,i\neq j\end{subarray}}^{n-1}(\kappa_{j}(s)-\kappa_{i}(s))}$ are the coefficients of the partial fraction decomposition
$1=\sum_{j=1}^{n-1}c_{j}\prod_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n-1}(1+\kappa_{i}(s)r).$
To ensure that $f(t)$ is bounded, we require $\sum_{j=1}^{n-1}\frac{c_{j}}{\kappa_{j}(s)}=0$. Thus, the asymptotic behaviour of $f(t)$ simplifies to
$\lim_{t\to\infty}f(t)=C_{1}+C_{2}\log\left(\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{c_{j}}{\kappa_{j}(s)}}\right),$
and the solution decays at infinity if and only if $C_{1}=-C_{2}\log\left(\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{c_{j}}{\kappa_{j}(s)}}\right)$. Because $\sum_{j=1}^{n-1}c_{j}=1$, this yields
$K\left(s,\Omega^{\mathrm{ext}}\right)=\frac{-f^{\prime}(0)}{f(0)}=\frac{1}{\log\left(\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{c_{j}}{\kappa_{j}(s)}}\right)}=\frac{1}{\sum_{j=1}^{n-1}\frac{c_{j}}{\kappa_{j}(s)}\log\left(\kappa_{j}(s)\right)}.$
Substituting the expression for $c_{j}$, we arrive at
$K\left(s,\Omega^{\mathrm{ext}}\right)=\frac{1}{\sum_{j=1}^{n-1}\frac{\kappa_{j}(s)^{n-3}}{\prod_{\begin{subarray}{c}i=1,\,i\neq j\end{subarray}}^{n-1}(\kappa_{j}(s)-\kappa_{i}(s))}\log\left(\kappa_{j}(s)\right)},$
which completes the proof when the curvatures $\kappa_{j}(s)$ are distinct. When they are not, we simply proceed by approximation. ∎
Proof of Corollary 1.12.
For distinct, positive $\alpha_{1},\ldots,\alpha_{n-1}$, by [Pi85, Proposition 4.1, Theorem 3.8],
$L(\alpha_{1},\ldots,\alpha_{2})\geq(n-2)\sqrt[n-1]{\prod_{j=1}^{n-1}|\alpha_{j}(s)|}.$
So (1.3) already implies (1.4) in the case of distinct principal curvatures. Situations where multiple $\kappa_{i}$’s share the same value can be treated as limiting cases of the above computation.
It remains to prove that equality holds only for balls. Suppose equality holds in (1.4). Then both inequalities in (6.4) must be equalities, and
$\inf_{s\in\partial\Omega}K\left(s,\Omega^{\mathrm{ext}}\right)=(n-2)\inf_{s\in\partial\Omega}\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{1}{n-1}}.$ (6.6)
The first inequality in (6.4) can be an equality only if $u$ is an eigenfunction; since the first eigenfunction does not change sign (see §6.1.1), the second inequality can only be equal if $K\left(s,\Omega^{\mathrm{ext}}\right)$ is constant. By (1.4) and (6.6), the geometric mean of the curvatures must also be constant. By [Mu87, Theorem 2], this forces $\partial\Omega$ to be a sphere. ∎
We now compare our lower bound (1.3) with the one from [Xio23, Theorem 1], considering prolate and oblate spheroids.
Example 6.3.
For $a\in(0,1)$, we consider two families of spheroids,
$\mathscr{p}_{a}:=\left\{x\in\mathbb{R}^{3}:\frac{x_{1}^{2}}{a^{2}}+\frac{x_{2}^{2}}{a^{2}}+x_{3}^{2}=1\right\}\qquad\text{ and }\qquad\mathscr{o}_{a}:=\left\{x\in\mathbb{R}^{3}:\frac{x_{1}^{2}}{a^{2}}+x_{2}^{2}+x_{3}^{2}=1\right\},$
and denote the interior and exterior, with respect to $\mathscr{p}_{a}$ and $\mathscr{o}_{a}$, domains in $\mathbb{R}^{3}$ by $\mathcal{P}_{a}$ and $\mathcal{P}^{\mathrm{ext}}_{a}$, and $\mathcal{O}_{a}$ and $\mathcal{O}^{\mathrm{ext}}_{a}$, respectively.
The principal curvatures of $\mathscr{p}_{a}$ and $\mathscr{o}_{a}$ are given by
$\kappa_{1}^{\mathscr{p}_{a}}(x)$ $=\frac{a}{\left(1-(1-a^{2})x_{3}^{2}\right)^{\frac{3}{2}}},\quad$ $\kappa_{2}^{\mathscr{p}_{a}}(x)=\frac{1}{a\left(1-(1-a^{2})x_{3}^{2}\right)^{\frac{1}{2}}},$
$\kappa_{1}^{\mathscr{o}_{a}}(x)$ $=\frac{a^{4}}{\left(a^{4}+(1-a^{2})x_{1}^{2}\right)^{\frac{3}{2}}},\quad$ $\kappa_{2}^{\mathscr{o}_{a}}(x)=\frac{a^{2}}{\left(a^{4}+(1-a^{2})x_{1}^{2}\right)^{\frac{1}{2}}}.$
Since $\kappa_{1}^{\mathscr{p}_{a}}(x)$ and $\kappa_{2}^{\mathscr{p}_{a}}(x)$ become minimal when $|x_{3}|=0$, we have for prolate spheroids
$\beta\left(\mathscr{p}_{a}\right)=\min_{x\in\mathscr{p}_{a}}L\left(\kappa_{1}^{\mathscr{p}_{a}}(x),\kappa_{2}^{\mathscr{p}_{a}}(x)\right)=L\left(a,a^{-1}\right)=\frac{1-a^{2}}{-2a\log(a)}$ (6.7)
and analogue for oblate spheroids,
$\beta\left(\mathscr{o}_{a}\right)=\min_{x\in\mathscr{o}_{a}}L\left(\kappa_{1}^{\mathscr{o}_{a}}(x),\kappa_{2}^{\mathscr{o}_{a}}(x)\right)=L\left(a,a\right)=a.$ (6.8)
We now compare our eigenvalue bound (1.3) with Xiong’s bound (1.5). A straightforward calculation yields
$\beta_{\mathrm{X}}\left(\mathscr{p}_{a}\right)=\beta_{\mathrm{X}}\left(\mathscr{o}_{a}\right)=\begin{cases}\frac{3\sqrt{3}}{2}\frac{a}{(1+a^{2})^{\frac{3}{2}}}\qquad&\text{if }0<a\leq\frac{1}{\sqrt{2}},\\ 1\qquad&\text{if }\frac{1}{\sqrt{2}}<a<1.\end{cases}$ (6.9)
The comparison of the bounds (6.7)–(6.9) and the first Steklov eigenvalues for prolate and oblate spheroids, computed numerically using the separation of variables technique [Gr24], is presented in Figure 7.
Figure 7: Comparison between the numerical results, our bounds (6.7), (6.8), and Xiong’s bound (6.9) for the exterior of prolate spheroids (top figure) and oblate spheroids (bottom figure).
Note that for oblate spheroids, our lower bound is weaker than the one from [Xio23], however, neither bound appears to be effective as $a\to 0$. In contrast, for prolate spheroids, the lower bound from [Xio23] vanishes as $a\to 0$, whereas our lower bound (6.7) exhibits an asymptotic behaviour that, up to a constant factor, agrees with the numerical results observed in [Gr24, Section II.F], see also Figure 8 below. In addition, Remark 6.10 will provide, later on, an upper bound that differs from the numerically observed asymptotic behaviour only by a logarithmic factor.
In order to better compare the behaviour of bounds for prolate spheroids for different values of $a$ and to illustrate Theorem 1.18 in dimension three, we normalise the solid prolate spheroids by volume, by setting
$\widetilde{\mathscr{p}}_{a}:=a^{-\frac{2}{3}}\mathscr{p}_{a},\qquad\widetilde{\mathcal{P}}_{a}:=a^{-\frac{2}{3}}\mathcal{P}_{a},\qquad\text{and}\qquad\widetilde{\mathcal{P}}^{\mathrm{ext}}_{a}:=a^{-\frac{2}{3}}\mathcal{P}^{\mathrm{ext}}_{a},$
so that $\left|\widetilde{\mathcal{P}}_{a}\right|=\frac{4\pi}{3}$ is constant for all $a\in(0,1)$. Then, bound (6.7) and the scaling $\sigma_{1}(\alpha\Omega^{\mathrm{ext}})=\alpha^{-1}\sigma_{1}(\Omega^{\mathrm{ext}})$ yield
$\sigma_{1}\left(\widetilde{\mathcal{P}}_{a}^{\mathrm{ext}}\right)\geq\beta\left(\widetilde{\mathscr{p}}_{a}\right)=a^{\frac{2}{3}}\beta\left(\mathscr{p}_{a}\right)=\frac{1-a^{2}}{-2a^{\frac{1}{3}}\log(a)}\xrightarrow[]{a\searrow 0}+\infty,$ (6.10)
and bound (6.9) becomes
$\sigma_{1}\left(\widetilde{\mathcal{P}}_{a}^{\mathrm{ext}}\right)\geq\beta_{\mathrm{X}}\left(\widetilde{\mathscr{p}}_{a}\right)=a^{\frac{2}{3}}\beta_{\mathrm{X}}\left(\mathscr{p}_{a}\right)=\begin{cases}\frac{3\sqrt{3}}{2}\frac{a^{\frac{4}{3}}}{(1+a^{2})^{\frac{3}{2}}}\qquad&\text{if }0<a\leq\frac{1}{\sqrt{2}},\\ a^{\frac{2}{3}}\qquad&\text{if }\frac{1}{\sqrt{2}}<a<1.\end{cases}$ (6.11)
In Figure 8, we compare both lower bounds and numerical computations of $\sigma_{1}\left(\widetilde{\mathcal{P}}_{a}^{\mathrm{ext}}\right)$.
Figure 8: Comparison between the numerical results, our bound (6.10), and Xiong’s bound (6.11) for the exterior of rescaled prolate spheroids. The inset on the left zooms onto very small values of $a$.
If we consider an arbitrary, not necessarily convex or star-shaped, domain, there are no nontrivial lower bounds, as the following example shows.
Definition 6.4.
Let $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$, be a bounded open set with a connected $\Omega^{\mathrm{ext}}$. We say that $\Omega^{\mathrm{ext}}$ has a passage of width $\varepsilon>0$ and base $\Gamma\subset\mathbb{R}^{n-1}$ if, subject to a rigid motion change of coordinates, there exists a bounded cylinder $\mathcal{C}_{\varepsilon,\Gamma}:=\Gamma\times(-\epsilon,\epsilon)$ such that $\mathcal{C}_{\varepsilon,\Gamma}\subset\Omega^{\mathrm{ext}}$ and $\partial\mathcal{C}_{\varepsilon,\Gamma}\cap\partial\Omega=\Gamma\times\{\pm\varepsilon\}$, see Figure 9. ∎
Figure 9: An example of $\Omega^{\mathrm{ext}}$ with a passage.
Proposition 6.5.
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded open set with Lipschitz boundary. If the exterior domain $\Omega^{\mathrm{ext}}$ has a passage of width $\varepsilon>0$ and base $\Gamma\subset\mathbb{R}^{n-1}$, then its exterior Steklov eigenvalues satisfy
$\sigma_{k}\left(\Omega^{\mathrm{ext}}\right)<\Lambda^{\mathrm{D}}_{k}(\Gamma)\varepsilon,\qquad k\in\mathbb{N},$
where $\Lambda^{\mathrm{D}}_{k}(\Gamma)$ is the $k$th Dirichlet eigenvalue of $\Gamma$.
Proof.
It follows from the variational characterisation (3.5), that
$\sigma_{k}\left(\Omega^{\mathrm{ext}}\right)\leq\sigma^{\mathrm{D}}_{k}\left(\mathcal{C}_{\varepsilon,\Gamma}\right),$
where $\sigma^{\mathrm{D}}_{k}\left(\mathcal{C}_{\varepsilon,\Gamma}\right)$ is the $k$th eigenvalue of the mixed Steklov–Dirichlet problem
$\begin{cases}\Delta u=0\qquad&\text{in }\mathcal{C}_{\varepsilon,\Gamma},\\ u=0\qquad&\text{on }\partial\Gamma\times[-\varepsilon,\varepsilon],\\ \partial_{\nu}u=\sigma^{\mathrm{D}}u&\text{on }\Gamma\times\{\pm\varepsilon\}.\end{cases}$ (6.12)
It is easily seen by separation of variables that the set of eigenvalues of (6.12) is given by
$\left\{\sqrt{\Lambda_{k}^{\mathrm{D}}(\Gamma)}\tanh\left(\varepsilon\sqrt{\Lambda_{k}^{\mathrm{D}}(\Gamma)}\right),k\in\mathbb{N}\right\}\cup\left\{\sqrt{\Lambda_{k}^{\mathrm{D}}(\Gamma)}\coth\left(\varepsilon\sqrt{\Lambda_{k}^{\mathrm{D}}(\Gamma)}\right),k\in\mathbb{N}\right\}.$
Ignoring the second subset, we conclude that $\sigma^{\mathrm{D}}_{k}\left(\mathcal{C}_{\varepsilon,\Gamma}\right)\leq\sqrt{\Lambda_{k}^{\mathrm{D}}(\Gamma)}\tanh\left(\varepsilon\sqrt{\Lambda_{k}^{\mathrm{D}}(\Gamma)}\right)<\Lambda_{k}^{\mathrm{D}}(\Gamma)\varepsilon$, and the result follows. ∎
If we now consider a family of $\varepsilon$-dependent exterior domains, each containing a passage with the same fixed base and of width $\varepsilon\searrow 0$, then the $k$th Steklov eigenvalue will tend to zero as well. For an illustration in dimension two, let $\Gamma:=\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$, and let $\Omega_{\varepsilon}\subset\mathbb{R}^{2}$ be a disjoint union of two identical squares of side $\pi$, $\Omega_{\varepsilon}:=\Gamma\times(-\pi-\varepsilon,-\varepsilon)\sqcup\Gamma\times(\varepsilon,\pi+\varepsilon)$. Then Proposition 6.5 implies that
$\sigma_{k}\left(\Omega_{\varepsilon}^{\mathrm{ext}}\right)<\varepsilon k^{2},\qquad k\in\mathbb{N}.$
Remark 6.6.
The same result holds for the interior Steklov problem: if $\Omega\subset\mathbb{R}^{n}$ is a bounded open set with Lipschitz boundary, which contains a passage of width $\varepsilon>0$ and base $\Gamma\subset\mathbb{R}^{n-1}$, then $\sigma_{k}\left(\Omega\right)<\Lambda^{\mathrm{D}}_{k}(\Gamma)\varepsilon$ for all $k\in\mathbb{N}$.
§6.3.  Shape optimisation and spectral asymptotics
6.3.1 Proof of Theorem 1.16
Assume for simplicity that $\left|\partial\Omega\right|=2\pi$. As in §3.4, we identify $\mathbb{R}^{2}$ with the complex plane $\mathbb{C}$ and consider the holomorphic map $\phi:\mathbb{C}\setminus\{0\}\to\mathbb{C}\setminus\{0\}$ defined by $\phi(z)=1/z$. Consider, as before, the bounded simply-connected domain $\Omega^{*}$ such that $\phi\left(\Omega^{*}\setminus\{0\}\right)=\Omega^{\mathrm{ext}}$. By the Riemann mapping theorem there exists a biholomorphic map $\psi:B_{1}\to\Omega^{*}$. Therefore, the exterior problem ($\mathrm{ES}$) is equivalent to
$\begin{cases}\Delta w=0\qquad&\text{ in }B_{1},\\ \partial_{\nu}w=\sigma\omega(z)w\qquad&\text{ on }\partial B_{1},\end{cases}$ (6.13)
where the function $\omega:\partial B_{1}\to\mathbb{R}$ is defined by
$\omega(z)=\left|\left(\phi(\psi(z))\right)^{\prime}\right|=\left|\left(\frac{1}{\psi(z)}\right)^{\prime}\right|=\left|\frac{\psi^{\prime}(z)}{\psi(z)^{2}}\right|,$
and satisfies $\int_{\partial B_{1}}\omega\,\mathrm{d}S=|\partial\Omega|=2\pi$. Following [We54, Theorem 4.12], we plan to use the coordinate functions $x_{1},x_{2}:B_{1}\to\mathbb{R}$ as trial functions for this problem. Indeed, using the Hersch trick, it is possible to choose the conformal map $\psi$ so that $\int_{\partial B_{1}}x_{i}\omega\,\mathrm{d}S=0$, leading to
$\sigma_{2}\int_{\partial B_{1}}x_{i}^{2}\omega\,\mathrm{d}S\leq\int_{B_{1}}\left|\nabla x_{i}\right|^{2}\,\mathrm{d}x=|B_{1}|=\pi.$
Summing over $i=1,2$, and using the fact that $x_{1}^{2}+x_{2}^{2}=1$ on $\partial B_{1}$, leads to
$\sigma_{2}\int_{\partial B_{1}}\omega\,\mathrm{d}S\leq 2\pi=\sigma_{2}\left|\partial\Omega\right|.$
In case of equality, arguing in the same was as in [GirPo10b, Section 2.3], we get that each $x_{i}$ is an eigenfunction of (6.13) with eigenvalue one, and it follows that $\omega(z)=1$ for all $z\in\partial B_{1}$. Let $a:=\psi^{-1}(0)\in B_{1}$, and consider the conformal automorphism $\delta_{a}:B_{1}\to B_{1}$ of the unit disk defined by
$\delta_{a}(z):=\frac{z-a}{1-\overline{a}z}.$
Because $\psi$ is a biholomorphic map, $\psi^{\prime}(z)$ does not vanish for $z\in B_{1}$. Moreover, $a$ is the unique zero of $\psi$, and it is simple. It follows that $\frac{1}{\psi(z)}$ is a meromorphic function with a simple pole at $a$. Therefore, the function $\upsilon(z):=\frac{\psi^{\prime}(z)}{\psi(z)^{2}}\delta_{a}(z)^{2}$ has a removable singularity at $a$, and hence it can be extended to a holomorphic function on $B_{1}$. Moreover, $|\upsilon(z)|=1$ for each $z\in\partial B_{1}$. By the maximum modulus principle, this implies that there is a constant $c$ with $|c|=1$, and such that $\upsilon(z)=c$ for all $z\in B_{1}$. As a consequence, for any $z\in B_{1}\setminus\{a\}$ we have,
$\left(\frac{1}{\psi(z)}\right)^{\prime}=c\left(\frac{1-\overline{a}z}{z-a}\right)^{2}.$ (6.14)
Let us compute the residues at $a$ of both sides of (6.14). A direct computation yields
$c\left(\frac{1-\overline{a}z}{z-a}\right)^{2}=c\left(\frac{(1-|a|^{2})^{2}}{(z-a)^{2}}-2\overline{a}\frac{1-|a|^{2}}{z-a}+\overline{a}^{2}\right).$
It follows that the residue of the right-hand side of (6.14) at $a$ is equal to $2c\overline{a}(1-|a|^{2})$. On the other hand, the left-hand side of (6.14) is the derivative of a meromorphic function with a pole at $a$. Using the Laurent series representation of $1/\psi$ we find that the residue of the left-hand side at $a$ is equal to zero. Combining these two calculations implies that $a=0$. Hence,
$\left(\frac{1}{\psi(z)}\right)^{\prime}=\frac{c}{z^{2}}.$
Integrating, we obtain
$\frac{1}{\psi(z)}=\frac{-c}{z}+C,$
where $C\in\mathbb{C}$ is a constant of integration. Therefore,
$\psi(z)=\frac{z}{Cz-c}=\frac{1}{C-\frac{c}{z}},$
and $|C|<1$ since $\psi$ is holomorphic. It is now easy to see that $\Omega^{*}=\psi(B_{1})$ is a disk, and hence $\Omega$ is a disk as well. This completes the proof of the theorem. ∎
6.3.2 Higher dimensions
In higher dimensions the exterior of a ball is known to be a local maximiser for the first eigenvalue among domains with given measure [Bun23], c.f. the eigenvalue curves in Figure 7 in the vicinity of $a=1$. However, even within the class of convex domains, no global isoperimetric bound holds. In this regard, the exterior problem differs significantly from the interior problem.
Recall Example 6.3: the normalised solid spheroids $\widetilde{\mathcal{P}}_{a}$ have constant volume, while the lower bound (6.10) shows that $\lim_{a\searrow 0}\sigma_{1}(\widetilde{\mathcal{P}}_{a}^{\mathrm{ext}})=+\infty$. Hence, the ball does not maximise the first Steklov exterior eigenvalue among domains of fixed volume. Moreover, if the ball were the maximiser among domains with fixed surface area, then by the classical isoperimetric inequality it would also maximise among domains with fixed volume, which contradicts the above example.
An analogous behaviour can be observed for higher-dimensional spheroids. The following example provides an immediate proof of Theorem 1.18 for all $n\geq 3$.
Example 6.7.
Let $n\geq 3$. For $a\in(0,1)$, and $k\in\mathbb{N}$, $2\leq k\leq n-1$, we consider a spheroid
$\mathscr{e}_{a}^{(k)}:=\left\{x\in\mathbb{R}^{n}:\sum_{j=1}^{k}\frac{x_{j}^{2}}{a^{2}}+\sum_{j=k+1}^{n}x_{j}^{2}=1\right\}.$
We write $x=\begin{pmatrix}x^{(1)}\\ x^{(2)}\end{pmatrix}$, with $x^{(1)}\in\mathbb{R}^{k}$ and $x^{(2)}\in\mathbb{R}^{n-k}$, and define $D(x):=1-(1-a^{2})\left|x^{(2)}\right|^{2}$. Then, the principal curvatures of $\mathscr{e}_{a}^{(k)}$ are given by
$\kappa_{1}(x)=\ldots=\kappa_{k-1}(x)=\frac{1}{a\sqrt{D(x)}},\qquad\kappa_{k}(x)=\ldots=\kappa_{n-2}(x)=\frac{a}{\sqrt{D(x)}},\qquad\kappa_{n-1}(x)=\frac{a}{D(x)^{\frac{3}{2}}}.$
Each principal curvature becomes minimal at $\left|x^{(2)}\right|=0$, so the logarithmic mean does also become minimal when $\left|x^{(2)}\right|=0$. Note that we have to understand $L(\kappa_{1},\ldots,\kappa_{n-1})$ as a limit. For $k=2$, this is
$\min_{x\in\mathscr{e}_{a}^{(2)}}L(\kappa_{1}(x),\ldots,\kappa_{n-1}(x))=\frac{1}{-\log(a)\frac{a}{(1-a^{2})^{n-2}}-\frac{a}{(n-3)!}\frac{\mathrm{d}^{n-3}}{\mathrm{d}t^{n-3}}\left(\frac{t^{n-3}\log(t)}{1-at}\right)\Big|_{t=a}}.$
For $1\leq i\leq m$,
$\frac{\mathrm{d}^{i}}{\mathrm{d}t^{i}}\log(t)=(-1)^{i-1}\frac{(i-1)!}{t^{i}},\qquad\frac{\mathrm{d}^{i}}{\mathrm{d}t^{i}}t^{m}=\frac{m!}{(m-i)!}t^{m-i},\qquad\frac{\mathrm{d}^{i}}{\mathrm{d}t^{i}}\frac{1}{1-at}=\frac{a^{i}i!}{(1-at)^{i+1}},$
and applying the Leibniz rule, we obtain
$\frac{\mathrm{d}^{m}}{\mathrm{d}t^{m}}\left(\frac{t^{m}\log t}{1-at}\right)=\frac{m!}{(1-at)^{m+1}}\log(t)+O(t)\qquad\text{as }t\searrow 0.$
Therefore,
$\begin{split}\min_{x\in\mathscr{e}_{a}^{(2)}}L(\kappa_{1}(x),\ldots,\kappa_{n-1}(x))&=\frac{1}{-\log(a)\frac{a}{(1-a^{2})^{n-2}}-\frac{a}{(1-at)^{n-2}}\log(a)+O(a)}\\ &=\frac{(1-a^{2})^{n-2}}{-2a\log(a)+O(a)}\qquad\text{as }a\searrow 0.\end{split}$
So, bound (1.3) implies that for the normalised solid spheroid
$\widetilde{\mathcal{E}}_{a}^{(2)}:=a^{-\frac{2}{n}}\left\{x\in\mathbb{R}^{n}:\frac{x_{1}^{2}+x_{2}^{2}}{a^{2}}+x_{3}^{3}+\ldots+x_{n}^{2}\leq 1\right\},$
with volume $\left|\widetilde{\mathcal{E}}_{a}^{(2)}\right|=\omega_{n}$, we have
$\sigma_{1}\left(\left(\widetilde{\mathcal{E}}_{a}^{(2)}\right)^{\mathrm{ext}}\right)\geq\frac{(1-a^{2})^{n-2}}{-2a^{\frac{n-2}{n}}\log(a)+O\left(a^{\frac{n-2}{n}}\right)}\to+\infty\qquad\text{as }a\searrow 0.$
For $k>2$, the argument carries over in the same way, leading again to $\sigma_{1}\left((\tilde{\mathcal{E}}_{a}^{(k)})^{\mathrm{ext}}\right)\to\infty$ as $a\searrow 0$. In fact, the divergence is faster, which can be understood informally from the fact that the principal curvature $\frac{1}{a\sqrt{D(x)}}$ is weighted more heavily in the logarithmic mean as $k$ increases. ∎
Another interesting non-convex example demonstrating the absence of an isoperimetric inequality is presented below.
Example 6.8.
Let $M$ be a smooth compact Riemannian manifold of dimension $m\geq 3$, without boundary. Let $N\subset M$ be a closed submanifold of positive codimension and for $\varepsilon>0$, consider the tubular neighbourhood $N_{\varepsilon}=\{x\in M:\operatorname{dist}(x,N)<\varepsilon\}$. Brisson [Bri22] studied the Steklov spectrum of $N_{\varepsilon}^{\mathrm{ext}}:=M\setminus N_{\varepsilon}$. She proved that
$\lim_{\varepsilon\to 0}\sigma_{1}\left(N_{\varepsilon}^{\mathrm{ext}}\right)\left|\partial N_{\varepsilon}^{\mathrm{ext}}\right|^{\frac{1}{m-1}}=\infty.$
Because the approach in [Bri22] is based on quasi-isometries and mixed Steklov–Neumann eigenvalue problems, it can be adapted verbatim in the setting of this paper to obtain the following result, via the mixed approach of Theorem 3.9.
Proposition 6.9.
Let $m\geq 3$. Consider a compact smooth submanifold $N\subset\mathbb{R}^{m}$ of dimension $n\leq m-2$. Then, the exterior Steklov eigenvalues of the tubular neighbourhood $N_{\varepsilon}$ satisfy
$\lim_{\varepsilon\to 0}\sigma_{1}\left(N_{\varepsilon}^{\mathrm{ext}}\right)\left|\partial N_{\varepsilon}\right|^{\frac{1}{m-1}}=\infty.$
This applies, in particular, to neighbourhoods of closed curves in $\mathbb{R}^{3}$. ∎
Remark 6.10.
In [Xio23, Theorem 4], several upper bounds for $\sigma_{1}$ are established. Another such bound was obtained in [Bun25, Theorem 3], following the approach of [GioSm05, Section 2], which shows the following. Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded Lipschitz domain, and $B_{\rho}\subset\Omega$ for some $\rho>0$. Then $\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)\leq\sigma_{1}\left(B_{\rho}^{\mathrm{ext}}\right)$. As an immediate consequence of Theorem 1.16, the same inequality also holds for the first nontrivial Steklov eigenvalue in dimension two: $\sigma_{2}\left(\Omega^{\mathrm{ext}}\right)\leq\sigma_{2}(B_{\rho}^{\mathrm{ext}})=\frac{1}{\rho}$.
The same idea as in [Bun25, Theorem 3] can be used to prove, for any $\Lambda>0$, that
$B_{\rho}\subset\Omega\quad\implies\quad\mu_{1}\left(\Omega^{\mathrm{ext}},\Lambda\right)\leq\mu_{1}\left(B_{\rho}^{\mathrm{ext}},\Lambda\right).$
Let $u(x)=f(|x|)$ denote the eigenfunction corresponding to $\mu_{1}\left(B_{\rho}^{\mathrm{ext}},\Lambda\right)=\mu_{(0)}\left(B_{\rho}^{\mathrm{ext}},\Lambda\right)$, see (3.9). We define the function
$\widehat{\mu}:\partial\Omega\to\mathbb{R},\quad y\mapsto\frac{\left\langle\nabla f(|y|),\nu_{y}\right\rangle}{f(|y|)}.$
Note that $f^{\prime}(\rho)<0$, and that $\frac{f^{\prime}(r)}{f(r)}=\frac{-\Lambda K_{\frac{n}{2}}(\Lambda r)}{K_{\frac{n-2}{2}}(\Lambda r)}$ is monotonically increasing in $r$, see [KrLo20, proof of Proposition 3]. Thus,
$\widehat{\mu}(y)\leq\frac{-f^{\prime}(|y|)}{f(|y|)}\leq\frac{-f^{\prime}(\rho)}{f(\rho)}=\mu_{1}(B_{\rho}^{\mathrm{ext}},\Lambda)\qquad\text{for all }y\in\partial\Omega.$
Furthermore, $u$ solves
$\begin{cases}(\Lambda^{2}-\Delta)u=0\qquad&\text{ in }\Omega^{\text{ext}}\subset B_{\rho}^{\mathrm{ext}},\\ \partial_{\nu}u=\widehat{\mu}u\qquad&\text{ on }\partial\Omega.\end{cases}$
Thus, integration by parts yields
$\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}u^{2}\,\mathrm{d}x=-\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x+\int_{\partial\Omega}\widehat{\mu}u^{2}\,\mathrm{d}S\leq-\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x+\mu_{1}(B_{\rho}^{\mathrm{ext}},\Lambda)\int_{\partial\Omega}u^{2}\,\mathrm{d}S.$
Therefore,
$\mu_{1}(B_{\rho}^{\mathrm{ext}},\Lambda)\geq\frac{\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}u^{2}\,\mathrm{d}x+\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}u^{2}\,\mathrm{d}S}\geq\mu_{1}\left(\Omega^{\mathrm{ext}},\Lambda\right),$
which completes the proof. ∎
6.3.3 Proofs of eigenvalue asymptotics
Proof of Proposition 1.20.
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded open set with $C^{2,\alpha}$ boundary for some $\alpha>0$. For $R>R_{0}(\Omega)$, it easily follows from Lemma 3.8 (see also Lemma 5.2) that
$\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)\leq\sigma_{k}\left(\Omega^{\mathrm{ext}}\right)\leq\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ (6.15)
for every $k\geq 1$. Inspecting the proof of [GirKLP22, Theorem 1.11], we conclude that it can be applied verbatim to the mixed boundary problems on $\Omega^{\mathrm{ext}}_{R}$ with Steklov condition on $\partial\Omega$ and either Dirichlet or Neumann condition on $\partial B_{R}$. Therefore,
$\left\{k\in\mathbb{N}\,:\,\sigma_{k}^{D,N}\left(\Omega^{\mathrm{ext}}_{R}\right)<\sigma\right\}=\frac{\omega_{n-1}}{(2\pi)^{n-1}}|\partial\Omega|\sigma^{n-1}+O\left(\sigma^{n-2}\right)\qquad\text{as }\sigma\to+\infty.$
Combining this with (6.15) yields the result. ∎
Proof of Proposition 1.22.
Let us use the conformal approach and consider the spectrum of the weighted interior problem ($\mathrm{CT}$). Note that $\Omega^{*}$ is a Lipschitz domain, since an inversion with the center away from the boundary preserves the Lipschitz regularity of the boundary. Therefore, it follows from [KarLaPo23, Theorem 1.8] that the eigenvalue counting function for the problem ($\mathrm{CT}$) satisfies the asymptotics (1.11). ∎